### v1

# Unifying KV Cache Compression for Large Language Models with LeanKV

Abstract.
Large language models (LLMs) demonstrate exceptional performance but incur high serving costs due to substantial memory demands, with the key-value (KV) cache being a primary bottleneck. Existing KV cache compression methods, including quantization and pruning, struggle with limitations such as uniform treatment of keys and values and static memory allocation across attention heads. To address these challenges, we introduce LeanKV, a unified KV cache compression framework that enhances LLM serving efficiency without compromising accuracy through three innovations: (1) Hetero-KV quantization, which stores keys at a higher precision than values to reflect their greater impact on attention computations; (2) per-head dynamic sparsity, which allocates memory based on token importance per head and per request; and (3) unified KV compression, integrating mixed-precision quantization and selective pruning to enable a smooth tradeoff between model accuracy and memory efficiency. To efficiently support these techniques, LeanKV introduces systems optimizations including unified paging and on-GPU parallel memory management. Implemented on vLLM, LeanKV compresses the KV cache by 3.0×3.0$\times$3.0 × to 5.0×5.0$\times$5.0 × without accuracy loss and up to 11.0×11.0$\times$11.0 × with under 5% accuracy loss, enhancing throughput by 1.9×1.9$\times$1.9 × to 2.5×2.5$\times$2.5 ×, and up to 6.9×6.9$\times$6.9 ×.
*Equal contribution.
1. Introduction
Large language models (LLMs) like GPT (brown2020language, ; achiam2023gpt, ; ouyang2022training, ) and Gemini (team2023gemini, ) have demonstrated significant potential to impact our daily lives, offering promising applications in areas including chatbots (vicuna, ; chiang2024chatbot, ), programming (gitcopilot, ; guo2024deepseek, ), mathematics (shao2024deepseekmath, ; zhong2024evaluation, ) and medical assistance (saab2024capabilities, ; xie2024preliminary, ). Despite their exceptional performance, hosting LLMs is costly due to their large model sizes, demanding extensive hardware resources. Given the pervasive adoption of LLMs, enhancing serving efficiency has become critically important (aminabadi2022deepspeed, ; fang2021turbotransformers, ; kao2023flat, ; shi2023welder, ; zhang2023shepherd, ; han2022microsecond, ; pope2023efficiently, ; li2023alpaserve, ).
LLMs are fundamentally based on the Transformer architecture (vaswani2017attention, ; wolf2020transformers, ), which employs the attention mechanism to generate tokens sequentially, based on the input prompt and previously generated tokens. Within the attention mechanism, each token is transformed into query, key and value feature vectors. The attention scores, derived from the dot product of the query and key vectors, enable the model to weigh the significance of different tokens. Subsequently, the output is computed as the weighted sum of the value vectors, based on the attention scores. To capture diverse relationships between tokens, the attention mechanism is typically divided into multiple heads (vaswani2017attention, ; shazeer2019fast, ; ainslie2023gqa, ). Each head independently computes attention, and their outputs are concatenated and transformed to create a richer representation of the input data
To avoid redundant computation, LLM inference frameworks typically cache intermediate key and value tensors in memory, commonly referred to as the KV cache (kwon2023efficient, ). The size of the KV cache scales linearly with both sequence length and the number of concurrent requests, often comprising over 90% of total memory consumption (kwon2023efficient, ; yu2022orca, ; gao2018low, ; yang2022infless, ). As state-of-the-art models continue to support longer sequences (dubey2024llama, ; jiang2024mixtral, ; liu2024deepseek, ; achiam2023gpt, ; team2023gemini, ; touvron2023llama, ), the KV cache has emerged as a critical bottleneck for LLM serving efficiency. It limits the number of concurrent requests that can fit in memory and increases attention computation latency due to its memory bandwidth-bound nature (dao2022flashattention, ; dao2023flashattention, ).
Recognizing the critical role of the KV cache in limiting serving efficiency, researchers have investigated various compression techniques, primarily focused on quantization (lin2024qserve, ; flashinfer, ; hooper2024kvquant, ) and pruning (zhang2024h2o, ; xiao2023efficient, ; zhang2024pyramidkv, ; ge2023model, ; li2024snapkv, ). Quantization reduces the size of the KV cache by converting floating-point feature values into integers, representing them based on pre-defined intervals. Pruning, on the other hand, reduces the token sequence length by eliminating tokens deemed insignificant based on attention scores. Despite the promise of these techniques, both of them face limitations in efficiency and practicality. First and foremost, existing quantization methods apply the same treatment to key and value vectors of all tokens, disregarding the different roles of keys and values in attention computation, as well as the varying significance of different tokens. Second, current pruning methods statically allocate memory across attention heads, failing to account for the dynamic attention patterns across heads and requests. Moreover, these pruning methods have not been comprehensively evaluated on workloads requiring advanced reasoning, such as programming and mathematics. Lastly, prior work has largely overlooked the potential of unifying quantization and pruning to achieve even greater serving efficiency.
In this paper, we propose LeanKV, a unified KV cache compression framework that addresses the limitations of existing methods and enhances LLM serving efficiency. LeanKV introduces three algorithmic principles :
1. Hetero-KV. LeanKV compresses key and value vectors using heterogeneous precisions, with key vectors stored at higher precision and value vectors at lower precision. This design leverages the observation that key vectors have a greater influence on attention calculations, as they affect the attention scores of all tokens, whereas value vectors only impact the contributions of individual tokens to the attention output.
2. Per-head Dynamic Sparsity. LeanKV dynamically identifies critical tokens on a per-head and per-request basis, adjusting memory allocation accordingly. This design stems from the observation that attention score distributions exhibit dynamic sparsity across both heads and requests. Within a single request, the number of critical tokens varies across heads, and for the same head, the number of critical tokens also varies across different requests.
3. Unified KV Compression. Built on Hetero-KV quantization and per-head dynamic sparsity, LeanKV unifies KV compression strategies by applying mixed-precision quantization and selective pruning on a per-head and per-request basis, with token importance serving as a unified metric. Critical tokens are stored at higher precision, less important tokens are quantized at lower precision, and the least significant tokens are pruned. This integrated approach enables a smooth tradeoff between model accuracy and memory efficiency.
Implementing LeanKV’s algorithmic innovations presents new challenges in system design, particularly regarding memory management. Unlike traditional attention mechanisms that adopt uniform memory allocation across attention heads, LeanKV requires dynamic, per-head memory allocation for each request, significantly increasing system complexity. Moreover, LeanKV must flexibly configure memory to accommodate varying numbers of tokens stored at different precision levels or pruned entirely, further complicating memory management.
To address these challenges, LeanKV introduces two novel techniques. First, LeanKV employs unified paging, which divides memory into fixed-size pages containing vectors stored at various precisions. Each page is dynamically parsed based on the precision configuration required by the head and request, while an attention GPU kernel is co-designed with the page layout to maximize memory bandwidth utilization. Second, LeanKV introduces an on-GPU memory management mechanism that accelerates memory management by parallelizing it directly on the GPU with minimal metadata overhead.
We implement LeanKV on vLLM (kwon2023efficient, ), a state-of-the-art LLM inference system, and evaluate it across three prominent model families using representative workloads, including tasks that demand advanced reasoning such as mathematics and coding. Our experimental results demonstrate that LeanKV offers a superior cost-accuracy tradeoff compared to existing KV cache management methods. Specifically, LeanKV compresses the KV cache by 3.0×3.0$\times$3.0 × to 5.0×5.0$\times$5.0 × without loss in model accuracy, and by up to 11.0×11.0$\times$11.0 × with a marginal accuracy reduction of under 5%, resulting in throughput improvement of 1.9×1.9$\times$1.9 × to 2.5×2.5$\times$2.5 ×, and up to 6.9×6.9$\times$6.9 × correspondingly.
2. Background and Related Work
In this section, we describe Transformer-based large language models and review existing techniques for KV cache management aimed at improving LLM serving efficiency.
2.1. Large Language Models
Transformer Architecture. Large language models (LLMs) are predominantly built upon the Transformer architecture (vaswani2017attention, ), which is used to capture complex long-range dependencies in sequential data. At the core of Transformer architecture is the attention mechanism, which allows each token in a sequence to weigh the significance of other tokens when constructing its contextualized representation. Mathematically, the attention computation is defined as:
(1) $\operatorname{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})_{i}$ $=\sum_{j=1}^{i}\operatorname{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)_{ij}\mathbf{v}_{j}$
$=\sum_{j=1}^{i}\frac{\exp\left(\frac{\mathbf{q}_{i}\cdot\mathbf{k}_{j}}{\sqrt{d}}\right)}{\sum_{n=1}^{i}\exp\left(\frac{\mathbf{q}_{i}\cdot\mathbf{k}_{n}}{\sqrt{d}}\right)}\mathbf{v}_{j}$
$\mathbf{Q}$ and $\mathbf{K}$ are matrices of size $l\times d$, representing the queries and keys, respectively, where $l$ denotes the number of tokens and $d$ represents the feature dimensionality. Vectors $\mathbf{q}_{i}$, $\mathbf{k}_{i}$, and $\mathbf{v}_{i}$ correspond to the query, key, and value of the $i^{\text{th}}$ token. The dot product between $\mathbf{Q}$ and $\mathbf{K}$ measures the relevance between each pair of query and key. The softmax function normalizes these dot products to form a probability distribution, referred to as attention scores, which are used to compute a weighted sum of the value vectors and produce the output.
To capture a broader range of interactions between tokens, Transformer models employ multi-head attention (MHA), where each head independently computes attention using distinct projections of queries, keys, and values. The outputs from different heads are concatenated and linearly transformed to create a more expressive representation. Grouped-query attention (GQA) (ainslie2023gqa, ) improves the efficiency of MHA by allowing a group of query heads to share the same projection of keys and values, referred to as a KV head.
Autoregressive Generation and KV Cache. The autoregressive generation process of LLMs consists of two phases: the prompt phase, during which the model computes latent representations for all tokens in the prompt, the user-provided context, and generates the first new token, and the generation phase, where the model iteratively generates subsequent tokens by attending to both the prompt and previously generated tokens. To avoid redundant computations across generation steps, KV cache is introduced to store the keys and values of all previous tokens. However, the size of the KV cache grows linearly with both the sequence length and batch size, quickly becoming a bottleneck for inference throughput (yu2022orca, ; kwon2023efficient, ). Therefore, efficient KV cache management is critical for alleviating memory bottlenecks and improving LLM serving efficiency.
2.2. KV Cache Optimization
PagedAttention. Static KV cache management systems (yu2022orca, ; FasterTransformer, ) reserve memory for the maximum possible sequence length, leading to considerable memory waste when actual sequences are shorter. To mitigate this inefficiency, vLLM (kwon2023efficient, ) introduces PagedAttention, which partitions the KV cache into pages, each containing a fixed number of tokens, and allocates the pages on-demand as the sequence length grows. By managing the KV cache at the granularity of pages, vLLM reduces memory waste and enables larger batch sizes to improve serving efficiency.
KV Cache Quantization. Quantization (nagel2021white, ; dettmers2022gpt3, ; guo2023olive, ; hooper2024kvquant, ) reduce the KV cache size by approximating high-precision floating points with discrete low-bit values. To quantize a given tensor $\mathbf{X}$, we first calculate the scale $s$ and zero point $z$ as shown in Eq. (2), where $\mathbf{X}_{\text{max}}$ and $\mathbf{X}_{\text{min}}$ represent the maximum and minimum values in $\mathbf{X}$, and $b$ is the bit-width. Then, quantization is applied element-wise as: $\mathbf{Q}=\operatorname{round}\left(\frac{\mathbf{X}+z}{s}\right)$. During inference, the original values are approximately reconstructed via dequantization: $\hat{\mathbf{X}}=s\cdot\mathbf{Q}-z$. Unlike the quantized tensor $\mathbf{Q}$, the metadata $s$ and $z$ are kept in higher precision (e.g., FP16) to ensure more accurate dequantization.
(2) $s=\frac{\mathbf{X}_{\text{max}}-\mathbf{X}_{\text{min}}}{2^{b}-1},\quad z=-\mathbf{X}_{\text{min}}$
State-of-the-art KV cache quantization methods such as Atom (zhao2024atom, ) and Qserve (lin2024qserve, ) apply this process to each key and value vector independently. However, these uniform quantization methods, which apply a fixed bit-width across keys and values of all tokens, may not be optimal. They overlook the varying importance of tokens as well as the different roles of keys and values in the attention calculation.
KV Cache Pruning. Several recent works (zhang2024h2o, ; xiao2023efficient, ; li2024snapkv, ; zhang2024pyramidkv, ; liu2024scissorhands, ; ge2023model, ) have explored KV cache pruning to alleviate the memory bottleneck in LLM inference, which can be considered as an extreme case of quantization. These methods use attention scores to assess token importance and evict less important tokens, thereby reducing the KV cache size. The effectiveness of these methods, however, is limited by their static and inflexible allocation of memory resources. Specifically, H2O (zhang2024h2o, ) and SnapKV (li2024snapkv, ) allocate the same memory budget uniformly across all attention heads and layers. PyramidKV (zhang2024pyramidkv, ) allocates more memory in lower layers and less in higher ones, based on the empirical observation that the number of important tokens shrinks as the model depth increases, but still relies on static heuristics and cannot dynamically adjust cache allocations to suit various workloads. Complementary to KV cache pruning, another line of work (sheng2023flexgen, ; lee2024infinigen, ) addresses GPU memory constraints by offloading the entire KV cache to CPU memory and selectively loading important tokens to the GPU for attention computations. However, these methods do not inherently reduce the KV cache size and incur significant latency due to data transfers between the CPU and GPU.
3. Motivation
In this section, we present the key findings that underpin the design of LeanKV, as well as their implications on KV cache management.
3.1. Impact of Key And Value
While the significance of different tokens can be directly inferred from attention scores, the roles of the key and value vectors within a token are less obvious. However, examining Equation 1 reveals a divergence in their impacts. Each token’s contribution to the attention output depends on two factors: the attention score from softmax and the value vector. Key vectors, as part of the shared softmax denominator, influence the attention scores of all tokens, whereas value vectors only affect their respective token’s contribution to the output.
Beyond the broader impact of key vectors, it is crucial to establish the relative significance of attention scores, determined by key vectors, over value vectors. Without this, the significance of key vectors could be overshadowed by that of value vectors. To this end, we reformulate the attention mechanism as a weighted sum of unit vectors, as presented in Equation 3. This formulation decomposes each input token’s contribution into two components: the unit vector $\frac{v_{j}}{|v_{j}|}$, representing only the direction of the token in feature space, and a coefficient, defined as the product of the attention score and the norm of the value vector, which determines the relative significance of the token’s direction. This allows us to assess the relative importance of attention scores and value vectors by examining their contributions to these coefficients.
(3) $\operatorname{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})_{i}$ $=\sum_{j=1}^{i}\underbrace{\operatorname{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)_{ij}|\mathbf{v}_{j}|}_{\text{Coefficient}}\underbrace{\frac{\mathbf{v}_{j}}{|\mathbf{v}_{j}|}}_{\text{Unit vector}}$
Figure 1 presents the distributions of the average attention score per token and the norms of value vectors across three representative layers of the Llama3-8B model (dubey2024llama, ), evaluated on the first one thousand articles from the Wikitext dataset (merity2016pointer, ). Notably, the attention scores span seven orders of magnitude, vastly exceeding the range of value vector norms, which only cover at most two orders of magnitude. Similar patterns are also observed in the larger Llama3-70B model. This pronounced disparity highlights the pivotal role of attention scores in determining each token’s contribution to the attention output. We therefore conclude that key vectors exert a broader and more impactful influence than value vectors, motivating further exploration of processing key and value vectors at distinct precision levels.
Figure 1. Distribution of attention score and value vector norm in Llama3-8B.
3.2. Dynamic Sparsity Patterns
Existing approaches to KV cache pruning primarily rely on static memory allocation for each transformer layer in LLMs, either distributing memory uniformly (xiao2023efficient, ; zhang2024h2o, ; ge2023model, ; li2024snapkv, ) or allocating it based on offline profiling (zhang2024pyramidkv, ). However, LLMs exhibit dynamic sparsity patterns across both attention heads and individual requests, which static memory allocation cannot adequately handle. To evaluate the sparsity of attention computation, we analyze the minimum number of critical tokens required to retain the majority of information, specifically by preserving a target percentage (e.g., 95%) of the total attention score.
We begin by investigating dynamic sparsity patterns across layers. Using the first thousand articles from the Wikitext dataset, we evaluate the Llama3-8B model. Figure 2 illustrates the average number of critical tokens per layer, aggregated across all KV heads, required to retain 95% of the total attention score. Vertical bars represent the standard deviation, capturing variability across individual requests. Notably, the degree of sparsity varies considerably across layers, as indicated by differences in the number of critical tokens.
Next, we delve into the dynamic sparsity within individual layers. Figure 3 presents the average number of critical tokens per KV head for three representative layers, with horizontal bars indicating the standard deviation across requests. The sparsity pattern remains highly dynamic: within each layer, the number of critical tokens varies significantly across KV heads. Furthermore, within individual KV heads, the number of critical tokens can vary substantially across requests, as shown by the high standard deviation in some KV heads.
These highly dynamic sparsity patterns, across layers, attention heads, and individual requests, underscore the necessity of adaptive memory management on a per-head and per-request basis.
Figure 2. Number of critical tokens per layer in Llama3-8B to preserve 95% of the total attention score.
Figure 3. Number of critical tokens per KV head in Llama3-8B to preserve 95% of the total attention score.
3.3. Unifying Quantization and Pruning
Quantization compresses the KV cache by representing feature values with a reduced bit width, whereas pruning can be viewed as an extreme form of quantization: effectively storing feature values with zero bits. By using attention scores to measure the significance of tokens, we can unify the spectrum of KV cache compression by storing progressively less critical tokens at decreasing precision levels, ultimately leading to the pruning of the least critical tokens.
To investigate the feasibility of this unified approach, we experiment with a simple design: quantizing both the key and value vectors of significant tokens to 8 bits while pruning insignificant ones. Figure 4 compares the performance of the FP16 baseline, 8-bit quantization, pruning, and the combination of 8-bit quantization with pruning on the CNN/Daily Mail summarization dataset using the Llama3-8B model. For both pruning and the combined approach, approximately 20% of insignificant tokens are pruned. The ROUGE-1 score, which measures the fidelity of the LLM-generated responses to reference answers, is nearly indistinguishable among the four approaches. However, the combination of 8-bit quantization and pruning achieves the lowest memory usage. The improved efficiency of this combined approach motivates us to further explore the spectrum of KV cache compression by introducing varying levels of precision to accommodate tokens of differing significance.
Figure 4. Rogue-1 score and normalized KV cache memory usage of the Llama3-8B model on CNN/DailyMail.
3.4. Key Findings and Implications
In this section, we outline the key findings that motivate the design of LeanKV. Specifically, we demonstrate that within individual tokens, key vectors have a greater impact than value vectors, inspiring us to propose storing key vectors at higher precision than value vectors. Beyond that, the significance of different tokens, as indicated by softmax scores, exhibits dynamic sparsity patterns across both heads and requests. These patterns necessitate flexible memory management on a per-head and per-request basis. Lastly, we show that unifying quantization and pruning can outperform either approach when used in isolation.
These findings motivate us to further explore the spectrum of KV cache compression, on a per-head and per-request basis to achieve the maximum memory efficiency. Specifically, each attention head should be allowed to store tokens at varying precision levels and to prune the least important ones, according to the sparsity patterns of individual requests. Additionally, within each token, the key and value vectors should be stored in different precisions. However, this unified approach presents significant challenges on memory management systems due to the resulting irregular memory layout, where the KV cache of each head and request consists of varying number of tokens stored at different precisions.
4. Unified Compression Algorithm
In this section, we present the design of LeanKV’s unified compression algorithm, which leverages the insights from Section 3 to optimize memory efficiency. To account for the greater impact of key vectors compared to value vectors, we propose Hetero-KV, a strategy that quantizes keys at a higher precision than values. For instance, key and value vectors can be quantized to 8 and 4 bits (K8V4), or to 4 and 2 bits (K4V2). Additionally, we propose per-head memory management tailored to dynamic sparsity patterns. Rather than imposing a fixed memory budget, each head adaptively determines its memory allocation by quantizing tokens at various precision levels or pruning, based on the attention score distributions of individual requests. Specifically, tokens are categorized into three significance classes: the most important tokens are quantized at high precision (e.g., K8V4), moderately important tokens at low precision (e.g., K4V2), and the least important tokens are pruned.
Next, we describe the unified compression algorithm in greater detail, addressing the prompt phase and the generation phase separately. Notably, in both phases, compression is conducted on a per-head and per-request basis.
Prompt Phase. In the prompt phase, key and value vectors for all tokens in the prompt are computed. The compression algorithm then determines the appropriate precision for storing each token based on its significance. The significance of the $i^{th}$ token is calculated by averaging the $N-i$ attention scores it receives from subsequent tokens, where $N$ denotes the sequence length. This average is then normalized to ensure that the total significance across all tokens sums to one.
To prevent premature compression, the algorithm reserves a recent window of size $W$, where all tokens are quantized at a high precision. $W$ is predefined and typically set to 64. The remaining tokens are then sorted by significance. The least important $N_{p}$ tokens, whose cumulative significance is below the threshold $T_{l}$, are pruned. The next $N_{l}$ tokens, with a cumulative significance less than $T_{h}$, are quantized to low precision, while the most important $N_{h}$ tokens are quantized to high precision. Here, $N_{p}+N_{l}+N_{h}+W=N$, and $T_{l}$ and $T_{h}$ are predefined thresholds determined by offline profiling. As a result of this mixed-precision compression, the KV cache is conceptually divided into two parts: a high precision section $KV_{h}$ and a low precision section $KV_{l}$.
Generation Phase. In the generation phase, a single token is compressed at each step, aligning with the autoregressive nature of the generation process. The most recent token is added to the recent window to prevent premature compression, while the earliest token $t_{c}$ in the window becomes a candidate for more aggressive compression. The compression procedure can be divided into two parts. First, token $t_{c}$ is quantized at either a high or low precision and added to the corresponding section of the KV cache, or it may be completely pruned. Next, if $t_{c}$ is quantized, the least significant token $t_{v}$ in the corresponding precision section of the KV cache is either re-quantized to lower precision or pruned, depending on its significance. Essentially, this policy establishes a smooth downgrading path for the less important token: rather than being pruned directly, it is first re-quantized to low precision, with pruning occurring only if it remains insignificant.
The detailed procedure is outlined in Algorithm 1 and is governed by two predefined thresholds: the high precision threshold $T_{h}$ and the low precision threshold $T_{l}$. If the significance of $t_{c}$ is at least $T_{h}$, it is quantized at high precision and added to $KV_{h}$, the high precision section of the KV cache (line 6). Subsequently, the least significant token in $KV_{h}$ is identified as the victim token $t_{v}$ for more aggressive compression (line 7). If $t_{v}$’s significance exceeds $T_{h}$, it remains in $KV_{h}$; if it falls between $T_{h}$ and $T_{l}$, $t_{v}$ is re-quantized to low precision and moved to $KV_{l}$, the low precision section of the KV cache (line 9); otherwise, $t_{v}$ is pruned. Similarly, if $t_{c}$’s significance falls between $T_{h}$ and $T_{l}$, it is quantized to low precision and added to $KV_{l}$ (line 14). The least important token in $KV_{l}$ is designated as the victim $t_{v}$ (line 15), and is further pruned if its significance falls below $T_{l}$ (line 17).
Notably, the proposed unified compression algorithm is extensible beyond three precision levels, and we adopt three precision levels to minimize metadata overhead and improve system efficiency, as detailed in the following section.
1:  Input: Thresholds $T_{h}$ $T_{l}$; Candidate token $t_{c}$
2:  Input: High & low precision $P_{h}$ $P_{l}$
3:  Input: High & low precision KV cache $KV_{h}$ $KV_{l}$
4:  Function: Significance $Score$; Quantization $Quant$;
5:  if Score($t_{c}$) $\geq$ $T_{h}$ then
6:     $KV_{h}$.add(Quant($t_{c}$, $P_{h}$))
7:     $t_{v}$ = $argmin_{t\in KV_{h}}$(Score($t$))
8:     if $T_{l}\leq$ Score($t_{v}$) $<T_{h}$ then
9:        $KV_{h}$.remove($t_{v}$),  $KV_{l}$.add(Quant($t_{v}$, $P_{l}$))
10:     else if Score($t_{v}$) ¡ $T_{l}$ then
11:        $KV_{h}$.remove($t_{v}$)
12:     end if
13:  else if Score($t_{c}$) $\geq$ $T_{l}$ then
14:     $KV_{l}$.add(Quant($t_{c}$, $P_{l}$))
15:     $t_{v}$ = $argmin_{t\in KV_{l}}$(Score($t$))
16:     if Score ($t_{v}$) ¡ $T_{l}$ then
17:        $KV_{l}$.remove($t_{v}$)
18:     end if
19:  end if
Algorithm 1 Unified compression (generation phase)
5. System Design
In this section, we present the design of the LeanKV system that leverages the proposed unified KV cache compression algorithm to boost LLM inference efficiency. We first discuss the challenges that the compression algorithm poses to system performance, and then present novel techniques to tackle these challenges.
5.1. Challenges
The Need for Flexible Paging. In PagedAttention (kwon2023efficient, ), all tokens are stored at the same precision, making a fixed page format sufficient. However, Hetero-KV and mixed-precision quantization introduce varying precision levels across different tokens, rendering a fixed page format inadequate. Specifically, a fixed page format requires conservatively allocating high-precision slots for all tokens regardless of their actual precisions, leading to considerable memory wastage. Suppose memory is allocated to accommodate the highest precision K8V4, a token with K4V2 precision would waste 50% of the memory. Additionally, such a fixed page format results in misaligned memory accesses, which hinder memory bandwidth utilization and reduce overall computational efficiency. To fully realize the benefits of Hetero-KV and mixed-precision quantization, we need a more flexible paging design that can dynamically adapt to varying precision levels, while ensuring compact memory usage and efficient memory accesses.
Increased Memory Management Overhead. The per-head KV cache compression algorithm in LeanKV introduces both significant latency and metadata overhead in memory management. In PagedAttention, the memory management complexity is $O(\#requests)$ regardless of the number of heads, as it partitions memory uniformally across heads and assigns identical page IDs to each head. In contrast, LeanKV must handle varying numbers of high-precision and low-precision tokens across different heads, which increases the complexity to $O(\#requests\times\#heads)$. If not carefully optimized, this overhead could overshadow the performance gains from cache compression. Moreover, to track both high-precision and low-precision tokens, a straightforward approach would involve using separate data structures for each precision, resulting in increased metadata overhead.
5.2. On-GPU Memory Management
A detailed analysis of the memory management process reveals opportunities for parallelization and acceleration. Specifically, memory management for the unified compression algorithm consists of two phases: planning and coordination. In the planning phase, each head independently determines its memory allocation requirements. The subsequent coordination phase synchronizes these per-head requirements and maps them to the GPU’s physical memory. For instance, assume two attention heads require 16 and 32 bytes, respectively, as determined in the planning phase, with all GPU physical memory available. Then in the coordination phase, physical memory bytes 0 to 15 are allocated to the first head, and bytes 16 to 48 to the second. The primary source of increased memory management complexity lies in the planning phase, which is perfectly parallel and well-suited to the GPU’s parallel compute capabilities. Parallelizing the coordination phase is more challenging, as it requires synchronization across heads. However, we observe that this phase can be parallelized effectively via parallel prefix sum (nakano2012prefix-sum, ).
Thus, we design an on-GPU memory management system to efficiently handle per-head dynamic memory allocation and recycling. This system introduces three GPU-resident data structures that collectively tackle the aforementioned challenges: the $(1)$ unified paging to enable flexible paging, $(2)$ the circular free page list to efficiently execute the coordination phase in memory management, and $(3)$ the bidirectional page table to minimize metadata overhead for tracking high-precision and low-precision tokens.
Unified Paging. We partition the available GPU memory into evenly sized pages, each configured upon allocation to store tokens at a given precision. Each unified page consists of six segments: quantized keys, quantization metadata for keys, quantized values, quantization metadata for values, token scores, and positions. The quantization metadata includes scales and zero points for the key and value vectors. The number of tokens stored per page is adjusted according to the quantization configuration, ensuring compact memory usage. Unified paging allows flexible handling of tokens with varying precisions while maintaining a consistent memory layout, effectively supporting our unified compression algorithm. Furthermore, by consolidating keys, values, and their metadata into a single structure, unified paging enhances data locality and eliminates the need for scattered lookups, improving memory access efficiency during attention computation.
Circular Free Page List. To facilitate efficient page allocation and recycling, we manage all available page IDs in a centralized, GPU-resident data structure called circular free page list. The list tracks free pages using a pair of start and end pointers, which wrap to the beginning of the list upon reaching the end, maintaining a circular structure. When a page is allocated, the start pointer increments to the next available page ID; when a page is freed, the end pointer increments to add the freed page ID back to the list.
Notably, both the available and unavailable regions in the list remain contiguous, enabling the coordination phase in memory management to be parallelized via parallel prefix sum. Specifically, after each head determines the number of pages to be allocated or freed in the planning phase, a parallel prefix sum operation computes a unique offset for each head relative to the start or end pointer, ensuring non-conflicting operations within the list. For memory allocation, each head concurrently retrieves its new page IDs from its designated region in the list, with the start pointer incremented by the cumulative number of pages required across all heads. Similarly, for memory recycling, each head concurrently writes freed page IDs to it designated region, with the end pointer incremented by the total number of pages released.
Bidirectional Page Table. Similar to PagedAttention, LeanKV maintains a page table for each attention head, mapping requests to physical memory. Each entry in the page table is a list of page IDs allocated to the head for a specific request. To avoid the doubled metadata overhead associated with maintaining separate page tables for high- and low-precision pages, LeanKV introduces a unified, GPU-resident data structure called the bidirectional page table. This structure efficiently supports the use of three precision levels for KV cache compression, as proposed in Section 4, minimizing the metadata overhead. In each entry of the bidirectional page table, high-precision page IDs grow from the left side of the list, while low-precision page IDs grow from the right, dynamically adapting to the precision requirements of the workload. The length of each page table entry is determined by the maximum sequence length divided by the tokens per high-precision page, ensuring no overflow since low-precision pages always contain more tokens than high-precision ones. This unified approach not only minimizes metadata overhead but also eliminates the need for separate lookups based on precision levels, thereby enhancing memory access efficiency during attention computations. The memory overhead of the bidirectional page table is minimal: for example, with a batch size of 128 on Llama-2-7B, which has 32 layers and 32 attention heads per layer, the total size of all bidirectional page tables is only 64 MB. By contrast, the KV cache for a single request occupies 2 GB.
5.3. Memory Management Workflow
In this section, we delve into LeanKV’s memory management workflow, demonstrating how the proposed data structures interact to efficiently support unified KV cache compression. Notably, this memory management workflow is executed once per inference step for all requests in the batch and across all layers in the LLM. This design choice ensures sufficient parallelism for efficient GPU execution and amortizes the associated GPU kernel launch overheads.
Figure 5. Memory management flow in the prompt phase.
Prompt Phase. Figure 5 illustrates the memory management workflow during the prompt phase using an example request with eight tokens. In this setting, one high-precision page can store two tokens, and one low-precision page can store four. Since the exact number of high- and low-precision pages required by each head is unknown prior to running the compression algorithm, we conservatively allocate four unified pages per head, assuming that all tokens will be stored at high precision. Each head then independently applies the KV cache compression algorithm to determine which tokens to store at high-precision, low-precision, or to prune. This step constitutes the planning phase, during which each head calculates its specific memory needs based on its unique sparsity pattern. In this example, the first head uses one high-precision page, while the second head uses two; both heads also require one low-precision page. High-precision pages are allocated from the left side of the bidirectional page table, while low-precision pages are allocated from the right side. Following the planning phase, any unused pages are marked for recycling, triggering the coordination phase. Here, we employ a parallel prefix-sum-based approach, as described in the previous section, to facilitate efficient and conflict-free page recycling across heads.
Generation Phase. At each generation step, a head allocates a new page only if either its high-precision or low-precision pages are full, requiring at most one additional page per step. Each head independently checks its page availability and, if needed, allocates a new page in parallel using the prefix-sum-based approach. Unlike in the prompt phase, page recycling is not performed during generation, as the total number of stored tokens either remains the same if an old token is evicted or increases by one if no eviction occurs. Once a request is finished, all pages allocated for that request are recycled, freeing up memory for incoming requests.
6. Implementation
We implement LeanKV on top of vLLM, comprising 4.5K lines of CUDA/C++ code and 9K lines of Python. We first outline the architecture of LeanKV and then detail our custom GPU attention kernel, designed to efficiently support the unified cache compression algorithm.
6.1. LeanKV Architecture
The architecture of LeanKV is illustrated in Figure 6. At each inference step, the scheduler batches as many requests as possible within the available GPU memory to maximize throughput, sending the selected requests to all workers. Each GPU hosts one worker, responsible for executing a segment of the model. LeanKV utilizes tensor parallelism (shoeybi2019megatron, ), where each worker handles a partition of each matrix multiplication and a subset of the attention heads. Each worker includes a dedicated memory manager to oversee the KV cache for its assigned attention heads, using the method described in Section 5. Due to dynamic sparsity patterns, memory requirements vary per head for each request, leading to an irregular memory layout. Additionally, each worker incorporates an execution engine for model computation. To support unified KV cache compression, the execution engine integrates a KV compressor and a custom GPU attention kernel. After computing key and value vectors, the KV compressor is invoked to compress them using the algorithm in Section 4, storing the results in the KV cache. The custom GPU attention kernel then efficiently computes the attention output using the compressed KV cache.
Figure 6. LeanKV architecture.
6.2. Efficient Attention Kernel
We develop a custom GPU attention kernel that efficiently supports per-head dynamic sparsity and Hetero-KV quantization, fully harnessing the reduction in memory access volume and translating it into accelerated performance.
Overall, we assign each CUDA thread block to process a single attention head per sequence. However, mixed-precision quantization introduces potential load imbalance, as high-precision tokens demand more bandwidth, which is the bottleneck of attention, than low-precision ones. Thanks to the unified paging mechanism that stores tokens of the same precision together in a page, we mitigate this issue by having thread warps iterate over high-precision pages first, followed by low-precision pages, ensuring that threads within a warp remain load-balanced. Each page is processed in two phases: dot product between the query and keys to derive attention scores, and weighted sum of values. These two phases follow different computation patterns: reduction over the feature dimension and token dimension, respectively. To ensure coalesced and vectorized memory accesses for both phases, we design tailored data layouts and parallelization strategies, elaborated in the following paragraphs.
For key processing, each warp handles one page at a time, with threads in the warp divided into groups responsible for distinct keys. Within a group, each thread fetches its assigned keys in a vectorized manner, performs dequantization on the fly, and computes a partial dot product between the keys and corresponding query. Using a straightforward layout such as [$F$, $N_{\text{tokens}}$] for keys in each page, where $F$ denotes the feature dimension length, would cause threads within a group to fetch non-contiguous elements strided by $N_{\text{tokens}}$ when parallelizing across the feature dimension. Alternatively, using [$N_{\text{tokens}}$, $F$] would cause thread groups to fetch non-contiguous elements strided by $F$ when parallelizing across tokens. Both layouts suffer inefficient strided memory accesses, leading to low bandwidth utilization. Instead, we organize the layout of keys as [ $\frac{F}{K_{\text{vec}}\times K_{\text{group}}}$, $N_{\text{tokens}}$, $K_{\text{group}}$, $K_{\text{vec}}$], where $K_{\text{vec}}$ denotes the vectorization factor and $K_{\text{group}}$ denotes the number of threads per group. During each execution step, each thread fetches $K_{\text{vec}}$ consecutive elements, threads within a group collectively fetch $K_{\text{group}}$ adjacent chunks, and groups within the warp access contiguous chunks along the $N_{\text{tokens}}$ dimension. As a result, the combined memory accesses of all threads in a warp are contiguous, enabling memory coalescing and maximizing bandwidth utilization.
For value processing, tokens in a page are evenly distributed across thread groups in a warp; within each group, individual threads perform sum reductions over specific ranges of the feature dimension and save their accumulation results in registers. Once all groups finish, a tree reduction aggregates these partial results and produces the output. The number of registers required per thread is proportional to the feature dimension range assigned to it, while during key processing each thread only requires a single register to store the partial dot product. As a consequence, vectorization along the feature dimension, as used in key processing, is not suitable for value processing, because it would significantly increase register pressure by forcing each thread to handle a larger portion of the feature dimension. To better align with the computation pattern of reduction across tokens during value processing, we apply vectorization to the token dimension instead, which reduces register pressure and ensures more effective parallelization. We organize the layout of values as [ $\frac{F}{V_{\text{group}}}$, $\frac{N_{\text{tokens}}}{V_{\text{vec}}}$, $V_{\text{group}}$, $V_{\text{vec}}$] accordingly.
7. Evaluation
We first evaluate the effectiveness of LeanKV’s unified KV cache compression algorithm and analyze the sensitivity of different models to KV cache compression. Then, We demonstrate the improvements in throughput achieved by LeanKV.
7.1. Experiment Setup
We evaluate LeanKV on six models across three families: Llama2-7B and 70B (touvron2023llama, ), Llama3-8B and 70B (dubey2024llama, ), Mistral-7B (jiang2023mistral, ), and Mixtral-8x7B (jiang2024mixtral, ). Mixtral-8x7B employs a Mixture-of-Experts (MOE) architecture, while the remaining models are dense. Performance of LeanKV is evaluated on five benchmarks, grouped into two categories. The first includes tasks requiring advanced reasoning and logic: GSM8K (cobbe2021training, ) (mathematics), HumanEval (chen2021evaluating, ) (coding), and MMLU (hendrycks2020measuring, ) (academic level question answering). The second category includes traditional text processing tasks: SQuAD (rajpurkar2018know, ) (reading comprehension) and CNN-DailyMail (nallapati2016abstractive, ) (summarization). Throughput and latency is measured on Nvidia L40 GPUs, each with 48GB of memory (L40, ).
7.2. Evaluating Unified Cache Compression
We first demonstrate the effectiveness of Hetero-KV quantization and dynamic sparsity individually, and then evaluate the end-to-end benefits of the proposed unified KV cache compression. To mitigate the noise and randomness in floating-point computations, each experiment is repeated five times with the dataset randomly shuffled, and the reported results are the average of these runs.
Evaluating Hetero-KV. Figure 7 shows the performance of Hetero-KV quantization, namely K8V4 and K4V2, compared to FP16 baselines, on GSM8K and SQuAD, representing benchmarks from the two categories. To validate our intuition that key vectors have a more significant impact than value vectors (Section 3.1), we additionally evaluate the mirror configurations of Hetero-KV, namely K4V8 and K2V4, where values are stored with higher precision than keys.
K8V4 maintains the accuracy of the FP16 baseline across both benchmarks for all models, whereas its mirror configuration, K4V8, results in noticeable degradation, particularly on the reasoning-intensive GSM8K benchmark. The performance gap between K4V2 and K2V4 is even more pronounced. On GSM8K, K4V2 retains at least 77.8% of FP16 performance, while K2V4 reduces accuracy to nearly zero for all models. On SQuAD, K4V2 achieves 94.7% to 99.4% of FP16 performance, whereas K2V4 achieves only 28.2% to 60.0%. These results confirm that key vectors play a more critical role than value vectors, and demonstrate that Hetero-KV is effective in maintaining the accuracy of the FP16 baseline.
Figure 7. Performance of Hetero-KV.
Evaluating Dynamic Sparsity. We evaluate the effectiveness of per-head dynamic sparsity (Section 3.2) in identifying critical tokens, comparing it to the static sparsity method used in H2O (zhang2024h2o, ), which allocates an equal memory budget to all attention heads. The results are shown in Figure 8, where we use the Mistral-7B model as an example and all tokens are stored in FP16.
For reasoning-intensive workloads such as GSM8K and HumanEval, dynamic sparsity significantly outperforms static sparsity, with the performance gap widening as the fraction of pruned tokens increases. For traditional text processing tasks, however, the benefits of dynamic sparsity are less pronounced: it slightly outperforms static sparsity on SQuAD and performs similarly on CNN-DailyMail.
The difference between SQuAD and CNN-DailyMail can be attributed to their differing sensitivities to token pruning. SQuAD (reading comprehension) requires precise retrieval of specific words from the prompt, making it more sensitive to pruning critical tokens. In contrast, CNN-DailyMail (summarization) inherently involves discarding less relevant information to generate concise summaries, resulting in lower sensitivity to token pruning. To summarize, dynamic sparsity is superior to static sparsity for reasoning-intensive tasks and text processing that involves precise information retrieval.
Figure 8. Performance of dynamic vs. static sparsity.
Evaluating Unified KV Cache Compression. Finally, we evaluate the end-to-end performance of the unified KV cache compression described in Section 4. For each benchmark, we perform grid search over the high-precision threshold $T_{h}$ (with discrete values from {$0.5\%$, $1\%$, $5\%$, $10\%$, $20\%$}) and the low-precision threshold $T_{l}$ (with discrete values from $0\%$, $0.1\%$, $0.5\%$, $1\%$}), selecting the configurations that minimize the KV cache memory budget under different performance constraints. To account for the inherent randomness in floating-point computations, we consider a configuration to be performance-neutral if its average performance falls within three standard deviations of the FP16 baseline.
Table 1 reports the KV cache memory usage of LeanKV, normalized to vLLM (kwon2023efficient, ), across all evaluated models and benchmarks. The results are presented for two configurations: performance-neutral, which ensures no degradation in output quality, and less-than-$5\%$ performance degradation, a common target in prior work (lee2024infinigen, ; fu2024moa, ; zhao2024atom, ; lin2024qserve, ; lin2024awq, ; frantar2023sparsegpt, ) that enables more aggressive memory savings while preserving acceptable quality. Across models and workloads, LeanKV achieves performance neutral with $18.9\%$ to $28.9\%$ of the KV cache memory budget, and less-than-$5\%$ performance degradation with $7.6\%$ to $20.6\%$ of the memory budget.
Additionally, Figure 9 shows the Pareto performance-memory curves of LeanKV on the reasoning-intensive benchmark GSM8K and text processing benchmark SQuAD. The results are compared against the FP16 baseline, conventional 8-bit quantization (K8V8) and 4-bit quantization (K4V4), and the pruning-only technique H2O (zhang2024h2o, ). For GSM8K, K8V8 preserves the FP16 baseline performance, while K4V4 leads to a noticeable performance degradation of $2.8\%$ to $7.3\%$. In comparison, LeanKV achieves FP16-equivalent performance with only $19.8\%$ to $23.1\%$ of the memory budget, even less than K4V4, which requires $25\%$ memory. Moreover, LeanKV offers a more flexible memory-performance tradeoff, with performance degradation limited to less than $5\%$ while using just $9.4\%$ to $16.3\%$ of the memory. In contrast, under similar memory constraints, H2O results in near-zero accuracy. Similar trends are observed on the SQuAD benchmark: LeanKV matches the performance of FP16 and K8V8 with less memory than K4V4, which causes noticeable performance degradation, and outperforms H2O under comparable memory budgets.
Notably, across all evaluated models and benchmarks, LeanKV achieves performance-neutral with the K8V4-K4V2 configuration, where critical tokens are stored with K8V4 precision, non-critical tokens with K4V2, and no pruning is applied. For the less-than-$5\%$ degradation setting, the optimal configuration varies across models. For large 70B models, less-than-$5\%$ degradation is achieved with the more aggressive K4V2-K4V1 quantization without pruning, as pruning significantly degrades performance even on top of K8V4-K4V2. In contrast, for smaller models, less-than-$5\%$ degradation is achieved by further pruning the least important tokens on top of K8V4-K4V2. These varying configurations in low-memory regimes are also visualized in Figure 9, where each data point for LeanKV is annotated with its corresponding configuration. This disparity underscores the varying sensitivity of models to KV cache compression techniques, emphasizing the need for a unified framework capable of adapting to diverse scenarios to maximize efficiency.
In summary, LeanKV provides a superior performance-memory tradeoff compared to conventional methods. It maintains the performance of FP16 and 8-bit quantization while using even less memory than 4-bit quantization. Additionally, LeanKV offers a smooth tradeoff in the low-memory regime ($10\%$ to $20\%$ of memory), where pruning-only methods experience significant performance degradation.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 9. KV cache memory normalized to vLLM vs. benchmark performance tradeoff of LeanKV
Table 1. Memory usage of LeanKV normalized to vLLM across models and benchmarks.
Models Benchmarks GSM8K MMLU HumanEval SQUAD CNN-DailyMail Llama2-7B perf-neutral 18.9% 22.8% 23.1% 21.2% 21.6% 5% perf-loss 11.3% 11.5% 18.9% 14.6% 7.6% Llama2-70B perf-neutral 21.3% 21.4% 21.2% 21.1% 20.1% 5% perf-loss 19.5% 15.9% 18.2% 16.0% 15.7% Llama3-8B perf-neutral 20.7% 26.7% 26.6% 24.1% 20.0% 5% perf-loss 9.4% 20.6% 22.1% 12.8% 8.8% Llama3-70B perf-neutral 23.1% 21.5% 22.7% 22.7% 19.5% 5% perf-loss 16.3% 16.1% 21.2% 16.0% 15.7% Mistral-7B perf-neutral 26.8% 28.9% 23.0% 26.2% 19.4% 5% perf-loss 17.7% 13.6% 15.1% 15.9% 9.7% Mixtral-8x7B perf-neutral 19.8% 20.1% 21.4% 21.1% 20.0% 5% perf-loss 13.8% 8.9% 10.7% 8.9% 10.0%
7.3. Sensitivity to KV Cache Compression
We investigate the varying sensitivity of models to KV cache compression, using the Llama model family as an example, due to their similar training data and methodologies. Figure 10 illustrates the normalized performance degradation relative to the FP16 baseline for Llama2-7B, Llama3-8B, and Llama3-70B as the fraction of pruned tokens increases. Among the models, Llama2-7B exhibits the least performance degradation, even when compared to Llama3-8B, which has a comparable parameter size. This resilience can be attributed to Llama2-7B’s MHA architecture, which introduces greater redundancy in the KV cache compared to the more memory efficient GQA architecture utilized by Llama3-8B and Llama3-70B. Llama3-70B experiences the most significant performance degradation, suggesting that larger models distribute attention more evenly across tokens, making them more sensitive to pruning.
Figure 11 further illustrates the normalized performance degradation of the K4V2-K4V1 configuration as the fraction of tokens stored in K4V1 increases. The Llama3-70B model exhibits the greatest resilience to low-bit quantization, maintaining near-baseline performance even when 60% of tokens are stored in K4V1. In comparison, smaller models show noticeable performance degradation even when all tokens are stored in K4V2, highlighting their greater sensitivity to aggressive quantization.
These observations justify the two key trends outlined earlier: larger models and newer models employing more memory-efficient network architectures are more sensitive to pruning, while smaller models are more adversely impacted by low-bit quantization.
Figure 10. Sensitivity to KV cache pruning.
Figure 11. Sensitivity to low-bit KV cache quantization.
7.4. Evaluating Latency and Throughput
In this section, we provide a comprehensive evaluation of LeanKV’s performance, comparing it against state-of-the-art systems including vLLM and H2O across various models and configurations. All experiments were conducted on a machine equipped with eight NVIDIA L40 GPUs. For models that fit into a single GPU’s memory, such as Llama2-7B and Llama3-7B, we use one GPU. For larger models like Llama2-70B and Llama3-70B, we parallelize the computation across four GPUs. Our evaluation focuses on latency breakdown, attention kernel speedup, and end-to-end throughput to demonstrate the efficiency and scalability of LeanKV.
Memory Management Overhead. Figure 12 shows the latency breakdown of LeanKV during one inference step with a sequence length of 1024 tokens, evaluated across different batch sizes. The memory management overhead is remarkably low, contributing less than 0.2% of the total latency in the prompt phase and under 0.9% in the generation phase. Model execution dominates the latency in both phases, accounting for 96–97% of the latency in the prompt phase and 92–93% in the generation phase. Notably, as the batch size increases, the percentage of time spent on model execution rises slightly, reflecting the scalability of LeanKV. For comparison, in vLLM, model execution contributes 97% of the latency in the prompt phase and 95% in the generation phase. These results demonstrate that LeanKV effectively minimizes latency contributions from memory management and KV compression, ensuring that these auxiliary processes do not become performance bottlenecks.
Figure 12. Latency breakdown of LeanKV.
Attention Kernel Speedup. Figure 13 shows the speedup of the LeanKV’s custom attention kernel against vLLM under different quantization configurations. LeanKV achieves a near-linear speedup proportional to the reduction in KV cache size. For example, with K8V8, which halves the KV cache size relative to FP16 datatype, the theoretical speedup is 2$\times$, and LeanKV achieves 1.7$\times$. The slight gap is primarily due to the overhead of accessing quantization metadata and performing dequantization. Additionally, LeanKV achieves greater speedups on longer sequences, indicating that its bandwidth optimization techniques become increasingly effective as the sequence length grows.
Figure 13. Attention kernel speedup.
End-to-End Throughput. Figure 14 presents the end-to-end throughput and achieved batch sizes of LeanKV compared to vLLM and H2O. We conduct experiments on 1000 sequences sampled from the ShareGPT (ShareGPT, ) dataset, setting the generation length to 4096 tokens. The compression thresholds for LeanKV are selected based on the most challenging dataset, MMLU, making them the most conservative choices. In configurations with no accuracy loss, LeanKV achieves throughput speedups ranging from 2.0×2.0$\times$2.0 × to 2.5×2.5$\times$2.5 × compared to vLLM across various models. When allowing a 5% accuracy loss, LeanKV achieves even higher speedups, ranging from 3.0×3.0$\times$3.0 × to 6.9×6.9$\times$6.9 ×. These improvements are directly correlated with the larger batch sizes supported by LeanKV, enabled by its KV cache compression techniques. For instance, with the Llama2-7B model, LeanKV sustains a batch size of 62 without accuracy loss and 177 with 5% accuracy loss, far exceeding vLLM’s batch size of 22. The increased batch size enables LeanKV to maximize the utilization of GPU compute capacity, leading to higher throughput. Llama3-7B exhibits a less significant speedup compared to Llama2-7B because it utilizes GQA, which reduces its KV cache size by 4$\times$. As a result, vLLM already supports a relatively large batch size of 51.5 for Llama3-7B, narrowing the batch size advantage that LeanKV can leverage. Nevertheless, LeanKV still achieves up to 3.4$\times$ higher throughput over vLLM on Llama3-7B.
Notably, H2O supports even smaller batch sizes than vLLM and consequently has lower throughput. This is because H2O consumes additional memory during the prompt phase for intermediate results required to calculate compression metrics, limiting the batch size that can be supported. In contrast, we design a FlashAttention-compatible kernel using Triton, which avoids the additional memory overhead caused by intermediate results and achieves performance on par with vLLM.
Figure 14. Throughput and achieved batch size.
8. Conclusion
We propose LeanKV, a unified KV cache compression framework designed to enhance LLM serving efficiency through three key techniques: Hetero-KV, per-head dynamic sparsity, and unified KV compression. To support these innovations, LeanKV incorporates system optimizations including unified paging and on-GPU parallel memory management. Our experiments show that LeanKV achieves 1.9×1.9$\times$1.9 × to 2.5×2.5$\times$2.5 × higher throughput over state-of-the-art systems without compromising model accuracy, and up to 6.9×6.9$\times$6.9 × higher throughput with minimal accuracy degradation.

### v2

@@
 Unifying KV Cache Compression for Large Language Models with LeanKV
-Abstract.
-Large language models (LLMs) demonstrate exceptional performance but incur high serving costs due to substantial memory demands, with the key-value (KV) cache being a primary bottleneck.
-Existing KV cache compression methods, including quantization and pruning, struggle with limitations such as uniform treatment of keys and values and static memory allocation across attention heads.
-To address these challenges, we introduce LeanKV, a unified KV cache compression framework that enhances LLM serving efficiency without compromising accuracy through three innovations: (1) Hetero-KV quantization, which stores keys at a higher precision than values to reflect their greater impact on attention computations; (2) per-head dynamic sparsity, which allocates memory based on token importance per head and per request; and (3) unified KV compression, integrating mixed-precision quantization and selective pruning to enable a smooth tradeoff between model accuracy and memory efficiency.
-To efficiently support these techniques, LeanKV introduces systems optimizations including unified paging and on-GPU parallel memory management.
-Implemented on vLLM, LeanKV compresses the KV cache by 3.0×3.0$\times$3.0 × to 5.0×5.0$\times$5.0 × without accuracy loss and up to 11.0×11.0$\times$11.0 × with under 5% accuracy loss, enhancing throughput by 1.9×1.9$\times$1.9 × to 2.5×2.5$\times$2.5 ×, and up to 6.9×6.9$\times$6.9 ×.
-*Equal contribution.
-1. Introduction
-Large language models (LLMs) like GPT (brown2020language, ; achiam2023gpt, ; ouyang2022training, ) and Gemini (team2023gemini, ) have demonstrated significant potential to impact our daily lives, offering promising applications in areas including chatbots (vicuna, ; chiang2024chatbot, ), programming (gitcopilot, ; guo2024deepseek, ), mathematics (shao2024deepseekmath, ; zhong2024evaluation, ) and medical assistance (saab2024capabilities, ; xie2024preliminary, ).
+Abstract
+Large language models (LLMs) exhibit exceptional performance but incur significant serving costs due to their substantial memory requirements, with the key-value (KV) cache being a primary bottleneck.
+Existing KV cache compression techniques, such as quantization and pruning, apply uniform treatment to both keys and values, and discard unimportant tokens entirely, overlooking the fine-grained differences in significance of various components within the KV cache.
+To address these limitations, we introduce LeanKV, a framework that advances KV cache compression by exploiting three levels of differentiation in the KV cache: (1) the differing impact of keys and values on attention computation, (2) the varying importance of tokens, and (3) the diverse dynamic sparsity patterns across attention heads.
+At the core of LeanKV is an on-GPU memory manager that compacts fragmented free memory list into contiguous regions in parallel, effectively translating sparsity in the KV cache into performance gains.
+We evaluate LeanKV on several mainstream models, including the recent "thinking model".
+LeanKV is able to compress the KV cache by 2.7×2.7$\times$2.7 × to 5.7×5.7$\times$5.7 × with near-lossless accuracy on complex workloads requiring sophisticated reasoning and long-generation capabilities, and enhances throughput by 1.9×1.9$\times$1.9 × to 5.4×5.4$\times$5.4 ×.
+1 Introduction
+Large language models (LLMs) like GPT [9, 4, 46] and Gemini [57] have demonstrated significant potential to impact our daily lives, offering promising applications in areas including chatbots [59, 12], programming [21, 24], mathematics [50, 71] and medical assistance [49, 65].
 Despite their exceptional performance, hosting LLMs is costly due to their large model sizes, demanding extensive hardware resources.
-Given the pervasive adoption of LLMs, enhancing serving efficiency has become critically important (aminabadi2022deepspeed, ; fang2021turbotransformers, ; kao2023flat, ; shi2023welder, ; zhang2023shepherd, ; han2022microsecond, ; pope2023efficiently, ; li2023alpaserve, ).
-LLMs are fundamentally based on the Transformer architecture (vaswani2017attention, ; wolf2020transformers, ), which employs the attention mechanism to generate tokens sequentially, based on the input prompt and previously generated tokens.
+Given the pervasive adoption of LLMs, enhancing serving efficiency has become critically important [6, 18, 31, 53, 68, 25, 47, 36].
+LLMs are fundamentally based on the Transformer architecture [61, 62], which employs the attention mechanism to generate tokens sequentially, based on the input prompt and previously generated tokens.
 Within the attention mechanism, each token is transformed into query, key and value feature vectors.
@@
 Subsequently, the output is computed as the weighted sum of the value vectors, based on the attention scores.
-To capture diverse relationships between tokens, the attention mechanism is typically divided into multiple heads (vaswani2017attention, ; shazeer2019fast, ; ainslie2023gqa, ).
-Each head independently computes attention, and their outputs are concatenated and transformed to create a richer representation of the input data
-To avoid redundant computation, LLM inference frameworks typically cache intermediate key and value tensors in memory, commonly referred to as the KV cache (kwon2023efficient, ).
-The size of the KV cache scales linearly with both sequence length and the number of concurrent requests, often comprising over 90% of total memory consumption (kwon2023efficient, ; yu2022orca, ; gao2018low, ; yang2022infless, ).
-As state-of-the-art models continue to support longer sequences (dubey2024llama, ; jiang2024mixtral, ; liu2024deepseek, ; achiam2023gpt, ; team2023gemini, ; touvron2023llama, ), the KV cache has emerged as a critical bottleneck for LLM serving efficiency.
-It limits the number of concurrent requests that can fit in memory and increases attention computation latency due to its memory bandwidth-bound nature (dao2022flashattention, ; dao2023flashattention, ).
-Recognizing the critical role of the KV cache in limiting serving efficiency, researchers have investigated various compression techniques, primarily focused on quantization (lin2024qserve, ; flashinfer, ; hooper2024kvquant, ) and pruning (zhang2024h2o, ; xiao2023efficient, ; zhang2024pyramidkv, ; ge2023model, ; li2024snapkv, ).
+To capture diverse relationships between tokens, the attention mechanism is typically divided into multiple heads [61, 51, 5].
+Each head independently computes attention, and their outputs are concatenated and transformed to create a richer representation of the input data.
+To avoid redundant computation, LLM inference frameworks typically cache intermediate key and value tensors in memory, commonly referred to as the KV cache [32].
+The size of the KV cache scales linearly with both sequence length and the number of concurrent requests, often comprising over 90% of total memory consumption [32, 67, 19, 66].
+As state-of-the-art models continue to support longer sequences [17, 30, 38, 4, 57, 60], and with the rise of recent "thinking models" [23, 29, 58] that generate extended reasoning processes, the KV cache has emerged as a critical bottleneck for LLM serving efficiency.
+It limits the number of concurrent requests that can fit in memory and increases attention computation latency due to its memory bandwidth-bound nature [15, 14].
+Recognizing the critical role of the KV cache in limiting serving efficiency, researchers have investigated various compression techniques, primarily focused on quantization [37, 2, 27] and pruning [69, 64, 10, 20, 35].
 Quantization reduces the size of the KV cache by converting floating-point feature values into integers, representing them based on pre-defined intervals.
-Pruning, on the other hand, reduces the token sequence length by eliminating tokens deemed insignificant based on attention scores.
-Despite the promise of these techniques, both of them face limitations in efficiency and practicality.
-First and foremost, existing quantization methods apply the same treatment to key and value vectors of all tokens, disregarding the different roles of keys and values in attention computation, as well as the varying significance of different tokens.
-Second, current pruning methods statically allocate memory across attention heads, failing to account for the dynamic attention patterns across heads and requests.
+Pruning, on the other hand, reduces the token sequence length by eliminating unimportant tokens based on attention scores.
+While promising, these techniques overlook the fine-grained differentiation in the significance of various KV cache components, limiting their overall effectiveness.
+First and foremost, existing quantization methods save key and value vectors of all tokens at a uniform precision, disregarding the distinct roles of keys and values and the varying importance of tokens.
+Second, current pruning methods employ static memory allocation across attention heads, neglecting the dynamic and heterogeneous attention patterns observed across heads and requests.
 Moreover, these pruning methods have not been comprehensively evaluated on workloads requiring advanced reasoning, such as programming and mathematics.
-Lastly, prior work has largely overlooked the potential of unifying quantization and pruning to achieve even greater serving efficiency.
-In this paper, we propose LeanKV, a unified KV cache compression framework that addresses the limitations of existing methods and enhances LLM serving efficiency.
-LeanKV introduces three algorithmic principles :
+In this paper, we present LeanKV, a framework that advances KV cache compression by leveraging three levels of differentiation within the KV cache:
 1.
-Hetero-KV.
-LeanKV compresses key and value vectors using heterogeneous precisions, with key vectors stored at higher precision and value vectors at lower precision.
-This design leverages the observation that key vectors have a greater influence on attention calculations, as they affect the attention scores of all tokens, whereas value vectors only impact the contributions of individual tokens to the attention output.
+Differentiated Keys and Values.
+LeanKV assigns higher precision to key vectors than to value vectors, based on the observation that key vectors play a more significant role in attention computations.
 2.
+Token Importance Differentiation.
+Guided by attention scores, LeanKV stores tokens at varying precision levels based on their significance.
+Most critical tokens are quantized at higher precision, less critical tokens are quantized at lower precision, and the least significant tokens are pruned.
+This hierarchical approach enables a flexible tradeoff between model accuracy and memory efficiency.
+3.
 Per-head Dynamic Sparsity.
-LeanKV dynamically identifies critical tokens on a per-head and per-request basis, adjusting memory allocation accordingly.
-This design stems from the observation that attention score distributions exhibit dynamic sparsity across both heads and requests.
+Attention score distributions exhibit dynamic sparsity across both heads and requests.
 Within a single request, the number of critical tokens varies across heads, and for the same head, the number of critical tokens also varies across different requests.
+To address this variability, LeanKV dynamically identifies critical tokens per-head, per-request, adapting memory allocation accordingly.
+These differentiations wthin the KV cache pose new challenges in memory management.
+In LeanKV, each request requires varying amounts of memory for each attention head to store tokens at different precision levels.
+Moreover, key and value vectors of the same token are compressed at different precisions.
+These complexities significantly increase memory management overhead, compared to traditional attention mechanisms, where memory allocation is uniform across all heads.
+To address these challenges, LeanKV employs an on-GPU memory manager that efficiently handles the irregular memory usage patterns across heads and requests via parallel KV compaction.
+Rather than performing memory management operations sequentially on CPU for each head and request, parallel KV compaction optimizes memory allocation and recycling by packing fragmented free memory lists into contiguous regions directly on the GPU in parallel.
+Parallel KV compaction is supported by three key on-GPU data structures:
+1.
+Unified Pages.
+GPU memory is divided into fixed-size pages containing vectors stored at various precisions.
+Each page is dynamically parsed according to the precision requirements of the head and request, while an attention GPU kernel is co-designed with the page layout to maximize memory bandwidth utilization.
+2.
+Circular Free Page List.
+All page IDs are managed in a centralized, GPU-resident circular list.
+Both free and used pages are kept in contiguous regions, enabling efficient page allocation and recycling via parallel prefix sum.
 3.
-Unified KV Compression.
-Built on Hetero-KV quantization and per-head dynamic sparsity, LeanKV unifies KV compression strategies by applying mixed-precision quantization and selective pruning on a per-head and per-request basis, with token importance serving as a unified metric.
-Critical tokens are stored at higher precision, less important tokens are quantized at lower precision, and the least significant tokens are pruned.
-This integrated approach enables a smooth tradeoff between model accuracy and memory efficiency.
-Implementing LeanKV’s algorithmic innovations presents new challenges in system design, particularly regarding memory management.
-Unlike traditional attention mechanisms that adopt uniform memory allocation across attention heads, LeanKV requires dynamic, per-head memory allocation for each request, significantly increasing system complexity.
-Moreover, LeanKV must flexibly configure memory to accommodate varying numbers of tokens stored at different precision levels or pruned entirely, further complicating memory management.
-To address these challenges, LeanKV introduces two novel techniques.
-First, LeanKV employs unified paging, which divides memory into fixed-size pages containing vectors stored at various precisions.
-Each page is dynamically parsed based on the precision configuration required by the head and request, while an attention GPU kernel is co-designed with the page layout to maximize memory bandwidth utilization.
-Second, LeanKV introduces an on-GPU memory management mechanism that accelerates memory management by parallelizing it directly on the GPU with minimal metadata overhead.
-We implement LeanKV on vLLM (kwon2023efficient, ), a state-of-the-art LLM inference system, and evaluate it across three prominent model families using representative workloads, including tasks that demand advanced reasoning such as mathematics and coding.
+Bidirectional Page Table.
+Instead of maintaining separate page tables for high- and low-precision pages, LeanKV consolidates both into a single bi-directional page table to minimize metadata overhead.
+High precision page IDs grow from the left side of the table, and low-precision page IDs grow from the right side.
+We implement LeanKV on vLLM [32], a state-of-the-art LLM inference system, and evaluate it across three prominent model families including the emerging thinking model.
+The evaluation is conducted on representative workloads that demand advanced reasoning capabilities, such as mathematics and coding.
 Our experimental results demonstrate that LeanKV offers a superior cost-accuracy tradeoff compared to existing KV cache management methods.
-Specifically, LeanKV compresses the KV cache by 3.0×3.0$\times$3.0 × to 5.0×5.0$\times$5.0 × without loss in model accuracy, and by up to 11.0×11.0$\times$11.0 × with a marginal accuracy reduction of under 5%, resulting in throughput improvement of 1.9×1.9$\times$1.9 × to 2.5×2.5$\times$2.5 ×, and up to 6.9×6.9$\times$6.9 × correspondingly.
-2. Background and Related Work
+Specifically, LeanKV compresses the KV cache by 2.7×2.7$\times$2.7 × to 5.7×5.7$\times$5.7 × with near-lossless accuracy, resulting in throughput improvement of 1.9×1.9$\times$1.9 × to 5.4×5.4$\times$5.4 ×.
+The throughput improvement is primarily enabled by the parallel KV compaction mechanism, which reduces management overhead by three orders of magnitude, decreasing the proportion of memory management time in total execution latency from up to $76\%$ to less than $1\%$ compared to a sequential CPU-based implementation.
+2 Background and Related Work
 In this section, we describe Transformer-based large language models and review existing techniques for KV cache management aimed at improving LLM serving efficiency.
-2.1. Large Language Models
+2.1 Large Language Models
 Transformer Architecture.
-Large language models (LLMs) are predominantly built upon the Transformer architecture (vaswani2017attention, ), which is used to capture complex long-range dependencies in sequential data.
+Large language models (LLMs) are predominantly built upon the Transformer architecture [61], which is used to capture complex long-range dependencies in sequential data.
 At the core of Transformer architecture is the attention mechanism, which allows each token in a sequence to weigh the significance of other tokens when constructing its contextualized representation.
 Mathematically, the attention computation is defined as:
-(1) $\operatorname{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})_{i}$ $=\sum_{j=1}^{i}\operatorname{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)_{ij}\mathbf{v}_{j}$
+$\operatorname{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})_{i}$ $=\sum_{j=1}^{i}\operatorname{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)_{ij}\mathbf{v}_{j}$ (1)
 $=\sum_{j=1}^{i}\frac{\exp\left(\frac{\mathbf{q}_{i}\cdot\mathbf{k}_{j}}{\sqrt{d}}\right)}{\sum_{n=1}^{i}\exp\left(\frac{\mathbf{q}_{i}\cdot\mathbf{k}_{n}}{\sqrt{d}}\right)}\mathbf{v}_{j}$
@@
 The outputs from different heads are concatenated and linearly transformed to create a more expressive representation.
-Grouped-query attention (GQA) (ainslie2023gqa, ) improves the efficiency of MHA by allowing a group of query heads to share the same projection of keys and values, referred to as a KV head.
+Grouped-query attention (GQA) [5] improves the efficiency of MHA by allowing a group of query heads to share the same projection of keys and values, referred to as a KV head.
 Autoregressive Generation and KV Cache.
@@
 To avoid redundant computations across generation steps, KV cache is introduced to store the keys and values of all previous tokens.
-However, the size of the KV cache grows linearly with both the sequence length and batch size, quickly becoming a bottleneck for inference throughput (yu2022orca, ; kwon2023efficient, ).
+However, the size of the KV cache grows linearly with both the sequence length and batch size, quickly becoming a bottleneck for inference throughput [67, 32].
 Therefore, efficient KV cache management is critical for alleviating memory bottlenecks and improving LLM serving efficiency.
-2.2. KV Cache Optimization
+2.2 KV Cache Optimization
 PagedAttention.
-Static KV cache management systems (yu2022orca, ; FasterTransformer, ) reserve memory for the maximum possible sequence length, leading to considerable memory waste when actual sequences are shorter.
-To mitigate this inefficiency, vLLM (kwon2023efficient, ) introduces PagedAttention, which partitions the KV cache into pages, each containing a fixed number of tokens, and allocates the pages on-demand as the sequence length grows.
+Static KV cache management systems [67, 1] reserve memory for the maximum possible sequence length, leading to considerable memory waste when actual sequences are shorter.
+To mitigate this inefficiency, vLLM [32] introduces PagedAttention, which partitions the KV cache into pages, each containing a fixed number of tokens, and allocates the pages on-demand as the sequence length grows.
 By managing the KV cache at the granularity of pages, vLLM reduces memory waste and enables larger batch sizes to improve serving efficiency.
 KV Cache Quantization.
-Quantization (nagel2021white, ; dettmers2022gpt3, ; guo2023olive, ; hooper2024kvquant, ) reduce the KV cache size by approximating high-precision floating points with discrete low-bit values.
+Quantization [43, 16, 22, 27] reduce the KV cache size by approximating high-precision floating points with discrete low-bit values.
 To quantize a given tensor $\mathbf{X}$, we first calculate the scale $s$ and zero point $z$ as shown in Eq.
@@
 Unlike the quantized tensor $\mathbf{Q}$, the metadata $s$ and $z$ are kept in higher precision (e.g., FP16) to ensure more accurate dequantization.
-(2) $s=\frac{\mathbf{X}_{\text{max}}-\mathbf{X}_{\text{min}}}{2^{b}-1},\quad z=-\mathbf{X}_{\text{min}}$
-State-of-the-art KV cache quantization methods such as Atom (zhao2024atom, ) and Qserve (lin2024qserve, ) apply this process to each key and value vector independently.
+$s=\frac{\mathbf{X}_{\text{max}}-\mathbf{X}_{\text{min}}}{2^{b}-1},\quad z=-\mathbf{X}_{\text{min}}$ (2)
+State-of-the-art KV cache quantization methods such as Atom [70] and Qserve [37] apply this process to each key and value vector independently.
 However, these uniform quantization methods, which apply a fixed bit-width across keys and values of all tokens, may not be optimal.
@@
 KV Cache Pruning.
-Several recent works (zhang2024h2o, ; xiao2023efficient, ; li2024snapkv, ; zhang2024pyramidkv, ; liu2024scissorhands, ; ge2023model, ) have explored KV cache pruning to alleviate the memory bottleneck in LLM inference, which can be considered as an extreme case of quantization.
+Several recent works [69, 64, 35, 10, 40, 20] have explored KV cache pruning to alleviate the memory bottleneck in LLM inference, which can be considered as an extreme case of quantization.
 These methods use attention scores to assess token importance and evict less important tokens, thereby reducing the KV cache size.
 The effectiveness of these methods, however, is limited by their static and inflexible allocation of memory resources.
-Specifically, H2O (zhang2024h2o, ) and SnapKV (li2024snapkv, ) allocate the same memory budget uniformly across all attention heads and layers.
-PyramidKV (zhang2024pyramidkv, ) allocates more memory in lower layers and less in higher ones, based on the empirical observation that the number of important tokens shrinks as the model depth increases, but still relies on static heuristics and cannot dynamically adjust cache allocations to suit various workloads.
-Complementary to KV cache pruning, another line of work (sheng2023flexgen, ; lee2024infinigen, ) addresses GPU memory constraints by offloading the entire KV cache to CPU memory and selectively loading important tokens to the GPU for attention computations.
+Specifically, H2O [69] and SnapKV [35] allocate the same memory budget uniformly across all attention heads and layers.
+PyramidKV [10] allocates more memory in lower layers and less in higher ones, based on the empirical observation that the number of important tokens shrinks as the model depth increases, but still relies on static heuristics and cannot dynamically adjust cache allocations to suit various workloads.
+Complementary to KV cache pruning, another line of work [52, 33] addresses GPU memory constraints by offloading the entire KV cache to CPU memory and selectively loading important tokens to the GPU for attention computations.
 However, these methods do not inherently reduce the KV cache size and incur significant latency due to data transfers between the CPU and GPU.
-3. Motivation
-In this section, we present the key findings that underpin the design of LeanKV, as well as their implications on KV cache management.
-3.1. Impact of Key And Value
+3 The Case for Differentiated KV
+We present the case for three levels of differentiation within the KV cache, motivating the design of LeanKV.
+3.1 Differentiated Impacts of Keys and Values
 While the significance of different tokens can be directly inferred from attention scores, the roles of the key and value vectors within a token are less obvious.
@@
 This allows us to assess the relative importance of attention scores and value vectors by examining their contributions to these coefficients.
-(3) $\operatorname{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})_{i}$ $=\sum_{j=1}^{i}\underbrace{\operatorname{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)_{ij}|\mathbf{v}_{j}|}_{\text{Coefficient}}\underbrace{\frac{\mathbf{v}_{j}}{|\mathbf{v}_{j}|}}_{\text{Unit vector}}$
-Figure 1 presents the distributions of the average attention score per token and the norms of value vectors across three representative layers of the Llama3-8B model (dubey2024llama, ), evaluated on the first one thousand articles from the Wikitext dataset (merity2016pointer, ).
+$\operatorname{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})_{i}$ $=\sum_{j=1}^{i}\underbrace{\operatorname{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)_{ij}|\mathbf{v}_{j}|}_{\text{Coefficient}}\underbrace{\frac{\mathbf{v}_{j}}{|\mathbf{v}_{j}|}}_{\text{Unit vector}}$ (3)
+Figure 1 presents the distributions of the average attention score per token and the norms of value vectors across three representative layers of the Llama3-8B model [17], evaluated on the first one thousand articles from the Wikitext dataset [42].
 Notably, the attention scores span seven orders of magnitude, vastly exceeding the range of value vector norms, which only cover at most two orders of magnitude.
@@
 We therefore conclude that key vectors exert a broader and more impactful influence than value vectors, motivating further exploration of processing key and value vectors at distinct precision levels.
-Figure 1. Distribution of attention score and value vector norm in Llama3-8B.
-3.2. Dynamic Sparsity Patterns
-Existing approaches to KV cache pruning primarily rely on static memory allocation for each transformer layer in LLMs, either distributing memory uniformly (xiao2023efficient, ; zhang2024h2o, ; ge2023model, ; li2024snapkv, ) or allocating it based on offline profiling (zhang2024pyramidkv, ).
-However, LLMs exhibit dynamic sparsity patterns across both attention heads and individual requests, which static memory allocation cannot adequately handle.
+Figure 1: Distribution of attention score and value vector norm in Llama3-8B.
+3.2 Differentiated Token Importance
+Tokens contribute to attention outputs with varying degrees of importance, as reflected by their attention scores.
+By exploiting these differences, we can apply finer-grained compression strategies that go beyond uniformly quantizing all tokens [70, 37] or exclusively pruning the least important ones [69, 10].
+Specifically, we store progressively less important tokens at decreasing precision levels, ultimately leading to the pruning of the least important tokens.
+To investigate the feasibility of this hierarchical approach, we experiment with a simple design: quantizing both the key and value vectors of significant tokens to 8 bits while pruning insignificant ones.
+Figure 2 compares the performance of the FP16 baseline, 8-bit quantization, pruning, and the combination of 8-bit quantization with pruning on the general knowledge question answering benchmark MMLU [26] using the Llama3-8B model.
+For both pruning and the combined approach, approximately 50% of insignificant tokens are pruned.
+The resulting accuracy is nearly indistinguishable among the four approaches.
+However, the combination of 8-bit quantization and pruning achieves the lowest memory usage.
+The improved efficiency of this combined approach motivates us to further explore the spectrum of KV cache compression by introducing varying levels of precision to accommodate tokens of differing significance.
+Figure 2: Accuracy and normalized KV cache memory usage of the Llama3-8B model on MMLU
+3.3 Dynamic Sparsity Patterns
+Existing approaches to KV cache pruning primarily rely on static memory allocation for each transformer layer in LLMs, either distributing memory uniformly [64, 69, 20, 35] or allocating it based on offline profiling [10].
+However, LLMs exhibit dynamic sparsity patterns that vary across both attention heads and requests, which static memory allocation cannot adequately handle.
 To evaluate the sparsity of attention computation, we analyze the minimum number of critical tokens required to retain the majority of information, specifically by preserving a target percentage (e.g., 95%) of the total attention score.
@@
 These highly dynamic sparsity patterns, across layers, attention heads, and individual requests, underscore the necessity of adaptive memory management on a per-head and per-request basis.
-Figure 2. Number of critical tokens per layer in Llama3-8B to preserve 95% of the total attention score.
-Figure 3. Number of critical tokens per KV head in Llama3-8B to preserve 95% of the total attention score.
-3.3. Unifying Quantization and Pruning
-Quantization compresses the KV cache by representing feature values with a reduced bit width, whereas pruning can be viewed as an extreme form of quantization: effectively storing feature values with zero bits.
-By using attention scores to measure the significance of tokens, we can unify the spectrum of KV cache compression by storing progressively less critical tokens at decreasing precision levels, ultimately leading to the pruning of the least critical tokens.
-To investigate the feasibility of this unified approach, we experiment with a simple design: quantizing both the key and value vectors of significant tokens to 8 bits while pruning insignificant ones.
-Figure 4 compares the performance of the FP16 baseline, 8-bit quantization, pruning, and the combination of 8-bit quantization with pruning on the CNN/Daily Mail summarization dataset using the Llama3-8B model.
-For both pruning and the combined approach, approximately 20% of insignificant tokens are pruned.
-The ROUGE-1 score, which measures the fidelity of the LLM-generated responses to reference answers, is nearly indistinguishable among the four approaches.
-However, the combination of 8-bit quantization and pruning achieves the lowest memory usage.
-The improved efficiency of this combined approach motivates us to further explore the spectrum of KV cache compression by introducing varying levels of precision to accommodate tokens of differing significance.
-Figure 4. Rogue-1 score and normalized KV cache memory usage of the Llama3-8B model on CNN/DailyMail.
-3.4. Key Findings and Implications
-In this section, we outline the key findings that motivate the design of LeanKV.
-Specifically, we demonstrate that within individual tokens, key vectors have a greater impact than value vectors, inspiring us to propose storing key vectors at higher precision than value vectors.
-Beyond that, the significance of different tokens, as indicated by softmax scores, exhibits dynamic sparsity patterns across both heads and requests.
-These patterns necessitate flexible memory management on a per-head and per-request basis.
-Lastly, we show that unifying quantization and pruning can outperform either approach when used in isolation.
-These findings motivate us to further explore the spectrum of KV cache compression, on a per-head and per-request basis to achieve the maximum memory efficiency.
-Specifically, each attention head should be allowed to store tokens at varying precision levels and to prune the least important ones, according to the sparsity patterns of individual requests.
-Additionally, within each token, the key and value vectors should be stored in different precisions.
-However, this unified approach presents significant challenges on memory management systems due to the resulting irregular memory layout, where the KV cache of each head and request consists of varying number of tokens stored at different precisions.
-4. Unified Compression Algorithm
-In this section, we present the design of LeanKV’s unified compression algorithm, which leverages the insights from Section 3 to optimize memory efficiency.
-To account for the greater impact of key vectors compared to value vectors, we propose Hetero-KV, a strategy that quantizes keys at a higher precision than values.
-For instance, key and value vectors can be quantized to 8 and 4 bits (K8V4), or to 4 and 2 bits (K4V2).
-Additionally, we propose per-head memory management tailored to dynamic sparsity patterns.
-Rather than imposing a fixed memory budget, each head adaptively determines its memory allocation by quantizing tokens at various precision levels or pruning, based on the attention score distributions of individual requests.
-Specifically, tokens are categorized into three significance classes: the most important tokens are quantized at high precision (e.g., K8V4), moderately important tokens at low precision (e.g., K4V2), and the least important tokens are pruned.
-Next, we describe the unified compression algorithm in greater detail, addressing the prompt phase and the generation phase separately.
-Notably, in both phases, compression is conducted on a per-head and per-request basis.
+Figure 3: Number of critical tokens per layer in Llama3-8B to preserve 95% of the total attention score.
+Figure 4: Number of critical tokens per KV head in Llama3-8B to preserve 95% of the total attention score.
+3.4 Key Findings and Implications
+Our findings reveal three critical levels of differentiation within the KV cache:
+  • Keys exert a broader impact on attention computation than values, motivating differentiated precisions for keys and values.
+  • Tokens vary in importance as reflected by their attention scores, motivating a fine-grained, hierarchical compression strategy.
+  • Attention sparsity patterns vary across requests and attention heads, requiring dynamically managing memory resources on a per-request and per-head basis.
+Collectively, these insights guide the design of LeanKV, which integrates differentiated precision assignments for keys and values, hierarchical token importance handling, and per-request and per-head memory management.
+4 KV Compression Policy
+We present the KV compression policy of LeanKV that addresses the three levels of differentiation within the KV cache.
+First, to reflect the greater impact of keys on attention computation than values, we propose quantizing keys at a higher precision than values.
+For example, keys and values can be quantized to 8 and 4 bits (K8V4), or 4 and 2 bits (K4V2).
+Second, to account for the varying importance of tokens, we introduce a hierarchical compression strategy that classifies tokens into three significance levels: the most important tokens are quantized at high precision (e.g., K8V4), moderately important tokens at lower precision (e.g., K4V2), and the least important tokens are pruned.
+Finally, to address the dynamic sparsity patterns across requests and attention heads, we propose an adaptive memory management approach.
+Rather than imposing a fixed memory budget, our approach allows each attention head to determine its memory requirements dynamically based on its specific sparsity pattern.
+Next, we describe the KV compression policy in greater detail, addressing the prompt phase and the generation phase separately.
+In both phases, compression is applied per-request and per-head, ensuring that memory usage is tailored to the specific sparsity pattern of each request and attention head.
 Prompt Phase.
 In the prompt phase, key and value vectors for all tokens in the prompt are computed.
-The compression algorithm then determines the appropriate precision for storing each token based on its significance.
+The compression policy then determines the appropriate precision for storing each token based on its significance.
 The significance of the $i^{th}$ token is calculated by averaging the $N-i$ attention scores it receives from subsequent tokens, where $N$ denotes the sequence length.
-This average is then normalized to ensure that the total significance across all tokens sums to one.
-To prevent premature compression, the algorithm reserves a recent window of size $W$, where all tokens are quantized at a high precision.
-$W$ is predefined and typically set to 64.
-The remaining tokens are then sorted by significance.
-The least important $N_{p}$ tokens, whose cumulative significance is below the threshold $T_{l}$, are pruned.
-The next $N_{l}$ tokens, with a cumulative significance less than $T_{h}$, are quantized to low precision, while the most important $N_{h}$ tokens are quantized to high precision.
-Here, $N_{p}+N_{l}+N_{h}+W=N$, and $T_{l}$ and $T_{h}$ are predefined thresholds determined by offline profiling.
-As a result of this mixed-precision compression, the KV cache is conceptually divided into two parts: a high precision section $KV_{h}$ and a low precision section $KV_{l}$.
+In the case of GQA and MHA, scores from all attention heads associated with the KV head are aggregated using the maximum operation.
+To mitigate premature compression, the most recent $W$ tokens are always quantized at high precision, where $W$ is typically set to 64.
+For the remaining tokens, the policy determines the precision level of the $i^{th}$ token by comparing its significance score to the theoretical average $\frac{1}{i}$, based on the intuition that tokens with scores below average are less critical.
+Specifically, the $i^{th}$ token is quantized at high precision if its significance score exceeds $\frac{\alpha_{h}}{i}$, at low precision if its score lies within the interval $[\frac{\alpha_{l}}{i},\frac{\alpha_{h}}{i}]$, and is pruned otherwise.
+The parameters $\alpha_{l}$ and $\alpha_{h}$, which define the thresholds for high- and low-precision quantization, are determined offline through profiling on a calibration dataset.
+As a result of this hierarchical compression, the KV cache is conceptually divided into two parts: a high-precision section $KV_{h}$ and a low-precision section $KV_{l}$.
 Generation Phase.
@@
 Essentially, this policy establishes a smooth downgrading path for the less important token: rather than being pruned directly, it is first re-quantized to low precision, with pruning occurring only if it remains insignificant.
-The detailed procedure is outlined in Algorithm 1 and is governed by two predefined thresholds: the high precision threshold $T_{h}$ and the low precision threshold $T_{l}$.
-If the significance of $t_{c}$ is at least $T_{h}$, it is quantized at high precision and added to $KV_{h}$, the high precision section of the KV cache (line 6).
+The detailed procedure is outlined in Algorithm 1, following the same intuition and parameters as the prompt phase.
+Given the sequence length $N$, $t_{c}$ is quantized at high precision and added to $KV_{h}$, the high precision KV cache (line 6), if its significance exceeds $\frac{\alpha_{h}}{N}$.
 Subsequently, the least significant token in $KV_{h}$ is identified as the victim token $t_{v}$ for more aggressive compression (line 7).
-If $t_{v}$’s significance exceeds $T_{h}$, it remains in $KV_{h}$; if it falls between $T_{h}$ and $T_{l}$, $t_{v}$ is re-quantized to low precision and moved to $KV_{l}$, the low precision section of the KV cache (line 9); otherwise, $t_{v}$ is pruned.
-Similarly, if $t_{c}$’s significance falls between $T_{h}$ and $T_{l}$, it is quantized to low precision and added to $KV_{l}$ (line 14).
-The least important token in $KV_{l}$ is designated as the victim $t_{v}$ (line 15), and is further pruned if its significance falls below $T_{l}$ (line 17).
-Notably, the proposed unified compression algorithm is extensible beyond three precision levels, and we adopt three precision levels to minimize metadata overhead and improve system efficiency, as detailed in the following section.
-1:  Input: Thresholds $T_{h}$ $T_{l}$; Candidate token $t_{c}$
-2:  Input: High & low precision $P_{h}$ $P_{l}$
+If $t_{v}$’s significance exceeds $\frac{\alpha_{h}}{N}$, it remains in $KV_{h}$; if it falls within $[\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, $t_{v}$ is re-quantized to low precision and moved to $KV_{l}$, the low precision KV cache (line 9); otherwise, $t_{v}$ is pruned.
+Similarly, if $t_{c}$’s significance lies within $[\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, it is quantized to low precision and added to $KV_{l}$ (line 14).
+The least important token in $KV_{l}$ is designated as the victim $t_{v}$ (line 15), and is further pruned if its significance falls below $\frac{\alpha_{l}}{N}$ (line 17).
+Discussion.
+The proposed KV compression policy is highly extensible, allowing for the potential use of more than three precision levels.
+We adopt three precision levels primarily to minimize metadata overhead and improve system efficiency, as will be detailed in the following section.
+Additionally, we use a shared set of thresholds for all attention heads, as our empirical studies demonstrate that this approach is sufficient to capture the varying sparsity patterns across different heads.
+Notably, LeanKV can efficiently support more flexible compression policies, such as tuning thresholds for each head individually, which could potentially further optimize the balance between model accuracy and memory efficiency.
+We leave the exploration of such flexible policies for future work.
+1:  Input: Parameters $\alpha_{h}$, $\alpha_{l}$; High & low precision $P_{h}$ $P_{l}$
+2:  Input: Candidate token $t_{c}$; Sequence length $N$
 3:  Input: High & low precision KV cache $KV_{h}$ $KV_{l}$
 4:  Function: Significance $Score$; Quantization $Quant$;
-5:  if Score($t_{c}$) $\geq$ $T_{h}$ then
+5:  if Score($t_{c}$) $\geq$ $\frac{\alpha_{h}}{N}$ then
 6:     $KV_{h}$.add(Quant($t_{c}$, $P_{h}$))
@@
 9:        $KV_{h}$.remove($t_{v}$),  $KV_{l}$.add(Quant($t_{v}$, $P_{l}$))
-10:     else if Score($t_{v}$) ¡ $T_{l}$ then
+10:     else if Score($t_{v}$) < $T_{l}$ then
 11:        $KV_{h}$.remove($t_{v}$)
 12:     end if
-13:  else if Score($t_{c}$) $\geq$ $T_{l}$ then
+13:  else if Score($t_{c}$) $\geq$ $\frac{\alpha_{l}}{N}$ then
 14:     $KV_{l}$.add(Quant($t_{c}$, $P_{l}$))
 15:     $t_{v}$ = $argmin_{t\in KV_{l}}$(Score($t$))
-16:     if Score ($t_{v}$) ¡ $T_{l}$ then
+16:     if Score ($t_{v}$) < $T_{l}$ then
 17:        $KV_{l}$.remove($t_{v}$)
@@
 19:  end if
-Algorithm 1 Unified compression (generation phase)
-5. System Design
-In this section, we present the design of the LeanKV system that leverages the proposed unified KV cache compression algorithm to boost LLM inference efficiency.
-We first discuss the challenges that the compression algorithm poses to system performance, and then present novel techniques to tackle these challenges.
-5.1. Challenges
+Algorithm 1 KV compression policy (generation)
+5 Memory Management
+In this section, we describe LeanKV’s memory management mechanism which efficiently supports the three levels of differentiation in the KV cache.
+We first highlight the challenges posed by these differentiations in memory management, and then present parallel KV compaction, a novel memory management technique that effectively tackles the challenges.
+5.1 Challenges
 The Need for Flexible Paging.
-In PagedAttention (kwon2023efficient, ), all tokens are stored at the same precision, making a fixed page format sufficient.
-However, Hetero-KV and mixed-precision quantization introduce varying precision levels across different tokens, rendering a fixed page format inadequate.
+In PagedAttention [32], all tokens are stored at the same precision, allowing for a fixed page format.
+However, the differentiation of key and value vectors, along with varying token importance, introduces multiple precision levels both within individual tokens and across different tokens, making a fixed page format insufficient.
 Specifically, a fixed page format requires conservatively allocating high-precision slots for all tokens regardless of their actual precisions, leading to considerable memory wastage.
 Suppose memory is allocated to accommodate the highest precision K8V4, a token with K4V2 precision would waste 50% of the memory.
-Additionally, such a fixed page format results in misaligned memory accesses, which hinder memory bandwidth utilization and reduce overall computational efficiency.
-To fully realize the benefits of Hetero-KV and mixed-precision quantization, we need a more flexible paging design that can dynamically adapt to varying precision levels, while ensuring compact memory usage and efficient memory accesses.
-Increased Memory Management Overhead.
-The per-head KV cache compression algorithm in LeanKV introduces both significant latency and metadata overhead in memory management.
-In PagedAttention, the memory management complexity is $O(\#requests)$ regardless of the number of heads, as it partitions memory uniformally across heads and assigns identical page IDs to each head.
-In contrast, LeanKV must handle varying numbers of high-precision and low-precision tokens across different heads, which increases the complexity to $O(\#requests\times\#heads)$.
-If not carefully optimized, this overhead could overshadow the performance gains from cache compression.
+Worse, such a fixed page format results in misaligned memory accesses, which hinder memory bandwidth utilization and reduce overall computational efficiency.
+Substantial KV Compaction Overhead.
+In PagedAttention, memory management complexity is $O(\#requests)$, as memory is partitioned uniformly across attention heads, with identical page IDs assigned to each head.
+In contrast, LeanKV must handle varying numbers of high- and low-precision tokens across different heads, mapping these irregular memory requirements to physical memory.
+We term this process KV compaction, whose complexity is $O(\#requests\times\#heads)$, significantly higher than that of PagedAttention.
+If KV compaction were performed sequentially on CPUs, its overhead could overshadow the performance gains from KV cache compression.
 Moreover, to track both high-precision and low-precision tokens, a straightforward approach would involve using separate data structures for each precision, resulting in increased metadata overhead.
-5.2. On-GPU Memory Management
-A detailed analysis of the memory management process reveals opportunities for parallelization and acceleration.
-Specifically, memory management for the unified compression algorithm consists of two phases: planning and coordination.
+5.2 Parallel KV Compaction
+A detailed analysis of the KV compaction process reveals opportunities for parallelization and performance optimization.
+KV compaction can be divided into two phases: planning and coordination.
 In the planning phase, each head independently determines its memory allocation requirements.
 The subsequent coordination phase synchronizes these per-head requirements and maps them to the GPU’s physical memory.
-For instance, assume two attention heads require 16 and 32 bytes, respectively, as determined in the planning phase, with all GPU physical memory available.
+For instance, consider two attention heads require 16 and 32 bytes, respectively, as determined in the planning phase, with all GPU physical memory available.
 Then in the coordination phase, physical memory bytes 0 to 15 are allocated to the first head, and bytes 16 to 48 to the second.
@@
 Parallelizing the coordination phase is more challenging, as it requires synchronization across heads.
-However, we observe that this phase can be parallelized effectively via parallel prefix sum (nakano2012prefix-sum, ).
-Thus, we design an on-GPU memory management system to efficiently handle per-head dynamic memory allocation and recycling.
-This system introduces three GPU-resident data structures that collectively tackle the aforementioned challenges: the $(1)$ unified paging to enable flexible paging, $(2)$ the circular free page list to efficiently execute the coordination phase in memory management, and $(3)$ the bidirectional page table to minimize metadata overhead for tracking high-precision and low-precision tokens.
-Unified Paging.
-We partition the available GPU memory into evenly sized pages, each configured upon allocation to store tokens at a given precision.
-Each unified page consists of six segments: quantized keys, quantization metadata for keys, quantized values, quantization metadata for values, token scores, and positions.
+However, we observe that this phase can be parallelized effectively via parallel prefix sum [44], provided that the free memory region is contiguous.
+Thus, we propose parallel KV compaction, a novel management technique that efficiently performs per-head dynamic memory allocation and recycling in parallel directly on the GPU.
+Parallel KV compaction is enabled by three GPU-resident data structures collectively: the $(1)$ unified pages to enable flexible paging, $(2)$ the circular free page list to efficiently parallelize memory management, and $(3)$ the bidirectional page table to minimize metadata overhead for tracking tokens of differentiated precisions.
+Unified Pages.
+Unified pages abstract away the complexity of differentiated precisions both within individual tokens and across different tokens, thereby simplifying the implementation of parallel KV compaction.
+Specifically, we partition the available GPU memory into evenly sized pages, each configured upon allocation to store tokens at a given precision.
+Each unified page is organized into six segments: quantized keys, quantization metadata for keys, quantized values, quantization metadata for values, token scores, and positions.
 The quantization metadata includes scales and zero points for the key and value vectors.
 The number of tokens stored per page is adjusted according to the quantization configuration, ensuring compact memory usage.
-Unified paging allows flexible handling of tokens with varying precisions while maintaining a consistent memory layout, effectively supporting our unified compression algorithm.
-Furthermore, by consolidating keys, values, and their metadata into a single structure, unified paging enhances data locality and eliminates the need for scattered lookups, improving memory access efficiency during attention computation.
+Furthermore, by consolidating keys, values, and their metadata into a single structure, unified pages enhance data locality and eliminate the need for scattered lookups, improving memory access efficiency during attention computation.
 Circular Free Page List.
-To facilitate efficient page allocation and recycling, we manage all available page IDs in a centralized, GPU-resident data structure called circular free page list.
-The list tracks free pages using a pair of start and end pointers, which wrap to the beginning of the list upon reaching the end, maintaining a circular structure.
-When a page is allocated, the start pointer increments to the next available page ID; when a page is freed, the end pointer increments to add the freed page ID back to the list.
-Notably, both the available and unavailable regions in the list remain contiguous, enabling the coordination phase in memory management to be parallelized via parallel prefix sum.
-Specifically, after each head determines the number of pages to be allocated or freed in the planning phase, a parallel prefix sum operation computes a unique offset for each head relative to the start or end pointer, ensuring non-conflicting operations within the list.
+The circular free page list serves as the cornerstone of parallel KV compaction, facilitating the parallelization of memory allocation and recycling by maintaining both free and used pages in contiguous regions.
+This centralized, GPU-resident data structure contains all Page IDs and tracks free pages through a pair of pointers: a start pointer for allocations and an end pointer for recycling.
+These pointers wrap around to the beginning of the list upon reaching the end, forming a circular structure.
+When a page is allocated, the start pointer advances to the next available Page ID; when a page is freed, the end pointer advances to add the released Page ID back into the list.
+Both the available and unavailable regions in the list remain contiguous, enabling the coordination phase in memory management to be parallelized via parallel prefix sum.
+In parallel KV compaction, after each head determines the number of pages to be allocated or freed, a parallel prefix sum operation computes a unique offset for each head relative to the start or end pointer, ensuring non-conflicting operations within the list.
 For memory allocation, each head concurrently retrieves its new page IDs from its designated region in the list, with the start pointer incremented by the cumulative number of pages required across all heads.
-Similarly, for memory recycling, each head concurrently writes freed page IDs to it designated region, with the end pointer incremented by the total number of pages released.
+Similarly, for memory recycling, each head concurrently writes freed page IDs to its designated region, with the end pointer incremented by the total number of pages released.
 Bidirectional Page Table.
-Similar to PagedAttention, LeanKV maintains a page table for each attention head, mapping requests to physical memory.
-Each entry in the page table is a list of page IDs allocated to the head for a specific request.
+Similar to PagedAttention, LeanKV maintains a page table for each attention head, mapping each request to its corresponding list of page IDs.
 To avoid the doubled metadata overhead associated with maintaining separate page tables for high- and low-precision pages, LeanKV introduces a unified, GPU-resident data structure called the bidirectional page table.
-This structure efficiently supports the use of three precision levels for KV cache compression, as proposed in Section 4, minimizing the metadata overhead.
+This structure efficiently supports the use of three precision levels for KV cache, as proposed in Section 4, minimizing the metadata overhead.
 In each entry of the bidirectional page table, high-precision page IDs grow from the left side of the list, while low-precision page IDs grow from the right, dynamically adapting to the precision requirements of the workload.
@@
 By contrast, the KV cache for a single request occupies 2 GB.
-5.3. Memory Management Workflow
-In this section, we delve into LeanKV’s memory management workflow, demonstrating how the proposed data structures interact to efficiently support unified KV cache compression.
-Notably, this memory management workflow is executed once per inference step for all requests in the batch and across all layers in the LLM.
-This design choice ensures sufficient parallelism for efficient GPU execution and amortizes the associated GPU kernel launch overheads.
-Figure 5. Memory management flow in the prompt phase.
+5.3 KV Compaction Workflow
+We delve into the KV compaction workflow during the prompt and generation phases, illustrating how the three data structures interact to enable differentiated KV cache compression.
+Notably, KV compaction is executed once per inference step for all requests in the batch and all layers in the LLM.
+This design ensures sufficient parallelism for efficient GPU execution of parallel KV compaction, while amortizing the associated GPU kernel launch overheads.
+Figure 5: Memory management flow in the prompt phase.
 Prompt Phase.
-Figure 5 illustrates the memory management workflow during the prompt phase using an example request with eight tokens.
+Figure 5 illustrates the KV compaction workflow during the prompt phase using an example request with eight tokens.
 In this setting, one high-precision page can store two tokens, and one low-precision page can store four.
@@
 Once a request is finished, all pages allocated for that request are recycled, freeing up memory for incoming requests.
-6. Implementation
+6 Implementation
 We implement LeanKV on top of vLLM, comprising 4.5K lines of CUDA/C++ code and 9K lines of Python.
-We first outline the architecture of LeanKV and then detail our custom GPU attention kernel, designed to efficiently support the unified cache compression algorithm.
-6.1. LeanKV Architecture
+We first outline the architecture of LeanKV and then detail our custom GPU attention kernel, designed to efficiently support the differentiated KV cache compression.
+6.1 LeanKV Architecture
 The architecture of LeanKV is illustrated in Figure 6.
@@
 Each GPU hosts one worker, responsible for executing a segment of the model.
-LeanKV utilizes tensor parallelism (shoeybi2019megatron, ), where each worker handles a partition of each matrix multiplication and a subset of the attention heads.
+LeanKV utilizes tensor parallelism [54], where each worker handles a partition of each matrix multiplication and a subset of the attention heads.
 Each worker includes a dedicated memory manager to oversee the KV cache for its assigned attention heads, using the method described in Section 5.
@@
 Additionally, each worker incorporates an execution engine for model computation.
-To support unified KV cache compression, the execution engine integrates a KV compressor and a custom GPU attention kernel.
-After computing key and value vectors, the KV compressor is invoked to compress them using the algorithm in Section 4, storing the results in the KV cache.
+To support differentiated KV cache compression, the execution engine integrates a KV compressor and a custom GPU attention kernel.
+After computing key and value vectors, the KV compressor is invoked to compress them following the policy described in Section 4, storing the results in the KV cache.
 The custom GPU attention kernel then efficiently computes the attention output using the compressed KV cache.
-Figure 6. LeanKV architecture.
-6.2. Efficient Attention Kernel
-We develop a custom GPU attention kernel that efficiently supports per-head dynamic sparsity and Hetero-KV quantization, fully harnessing the reduction in memory access volume and translating it into accelerated performance.
+Figure 6: LeanKV architecture.
+6.2 Efficient Attention Kernel
+We develop a custom GPU attention kernel that efficiently supports differentiated KV cache compression, fully harnessing the reduction in memory access volume and translating it into accelerated performance.
 Overall, we assign each CUDA thread block to process a single attention head per sequence.
 However, mixed-precision quantization introduces potential load imbalance, as high-precision tokens demand more bandwidth, which is the bottleneck of attention, than low-precision ones.
-Thanks to the unified paging mechanism that stores tokens of the same precision together in a page, we mitigate this issue by having thread warps iterate over high-precision pages first, followed by low-precision pages, ensuring that threads within a warp remain load-balanced.
+Thanks to unified pages that store tokens of the same precision together, we mitigate this issue by having thread warps iterate over high-precision pages first, followed by low-precision pages, ensuring that threads within a warp remain load-balanced.
 Each page is processed in two phases: dot product between the query and keys to derive attention scores, and weighted sum of values.
 These two phases follow different computation patterns: reduction over the feature dimension and token dimension, respectively.
-To ensure coalesced and vectorized memory accesses for both phases, we design tailored data layouts and parallelization strategies, elaborated in the following paragraphs.
+To ensure coalesced and vectorized memory accesses for both phases, we design tailored data layouts and parallelization strategies, elaborated as follows.
 For key processing, each warp handles one page at a time, with threads in the warp divided into groups responsible for distinct keys.
@@
 We organize the layout of values as [ $\frac{F}{V_{\text{group}}}$, $\frac{N_{\text{tokens}}}{V_{\text{vec}}}$, $V_{\text{group}}$, $V_{\text{vec}}$] accordingly.
-7. Evaluation
-We first evaluate the effectiveness of LeanKV’s unified KV cache compression algorithm and analyze the sensitivity of different models to KV cache compression.
-Then, We demonstrate the improvements in throughput achieved by LeanKV.
-7.1. Experiment Setup
-We evaluate LeanKV on six models across three families: Llama2-7B and 70B (touvron2023llama, ), Llama3-8B and 70B (dubey2024llama, ), Mistral-7B (jiang2023mistral, ), and Mixtral-8x7B (jiang2024mixtral, ).
-Mixtral-8x7B employs a Mixture-of-Experts (MOE) architecture, while the remaining models are dense.
-Performance of LeanKV is evaluated on five benchmarks, grouped into two categories.
-The first includes tasks requiring advanced reasoning and logic: GSM8K (cobbe2021training, ) (mathematics), HumanEval (chen2021evaluating, ) (coding), and MMLU (hendrycks2020measuring, ) (academic level question answering).
-The second category includes traditional text processing tasks: SQuAD (rajpurkar2018know, ) (reading comprehension) and CNN-DailyMail (nallapati2016abstractive, ) (summarization).
-Throughput and latency is measured on Nvidia L40 GPUs, each with 48GB of memory (L40, ).
-7.2. Evaluating Unified Cache Compression
-We first demonstrate the effectiveness of Hetero-KV quantization and dynamic sparsity individually, and then evaluate the end-to-end benefits of the proposed unified KV cache compression.
-To mitigate the noise and randomness in floating-point computations, each experiment is repeated five times with the dataset randomly shuffled, and the reported results are the average of these runs.
-Evaluating Hetero-KV.
-Figure 7 shows the performance of Hetero-KV quantization, namely K8V4 and K4V2, compared to FP16 baselines, on GSM8K and SQuAD, representing benchmarks from the two categories.
-To validate our intuition that key vectors have a more significant impact than value vectors (Section 3.1), we additionally evaluate the mirror configurations of Hetero-KV, namely K4V8 and K2V4, where values are stored with higher precision than keys.
-K8V4 maintains the accuracy of the FP16 baseline across both benchmarks for all models, whereas its mirror configuration, K4V8, results in noticeable degradation, particularly on the reasoning-intensive GSM8K benchmark.
-The performance gap between K4V2 and K2V4 is even more pronounced.
-On GSM8K, K4V2 retains at least 77.8% of FP16 performance, while K2V4 reduces accuracy to nearly zero for all models.
-On SQuAD, K4V2 achieves 94.7% to 99.4% of FP16 performance, whereas K2V4 achieves only 28.2% to 60.0%.
-These results confirm that key vectors play a more critical role than value vectors, and demonstrate that Hetero-KV is effective in maintaining the accuracy of the FP16 baseline.
-Figure 7. Performance of Hetero-KV.
+7 Evaluation
+We first evaluate the effectiveness of LeanKV’s differentiated KV cache compression policy and analyze the sensitivity of different models to KV cache compression.
+Next, we evaluate the efficiency of the memory manager as well as the end-to-end throughput improvements achieved by LeanKV.
+7.1 Experiment Setup
+We evaluate LeanKV on five models spanning three major families: Llama3-8B and 70B [60], Qwen2.5-7B and 32B [28], and the recently released thinking model QwQ-32B [58], which generates extended chains of thought and exhibits strong performance on complex reasoning tasks.
+Our evaluation covers a diverse set of mainstream benchmarks across multiple task domains, including general knowledge (MMLU [26] and MMLU-Pro [55]), mathematics (GSM8K [13] and MATH [34]), code generation (HumanEval+ [11, 39] and MBPP+ [7, 39]), and long-context understanding (LongBench [8]).
+In addition, we assess LeanKV’s performance on QwQ-32B using two particularly challenging benchmarks: AIME24 [3], a high-school level mathematics competition dataset, and GPQA [48], which evaluates graduate-level science reasoning.
+Throughput and latency are measured on NVIDIA L40 GPUs, each equipped with 48 GB of memory [45].
+7.2 Differentiated KV Compression Policy
+We first assess the effectiveness of differentiated KV quantization and dynamic sparsity individually, and then evaluate the combined benefits of the proposed differentiated KV cache compression.
+To ensure the reliability of our results and mitigate the influence of noise and randomness inherent in floating-point computations, each experiment is repeated five times with the dataset randomly shuffled, and the reported results represent the average across these runs.
+Evaluating Differentiated KV Quantization.
+Figure 7 shows the performance of differentiated KV quantization on GSM8K and HumanEval+, namely K8V4 and K4V2, compared to FP16 baselines, across Llama-3 8B & 70B and Qwen-2.5 7B.
+To validate our intuition that keys have a more significant impact than values (Section 3.1), we additionally evaluate the mirror configurations, namely K4V8 and K2V4, where values are stored with higher precision than keys.
+Results show that K8V4 matches the accuracy of the FP16 baseline across all models and benchmarks.
+In contrast, its mirror configuration, K4V8, exhibits noticeable performance degradation, particularly for Qwen2.5-7B, where accuracy drops to nearly zero on both tasks.
+This pronounced sensitivity to 4-bit key quantization in Qwen2.5-7B is likely due to its GQA architecture, which applies aggressive KV compression with a queries-per-KV ratio of 7, substantially higher than the ratio of 4 used in Llama-3 8B.
+Similarly, K4V2 retains over 65% of FP16 accuracy across both benchmarks on Llama-3 8B and 70B, while the mirror configuration K2V4 results in near-zero accuracy.
+These results confirm that keys play a more critical role than values, and demonstrate the effectiveness of differentiated KV quantization in preserving accuracy while enabling aggressive compression.
+Figure 7: Performance of differentiated KV quantization.
+Figure 8: Performance of dynamic vs. static sparsity.
 Evaluating Dynamic Sparsity.
-We evaluate the effectiveness of per-head dynamic sparsity (Section 3.2) in identifying critical tokens, comparing it to the static sparsity method used in H2O (zhang2024h2o, ), which allocates an equal memory budget to all attention heads.
-The results are shown in Figure 8, where we use the Mistral-7B model as an example and all tokens are stored in FP16.
-For reasoning-intensive workloads such as GSM8K and HumanEval, dynamic sparsity significantly outperforms static sparsity, with the performance gap widening as the fraction of pruned tokens increases.
-For traditional text processing tasks, however, the benefits of dynamic sparsity are less pronounced: it slightly outperforms static sparsity on SQuAD and performs similarly on CNN-DailyMail.
-The difference between SQuAD and CNN-DailyMail can be attributed to their differing sensitivities to token pruning.
-SQuAD (reading comprehension) requires precise retrieval of specific words from the prompt, making it more sensitive to pruning critical tokens.
-In contrast, CNN-DailyMail (summarization) inherently involves discarding less relevant information to generate concise summaries, resulting in lower sensitivity to token pruning.
-To summarize, dynamic sparsity is superior to static sparsity for reasoning-intensive tasks and text processing that involves precise information retrieval.
-Figure 8. Performance of dynamic vs. static sparsity.
-Evaluating Unified KV Cache Compression.
-Finally, we evaluate the end-to-end performance of the unified KV cache compression described in Section 4.
-For each benchmark, we perform grid search over the high-precision threshold $T_{h}$ (with discrete values from {$0.5\%$, $1\%$, $5\%$, $10\%$, $20\%$}) and the low-precision threshold $T_{l}$ (with discrete values from $0\%$, $0.1\%$, $0.5\%$, $1\%$}), selecting the configurations that minimize the KV cache memory budget under different performance constraints.
-To account for the inherent randomness in floating-point computations, we consider a configuration to be performance-neutral if its average performance falls within three standard deviations of the FP16 baseline.
-Table 1 reports the KV cache memory usage of LeanKV, normalized to vLLM (kwon2023efficient, ), across all evaluated models and benchmarks.
-The results are presented for two configurations: performance-neutral, which ensures no degradation in output quality, and less-than-$5\%$ performance degradation, a common target in prior work (lee2024infinigen, ; fu2024moa, ; zhao2024atom, ; lin2024qserve, ; lin2024awq, ; frantar2023sparsegpt, ) that enables more aggressive memory savings while preserving acceptable quality.
-Across models and workloads, LeanKV achieves performance neutral with $18.9\%$ to $28.9\%$ of the KV cache memory budget, and less-than-$5\%$ performance degradation with $7.6\%$ to $20.6\%$ of the memory budget.
-Additionally, Figure 9 shows the Pareto performance-memory curves of LeanKV on the reasoning-intensive benchmark GSM8K and text processing benchmark SQuAD.
-The results are compared against the FP16 baseline, conventional 8-bit quantization (K8V8) and 4-bit quantization (K4V4), and the pruning-only technique H2O (zhang2024h2o, ).
-For GSM8K, K8V8 preserves the FP16 baseline performance, while K4V4 leads to a noticeable performance degradation of $2.8\%$ to $7.3\%$.
-In comparison, LeanKV achieves FP16-equivalent performance with only $19.8\%$ to $23.1\%$ of the memory budget, even less than K4V4, which requires $25\%$ memory.
-Moreover, LeanKV offers a more flexible memory-performance tradeoff, with performance degradation limited to less than $5\%$ while using just $9.4\%$ to $16.3\%$ of the memory.
-In contrast, under similar memory constraints, H2O results in near-zero accuracy.
-Similar trends are observed on the SQuAD benchmark: LeanKV matches the performance of FP16 and K8V8 with less memory than K4V4, which causes noticeable performance degradation, and outperforms H2O under comparable memory budgets.
-Notably, across all evaluated models and benchmarks, LeanKV achieves performance-neutral with the K8V4-K4V2 configuration, where critical tokens are stored with K8V4 precision, non-critical tokens with K4V2, and no pruning is applied.
-For the less-than-$5\%$ degradation setting, the optimal configuration varies across models.
-For large 70B models, less-than-$5\%$ degradation is achieved with the more aggressive K4V2-K4V1 quantization without pruning, as pruning significantly degrades performance even on top of K8V4-K4V2.
-In contrast, for smaller models, less-than-$5\%$ degradation is achieved by further pruning the least important tokens on top of K8V4-K4V2.
-These varying configurations in low-memory regimes are also visualized in Figure 9, where each data point for LeanKV is annotated with its corresponding configuration.
-This disparity underscores the varying sensitivity of models to KV cache compression techniques, emphasizing the need for a unified framework capable of adapting to diverse scenarios to maximize efficiency.
-In summary, LeanKV provides a superior performance-memory tradeoff compared to conventional methods.
-It maintains the performance of FP16 and 8-bit quantization while using even less memory than 4-bit quantization.
-Additionally, LeanKV offers a smooth tradeoff in the low-memory regime ($10\%$ to $20\%$ of memory), where pruning-only methods experience significant performance degradation.
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-Figure 9. KV cache memory normalized to vLLM vs. benchmark performance tradeoff of LeanKV
-Table 1. Memory usage of LeanKV normalized to vLLM across models and benchmarks.
-Models Benchmarks GSM8K MMLU HumanEval SQUAD CNN-DailyMail Llama2-7B perf-neutral 18.9% 22.8% 23.1% 21.2% 21.6% 5% perf-loss 11.3% 11.5% 18.9% 14.6% 7.6% Llama2-70B perf-neutral 21.3% 21.4% 21.2% 21.1% 20.1% 5% perf-loss 19.5% 15.9% 18.2% 16.0% 15.7% Llama3-8B perf-neutral 20.7% 26.7% 26.6% 24.1% 20.0% 5% perf-loss 9.4% 20.6% 22.1% 12.8% 8.8% Llama3-70B perf-neutral 23.1% 21.5% 22.7% 22.7% 19.5% 5% perf-loss 16.3% 16.1% 21.2% 16.0% 15.7% Mistral-7B perf-neutral 26.8% 28.9% 23.0% 26.2% 19.4% 5% perf-loss 17.7% 13.6% 15.1% 15.9% 9.7% Mixtral-8x7B perf-neutral 19.8% 20.1% 21.4% 21.1% 20.0% 5% perf-loss 13.8% 8.9% 10.7% 8.9% 10.0%
-7.3. Sensitivity to KV Cache Compression
-We investigate the varying sensitivity of models to KV cache compression, using the Llama model family as an example, due to their similar training data and methodologies.
-Figure 10 illustrates the normalized performance degradation relative to the FP16 baseline for Llama2-7B, Llama3-8B, and Llama3-70B as the fraction of pruned tokens increases.
-Among the models, Llama2-7B exhibits the least performance degradation, even when compared to Llama3-8B, which has a comparable parameter size.
-This resilience can be attributed to Llama2-7B’s MHA architecture, which introduces greater redundancy in the KV cache compared to the more memory efficient GQA architecture utilized by Llama3-8B and Llama3-70B.
-Llama3-70B experiences the most significant performance degradation, suggesting that larger models distribute attention more evenly across tokens, making them more sensitive to pruning.
-Figure 11 further illustrates the normalized performance degradation of the K4V2-K4V1 configuration as the fraction of tokens stored in K4V1 increases.
-The Llama3-70B model exhibits the greatest resilience to low-bit quantization, maintaining near-baseline performance even when 60% of tokens are stored in K4V1.
-In comparison, smaller models show noticeable performance degradation even when all tokens are stored in K4V2, highlighting their greater sensitivity to aggressive quantization.
-These observations justify the two key trends outlined earlier: larger models and newer models employing more memory-efficient network architectures are more sensitive to pruning, while smaller models are more adversely impacted by low-bit quantization.
-Figure 10. Sensitivity to KV cache pruning.
-Figure 11. Sensitivity to low-bit KV cache quantization.
-7.4. Evaluating Latency and Throughput
-In this section, we provide a comprehensive evaluation of LeanKV’s performance, comparing it against state-of-the-art systems including vLLM and H2O across various models and configurations.
-All experiments were conducted on a machine equipped with eight NVIDIA L40 GPUs.
-For models that fit into a single GPU’s memory, such as Llama2-7B and Llama3-7B, we use one GPU.
-For larger models like Llama2-70B and Llama3-70B, we parallelize the computation across four GPUs.
+We evaluate the effectiveness of per-head dynamic sparsity (Section 3.3) in identifying critical tokens, comparing it to the static sparsity method used in SnapKV [35] and H2O [69], which allocates an equal memory budget to all attention heads.
+Figure 8 reports results on Qwen2.5-7B and Llama3-8B across GSM8K and HumanEval+.
+The x-axis indicates the percentage of tokens pruned, while the y-axis shows the resulting task accuracy.
+Dynamic sparsity significantly outperforms static sparsity in Llama3-8B, maintaining full accuracy with 50% tokens pruned on GSM8k, and 10% tokens pruned for the more sensitive HumanEval+.
+For Qwen2.5-7B, dynamic sparsity also consistently surpasses static sparsity, achieving accuracy higher than 80% on GSM8k with 50% tokens pruned.
+While the improvement is less pronounced on HumanEval+, dynamic sparsity still yields higher accuracy at equivalent pruning ratios.
+In summary, per-head dynamic sparsity is superior to static sparsity by leveraging head-specific significance.
+Figure 9: Calibrating the high- and low-precision threshold parameters $\alpha_{h}$ and $\alpha_{l}$ on the training split of the MATH dataset.
+Parameter Calibration.
+As described in Section 4, LeanKV compares the significance score of a token with its theoretical average, $\frac{1}{N}$, where $N$ denotes the sequence length.
+A token is quantized to high precision if its score exceeds $\frac{\alpha_{h}}{N}$, to low precision if its score falls within the interval $[\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, and is pruned otherwise.
+To calibrate the appropriate parameters $\alpha_{h}$ and $\alpha_{l}$, we utilize the training split of MATH [34].
+We select MATH as the calibration dataset as it is comprised of complex mathematical reasoning tasks with high text information density, where even the removal of a single symbol can invalidate an argument.
+This makes MATH a highly suitable choice for parameter fitting, as the resulting compression parameters are expected to generalize well across other tasks.
+Additionally, MATH provides a dedicated training split, ensuring that parameter fitting does not involve testing data and mitigating potential overfitting.
+We profile a consistent range of parameters across all evaluated models, as depicted in Figure 9.
+The x-axis represents the profiled parameter values, while the y-axis shows the resulting accuracy on the calibration dataset.
+Specifically, we profile $\alpha_{h}$ within the integer range $[1,5]$, progressively relaxing the high-precision threshold beyond the theoretical average.
+This adjustment accounts for aggregating scores from multiple attention heads by maximum for GQA, as mentioned in Section 4.
+The profiled values of $\alpha_{h}$ achieve FP16-equivalent accuracy across all models, except for Qwen2.5-7B, which exhibits heightened sensitivity to 4-bit key quantization.
+Based on our profiling, we select the values of $\alpha_{h}$ that yield the highest accuracy for each model: $\alpha_{h}=1$ for Llama-3 8B and 70B, $\alpha_{h}=3$ for Qwen2.5-32B and QwQ-32B, and $\alpha_{h}=0$ for Qwen2.5-7B.
+Subsequently, we profile the low-precision threshold $\alpha_{l}$, while keeping $\alpha_{h}$ fixed.
+We observe that setting $\alpha_{l}=0.1$ induces at least 5% accuracy degradation on the larger models.
+As a result, we profile five equally spaced values within the range $[0,0.1]$ and select the one yielding the highest accuracy for each model.
+The final selected values of $\alpha_{l}$ are $0.02$ for Llama3-8B, $0.04$ for Qwen2.5-7B, and 00 for the remaining models.
+While the profiled values cover a limited portion of the entire parameter space, as will be demonstrated in the following sections, LeanKV significantly outperforms existing baselines with these fitted parameters, highlighting its potential effectiveness.
+Figure 10: KV cache memory normalized to vLLM vs. benchmark performance tradeoff of LeanKV.
+Evaluating Differentiated Compression Policy.
+Finally, we evaluate the combined benefits of LeanKV’s KV cache compression policy.
+We compare LeanKV against several state-of-the-art baselines, including pruning-based methods such as H2O [69], SnapKV [35], and DuoAttention [63], quantization-based approaches such as 4-bit KV and KIVI [41], as well as KV partial loading techniques like Quest [56].
+Table 1 summarizes the accuracy and KV cache memory usage of LeanKV normalized to vLLM [32] baseline, with the parameters tuned as described, for the evaluated non-thinking models.
+For the pruning-based baselines and Quest, we tune their memory usage to 50% of the baseline.
+LeanKV achieves near-lossless accuracy relative to the FP16 baseline, with an average degradation of only $0.3\%$, while using $19.3\%$ to $36.7\%$ memory.
+In contrast, all other baselines incur more significant accuracy degradation at comparable or higher memory usage.
+Furthermore, LeanKV dynamically adapts its memory usage based on the task’s information density.
+For example, LeanKV allocates less memory for the 5-shot MMLU benchmark compared to the 0-shot HumanEval+, where the prompt consists solely of the function name and comments.
+LeanKV also outperforms baseline methods in long-context scenarios, as shown in Table 2, where we present results from one benchmark in each category of LongBench.
+For Llama3-8B and Qwen2.5-7B, LeanKV achieves memory usage of $14.9\%$ and $27.0\%$, respectively, while baseline methods are tuned to $25\%$ memory usage.
+Notably, LeanKV consistently achieves superior accuracy across all categories, with an average degradation of only $0.4\%$, while utilizing less or comparable memory compared to the baseline methods.
+Furthermore, the performance gap between LeanKV and baseline methods becomes even more pronounced for the thinking model QwQ-32B on more challenging benchmarks, as demonstrated in Table 3.
+LeanKV matches the FP16 baseline performance using only $27.4\%$ of the memory on average, while all other methods experience significant performance degradation.
+On the math competition benchmark AIME24, the most accurate baseline, 4-bit KV quantization incurs a $16.0\%$ accuracy degradation.
+On the graduate-level science benchmark GPQA, the accuracy of pruning-based and KV partial loading methods drops to almost zero.
+Lastly, Figure 10 depicts the performance-memory trade-offs of LeanKV across representative benchmarks under the profiled parameter values, namely $\alpha_{h}\in[0,5]$ (step size of 1) for all models and $\alpha_{l}\in[0.02,0.1]$ (step size of 0.02) for Llama3-8B and Qwen2.5-7B.
+We also include the accuracy of baseline methods within similar memory usage ranges for comparison.
+LeanKV consistently matches the FP16 baseline accuracy across different models and benchmarks within the profiled range of memory usage, demonstrating its robustness.
+In contrast, baseline methods experience significant accuracy degradation, particularly pruning-based and KV partial loading methods, which suffer rapid accuracy declines as memory usage decreases.
+In summary, LeanKV provides a superior performance-memory tradeoff compared to state-of-the-art methods.
+Table 1: Accuracy and memory usage of LeanKV and baseline methods across models and benchmarks.
+Numbers in parentheses represent memory usage normalized to FP16.
+The memory usage is 50% for SnapKV, Quest, H2O and DuoAttn, 25% for INT4, and 12.5% for KIVI due to 2-bit quantization.
+[…]
+Table 2: Evaluating LeanKV on LongBench. LeanKV uses 14.9% memory for Llama3.1-8B and 27.0% for Qwen2.5-7B. For Quest and SnapKV, the memory usage is 25%.
+[…]
+Table 3: Evaluating LeanKV on QwQ-32B thinking model.
+[…]
+7.3 System Performance
+In this section, we provide a comprehensive evaluation of LeanKV’s performance, comparing it against state-of-the-art systems including vLLM, Quest, and SnapKV across various models.
+All experiments are conducted on a machine equipped with eight NVIDIA L40 GPUs.
+For models that fit within the memory of a single GPU, such as Llama3-8B and Qwen2.5-7B, we use one GPU.
+For larger models, we parallelize the computation across multiple GPUs: four GPUs for Llama3-70B, and two GPUs for Qwen2.5-32B and QwQ-32B.
 Our evaluation focuses on latency breakdown, attention kernel speedup, and end-to-end throughput to demonstrate the efficiency and scalability of LeanKV.
 Memory Management Overhead.
-Figure 12 shows the latency breakdown of LeanKV during one inference step with a sequence length of 1024 tokens, evaluated across different batch sizes.
-The memory management overhead is remarkably low, contributing less than 0.2% of the total latency in the prompt phase and under 0.9% in the generation phase.
+Figure 11 compares LeanKV’s on-GPU parallel KV compaction with an alternative implementation that performs multi-threaded memory management on the CPU.
+Evaluated with a sequence length of 1024 tokens across different batch sizes, LeanKV reduces memory management latency by up to three orders of magnitude relative to the on-CPU approach.
+For one entire inference step, LeanKV significantly outperforms the on-CPU approach, particularly in the generation phase, where the memory management overhead on the CPU far exceeds the model execution time.
+Therefore, LeanKV’s on-GPU parallel memory management is essential for ensuring that the performance benefits of KV cache compression are fully realized without being overshadowed by memory management overhead.
+Figure 12 further illustrates the latency breakdown of LeanKV during one inference step.
+The memory management overhead is remarkably low thanks to LeanKV’s on-GPU parallel KV compaction, contributing less than 0.2% of the total latency in the prompt phase and under 0.9% in the generation phase.
 Model execution dominates the latency in both phases, accounting for 96–97% of the latency in the prompt phase and 92–93% in the generation phase.
 Notably, as the batch size increases, the percentage of time spent on model execution rises slightly, reflecting the scalability of LeanKV.
-For comparison, in vLLM, model execution contributes 97% of the latency in the prompt phase and 95% in the generation phase.
-These results demonstrate that LeanKV effectively minimizes latency contributions from memory management and KV compression, ensuring that these auxiliary processes do not become performance bottlenecks.
-Figure 12. Latency breakdown of LeanKV.
+(a) Memory management
+(b) One entire inference step
+Figure 11: Latency comparison between parallel KV compaction and on-CPU multi-threaded memory management.
+Figure 12: Latency breakdown of LeanKV.
 Attention Kernel Speedup.
@@
 Additionally, LeanKV achieves greater speedups on longer sequences, indicating that its bandwidth optimization techniques become increasingly effective as the sequence length grows.
-Figure 13. Attention kernel speedup.
+Figure 13: Attention kernel speedup.
 End-to-End Throughput.
-Figure 14 presents the end-to-end throughput and achieved batch sizes of LeanKV compared to vLLM and H2O.
-We conduct experiments on 1000 sequences sampled from the ShareGPT (ShareGPT, ) dataset, setting the generation length to 4096 tokens.
-The compression thresholds for LeanKV are selected based on the most challenging dataset, MMLU, making them the most conservative choices.
-In configurations with no accuracy loss, LeanKV achieves throughput speedups ranging from 2.0×2.0$\times$2.0 × to 2.5×2.5$\times$2.5 × compared to vLLM across various models.
-When allowing a 5% accuracy loss, LeanKV achieves even higher speedups, ranging from 3.0×3.0$\times$3.0 × to 6.9×6.9$\times$6.9 ×.
-These improvements are directly correlated with the larger batch sizes supported by LeanKV, enabled by its KV cache compression techniques.
-For instance, with the Llama2-7B model, LeanKV sustains a batch size of 62 without accuracy loss and 177 with 5% accuracy loss, far exceeding vLLM’s batch size of 22.
+Figure 14 presents the end-to-end throughput and achieved batch sizes of LeanKV compared to vLLM, Quest, and SnapKV.
+The maximum generation length is set to 16K tokens for QwQ-32B, 8K for Qwen2.5-32B, and 4K for the other models.
+We evaluate on 1000 sequences sampled from the MATH dataset [34], which naturally elicits chain-of-thought reasoning and typically leads to long generations reaching the specified limit.
+Compression thresholds for LeanKV are adopted from Figure 9.
+Across all models, LeanKV consistently achieves higher throughput than prior systems.
+Notably, LeanKV achieves a remarkable 5.4×5.4$\times$5.4 × higher throughput over vLLM on the QwQ-32B thinking model, while Quest and SnapKV achieve only 1.6×1.6$\times$1.6 × and 1.8×1.8$\times$1.8 × speedups, respectively.
+This result highlights the superior efficiency of LeanKV in processing long sequences, making it particularly well-suited for tasks that require extended reasoning.
+The throughput improvements are directly correlated with the larger batch sizes supported by LeanKV, enabled by its KV cache compression techniques.
+For instance, with the QwQ-32B model, LeanKV sustains a batch size of 15.9, far exceeding vLLM’s batch size of 2.7.
 The increased batch size enables LeanKV to maximize the utilization of GPU compute capacity, leading to higher throughput.
-Llama3-7B exhibits a less significant speedup compared to Llama2-7B because it utilizes GQA, which reduces its KV cache size by 4$\times$.
-As a result, vLLM already supports a relatively large batch size of 51.5 for Llama3-7B, narrowing the batch size advantage that LeanKV can leverage.
-Nevertheless, LeanKV still achieves up to 3.4$\times$ higher throughput over vLLM on Llama3-7B.
-Notably, H2O supports even smaller batch sizes than vLLM and consequently has lower throughput.
-This is because H2O consumes additional memory during the prompt phase for intermediate results required to calculate compression metrics, limiting the batch size that can be supported.
-In contrast, we design a FlashAttention-compatible kernel using Triton, which avoids the additional memory overhead caused by intermediate results and achieves performance on par with vLLM.
-Figure 14. Throughput and achieved batch size.
-8. Conclusion
-We propose LeanKV, a unified KV cache compression framework designed to enhance LLM serving efficiency through three key techniques: Hetero-KV, per-head dynamic sparsity, and unified KV compression.
-To support these innovations, LeanKV incorporates system optimizations including unified paging and on-GPU parallel memory management.
-Our experiments show that LeanKV achieves 1.9×1.9$\times$1.9 × to 2.5×2.5$\times$2.5 × higher throughput over state-of-the-art systems without compromising model accuracy, and up to 6.9×6.9$\times$6.9 × higher throughput with minimal accuracy degradation.
+In contrast, Quest supports the same batch size as vLLM because it retains the entire KV cache without reducing memory usage.
+Its speedup comes primarily from faster attention computation, as it processes only a subset of tokens deemed important.
+However, the process of estimating token importance incurs additional overhead, limiting the overall efficiency of Quest.
+Figure 14: Throughput and achieved batch size.
+8 Conclusion
+We propose LeanKV, a framework to enhance LLM serving efficiency by exploiting three levels of differentiation in KV cache: differentiated precision for keys and values, hierarchical compression based on varying token importance, and per-head dynamic sparsity.
+At the core of LeanKV is the parallel KV compaction technique that efficiently handles irregular memory requirements across requests and attention heads, effectively translating memory savings into performance gains.
+Our evaluation shows that LeanKV is able to compress the KV cache by 2.7×2.7$\times$2.7 × to 5.7×5.7$\times$5.7 × without compromising accuracy on complex workloads that require sophisticated reasoning and long-generation capabilities, and enhances throughput by 1.9×1.9$\times$1.9 × to 5.4×5.4$\times$5.4 ×.

### v3

@@
-Unifying KV Cache Compression for Large Language Models with LeanKV
+DiffKV: Differentiated Memory Management for Large Language Models with Parallel KV Compaction
 Abstract
-Large language models (LLMs) exhibit exceptional performance but incur significant serving costs due to their substantial memory requirements, with the key-value (KV) cache being a primary bottleneck.
-Existing KV cache compression techniques, such as quantization and pruning, apply uniform treatment to both keys and values, and discard unimportant tokens entirely, overlooking the fine-grained differences in significance of various components within the KV cache.
-To address these limitations, we introduce LeanKV, a framework that advances KV cache compression by exploiting three levels of differentiation in the KV cache: (1) the differing impact of keys and values on attention computation, (2) the varying importance of tokens, and (3) the diverse dynamic sparsity patterns across attention heads.
-At the core of LeanKV is an on-GPU memory manager that compacts fragmented free memory list into contiguous regions in parallel, effectively translating sparsity in the KV cache into performance gains.
-We evaluate LeanKV on several mainstream models, including the recent "thinking model".
-LeanKV is able to compress the KV cache by 2.7×2.7$\times$2.7 × to 5.7×5.7$\times$5.7 × with near-lossless accuracy on complex workloads requiring sophisticated reasoning and long-generation capabilities, and enhances throughput by 1.9×1.9$\times$1.9 × to 5.4×5.4$\times$5.4 ×.
+Large language models (LLMs) demonstrate remarkable capabilities but face substantial serving costs due to their high memory demands, with the key-value (KV) cache being a primary bottleneck.
+State-of-the-art KV cache compression techniques, such as quantization and pruning, apply uniform treatment to both keys and values, and discard unimportant tokens entirely, overlooking the fine-grained distinctions in the significance of individual KV cache components.
+To address such limitations, we introduce DiffKV, a novel framework for efficient KV cache compression that exploits three levels of differentiation in the KV cache: (1) the differing impact of keys and values on attention computation, (2) the varying importance of tokens, and (3) the diverse dynamic sparsity patterns across attention heads.
+These levels of differentiation introduce irregular memory usage patterns across different requests and attention heads, posing significant scalability challenges for memory management.
+To address these challenges, DiffKV proposes an on-GPU memory manager that compacts fragmented free memory list into contiguous regions in parallel, effectively translating sparsity in the KV cache into performance gains.
+We evaluate DiffKV on several mainstream LLMs, including the emerging thinking models that generate extended chains of thought.
+DiffKV is able to compress the KV cache by $2.7\times$ to $5.7\times$ with near-lossless accuracy on complex workloads requiring sophisticated reasoning and long-generation capabilities, and enhances throughput by $1.9\times$ to $5.4\times$.
+Source codes of DiffKV are available at https://github.com/zyqCSL/DiffKV.
 1 Introduction
@@
 Given the pervasive adoption of LLMs, enhancing serving efficiency has become critically important [7, 21, 38, 62, 75, 28, 56, 43].
-LLMs are fundamentally based on the Transformer architecture [61, 62], which employs the attention mechanism to generate tokens sequentially, based on the input prompt and previously generated tokens.
-Within the attention mechanism, each token is transformed into query, key and value feature vectors.
-The attention scores, derived from the dot product of the query and key vectors, enable the model to weigh the significance of different tokens.
-Subsequently, the output is computed as the weighted sum of the value vectors, based on the attention scores.
-To capture diverse relationships between tokens, the attention mechanism is typically divided into multiple heads [61, 51, 5].
-Each head independently computes attention, and their outputs are concatenated and transformed to create a richer representation of the input data.
 To avoid redundant computation, LLM inference frameworks typically cache intermediate key and value tensors in memory, commonly referred to as the KV cache [39].
 The size of the KV cache scales linearly with both sequence length and the number of concurrent requests, often comprising over 90% of total memory consumption [39, 74, 22, 73].
-As state-of-the-art models continue to support longer sequences [17, 30, 38, 4, 57, 60], and with the rise of recent "thinking models" [23, 29, 58] that generate extended reasoning processes, the KV cache has emerged as a critical bottleneck for LLM serving efficiency.
-It limits the number of concurrent requests that can fit in memory and increases attention computation latency due to its memory bandwidth-bound nature [15, 14].
-Recognizing the critical role of the KV cache in limiting serving efficiency, researchers have investigated various compression techniques, primarily focused on quantization [37, 2, 27] and pruning [69, 64, 10, 20, 35].
-Quantization reduces the size of the KV cache by converting floating-point feature values into integers, representing them based on pre-defined intervals.
-Pruning, on the other hand, reduces the token sequence length by eliminating unimportant tokens based on attention scores.
-While promising, these techniques overlook the fine-grained differentiation in the significance of various KV cache components, limiting their overall effectiveness.
-First and foremost, existing quantization methods save key and value vectors of all tokens at a uniform precision, disregarding the distinct roles of keys and values and the varying importance of tokens.
-Second, current pruning methods employ static memory allocation across attention heads, neglecting the dynamic and heterogeneous attention patterns observed across heads and requests.
-Moreover, these pruning methods have not been comprehensively evaluated on workloads requiring advanced reasoning, such as programming and mathematics.
-In this paper, we present LeanKV, a framework that advances KV cache compression by leveraging three levels of differentiation within the KV cache:
+As state-of-the-art models continue to support longer sequences [20, 36, 45, 4, 66, 68], and with the rise of recent thinking models [26, 35, 67] that generate extended reasoning processes, the KV cache has emerged as a critical bottleneck for LLM serving efficiency.
+It limits the number of concurrent requests and increases attention computation latency due to its memory bandwidth bound nature [17, 16].
+Researchers have investigated various compression techniques, primarily focused on pruning [76, 71, 11, 23, 42] and quantization [44, 2, 32].
+Pruning reduces the token sequence length by eliminating unimportant tokens based on attention scores.
+Quantization, on the other hand, reduces the size of the KV cache by converting floating-point feature values into lower-precision representations.
+While effective, these methods overlook the nuanced variations in importance across different components of the KV cache.
+First, existing pruning approaches employ a one-size-fits-all strategy, statically allocating memory uniformly across attention heads despite the per-head dynamic attention sparsity patterns.
+Second, current quantization schemes apply uniform precision to all key and value vectors, ignoring both the distinct roles of keys versus values in the attention mechanism, and the varying significance of different tokens.
+Finally, both pruning and quantization treat all requests uniformly, overlooking variability in information density across requests.
+To address these limitations and advance KV cache compression, we propose DiffKV, a novel framework that exploits three levels of differentiation within the KV cache:
 1.
 Differentiated Keys and Values.
-LeanKV assigns higher precision to key vectors than to value vectors, based on the observation that key vectors play a more significant role in attention computations.
+DiffKV assigns higher precision to key vectors than to value vectors, based on the observation that keys (critical for attention score computation) play a more significant role than values (used primarily for weighted aggregation) in attention computations.
 2.
 Token Importance Differentiation.
-Guided by attention scores, LeanKV stores tokens at varying precision levels based on their significance.
+Guided by attention scores, DiffKV stores tokens at varying precision levels based on their significance.
 Most critical tokens are quantized at higher precision, less critical tokens are quantized at lower precision, and the least significant tokens are pruned.
-This hierarchical approach enables a flexible tradeoff between model accuracy and memory efficiency.
 3.
 Per-head Dynamic Sparsity.
-Attention score distributions exhibit dynamic sparsity across both heads and requests.
-Within a single request, the number of critical tokens varies across heads, and for the same head, the number of critical tokens also varies across different requests.
-To address this variability, LeanKV dynamically identifies critical tokens per-head, per-request, adapting memory allocation accordingly.
-These differentiations wthin the KV cache pose new challenges in memory management.
-In LeanKV, each request requires varying amounts of memory for each attention head to store tokens at different precision levels.
-Moreover, key and value vectors of the same token are compressed at different precisions.
-These complexities significantly increase memory management overhead, compared to traditional attention mechanisms, where memory allocation is uniform across all heads.
-To address these challenges, LeanKV employs an on-GPU memory manager that efficiently handles the irregular memory usage patterns across heads and requests via parallel KV compaction.
-Rather than performing memory management operations sequentially on CPU for each head and request, parallel KV compaction optimizes memory allocation and recycling by packing fragmented free memory lists into contiguous regions directly on the GPU in parallel.
-Parallel KV compaction is supported by three key on-GPU data structures:
+Attention score distributions exhibit dynamic sparsity across both heads and requests: the number of important tokens varies across different heads, and for the same head under different requests.
+To address this variability, DiffKV dynamically identifies critical tokens per-head, per-request, adapting memory allocation accordingly.
+Figure 1: KV cache memory allocation patterns for (a) pruning, (b) uniform quantization, and (c) DiffKV across two attention heads in two 5-token requests.
+Boxes represent retained tokens (annotated with token IDs), with box size proportional to memory usage.
+Percentages indicate total memory usage relative to an uncompressed KV cache.
+Figure 1 compares the KV cache usage of pruning, quantization, and DiffKV across two heads in two 5-token requests.
+In both requests, pruning uniformly eliminates one token per head (reducing cache usage to 80%), failing to adapt to heads with naturally sparser attention patterns.
+Quantization, meanwhile, applies INT8 precision uniformly to all FP16 values, halving memory usage but missing opportunities for finger-grained savings by treating keys and values differently based on their relative importance.
+Moreover, both approaches treat two requests identically, overlooking their distinct sparsity patterns.
+In contrast, DiffKV exhibits more adaptive memory usage.
+In Request 1, for Head A, DiffKV prunes one token and further distinguishes precision across the rest: Token 0, deemed highly important, is stored at high precision (key in INT8, value in INT4), while Tokens 1, 3, and 4 are stored at lower precision (keys in INT4, values in INT2).
+This reduces usage to 25% for keys and 12.5% for values.
+For Head B, DiffKV instead identifies two tokens as highly important, resulting in 30% and 15% KV cache usage for keys and values, respectively.
+Averaged across both heads, DiffKV uses 20.6% of memory for Request 1, outperforming pruning and quantization.
+Additionally, DiffKV also adapts flexibly across requests.
+In Request 2, Head A stores one token at high precision and the remaining four at lower precision without pruning, whereas Head B prunes two tokens and stores two others at high precision.
+These examples highlight DiffKV’s ability to leverage both per-head dynamic sparsity and request-level variation, enabling more aggressive and flexible compression than prior methods.
+Although DiffKV achieves a high KV cache compression ratio, it introduces significant memory usage irregularity.
+As depicted in Figure 1, pruning and quantization maintain uniform KV cache utilization across attention heads.
+In contrast, DiffKV exhibits substantial variability, not only across different heads but also between the key and value within individual heads.
+To mitigate fragmentation, the memory manager must precisely track and allocate memory per head and per request.
+However, this introduces a scalability challenge: even a modestly sized LLM like Llama3-8B [20] comprises hundreds of attention heads, and with hundreds of concurrent requests, the memory manager must handle tens of thousands of heterogeneous memory regions during each inference step.
+Considering that model execution time per step is on the order of tens of milliseconds, improper memory management could render the overhead of dynamic allocation prohibitive, negating the benefits of KV cache compression.
+To overcome the scalability challenge, DiffKV introduces an on-GPU memory manager that efficiently handles the irregular memory usage patterns via parallel KV compaction, which optimizes memory allocation and recycling by packing fragmented free memory lists into contiguous regions directly on the GPU in parallel.
+Parallel KV compaction is enabled by three essential on-GPU data structures:
 1.
 Unified Pages.
-GPU memory is divided into fixed-size pages containing vectors stored at various precisions.
-Each page is dynamically parsed according to the precision requirements of the head and request, while an attention GPU kernel is co-designed with the page layout to maximize memory bandwidth utilization.
+GPU memory is partitioned into fixed-size pages, each storing tokens at a specific precision.
+For example, a page may hold tokens with keys in INT8 and values in INT4.
+Each page is dynamically configured and parsed according to the precision requirements of the head and request.
 2.
@@
 All page IDs are managed in a centralized, GPU-resident circular list.
-Both free and used pages are kept in contiguous regions, enabling efficient page allocation and recycling via parallel prefix sum.
+Both free and used pages are kept in contiguous regions, enabling efficient page allocation and recycling via parallel prefix sum [29].
 3.
 Bidirectional Page Table.
-Instead of maintaining separate page tables for high- and low-precision pages, LeanKV consolidates both into a single bi-directional page table to minimize metadata overhead.
-High precision page IDs grow from the left side of the table, and low-precision page IDs grow from the right side.
-We implement LeanKV on vLLM [32], a state-of-the-art LLM inference system, and evaluate it across three prominent model families including the emerging thinking model.
-The evaluation is conducted on representative workloads that demand advanced reasoning capabilities, such as mathematics and coding.
-Our experimental results demonstrate that LeanKV offers a superior cost-accuracy tradeoff compared to existing KV cache management methods.
-Specifically, LeanKV compresses the KV cache by 2.7×2.7$\times$2.7 × to 5.7×5.7$\times$5.7 × with near-lossless accuracy, resulting in throughput improvement of 1.9×1.9$\times$1.9 × to 5.4×5.4$\times$5.4 ×.
-The throughput improvement is primarily enabled by the parallel KV compaction mechanism, which reduces management overhead by three orders of magnitude, decreasing the proportion of memory management time in total execution latency from up to $76\%$ to less than $1\%$ compared to a sequential CPU-based implementation.
+Instead of maintaining separate page tables for high- and low-precision pages, DiffKV consolidates both into a single bi-directional page table to minimize metadata overhead.
+High-precision page IDs grow from the left side of the table, and low-precision page IDs grow from the right side.
+Additional precision levels can also be efficiently accommodated by employing multiple bi-directional page tables.
+We have implemented DiffKV on vLLM [39] and evaluate it across multiple LLMs, including the emerging thinking models QwQ-32B [67], R1-Distill-Qwen-14B and R1-Distill-Llama-8B [26].
+Our experiment results demonstrate that Diff-
+KV offers a superior cost-accuracy tradeoff. Specifically, Diff-
+KV compresses the KV cache by $2.7\times$ to $5.7\times$ with near-lossless accuracy, resulting in throughput improvement of $1.9\times$ to $5.4\times$.
+Notably, DiffKV is, to the best of our knowledge, the first KV cache compression framework evaluated on thinking models and complex reasoning tasks that require advanced chain-of-thought (CoT) capabilities, while achieving FP16-comparable generation quality.
 2 Background and Related Work
-In this section, we describe Transformer-based large language models and review existing techniques for KV cache management aimed at improving LLM serving efficiency.
+In this section, we describe Transformer-based large language models and review existing techniques for KV cache management.
 2.1 Large Language Models
-Transformer Architecture.
-Large language models (LLMs) are predominantly built upon the Transformer architecture [61], which is used to capture complex long-range dependencies in sequential data.
-At the core of Transformer architecture is the attention mechanism, which allows each token in a sequence to weigh the significance of other tokens when constructing its contextualized representation.
-Mathematically, the attention computation is defined as:
+Large language models (LLMs) are predominantly built upon the Transformer architecture [69].
+At the core of Transformer is the attention mechanism, which allows each token in a sequence to weigh the significance of other tokens when constructing its contextualized representation.
+During autoregressive inference, the attention mechanism operates in a causal manner, attending only to preceding tokens.
+Mathematically, the standard attention computation is defined as:
 $\operatorname{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})_{i}$ $=\sum_{j=1}^{i}\operatorname{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)_{ij}\mathbf{v}_{j}$ (1)
 $=\sum_{j=1}^{i}\frac{\exp\left(\frac{\mathbf{q}_{i}\cdot\mathbf{k}_{j}}{\sqrt{d}}\right)}{\sum_{n=1}^{i}\exp\left(\frac{\mathbf{q}_{i}\cdot\mathbf{k}_{n}}{\sqrt{d}}\right)}\mathbf{v}_{j}$
-$\mathbf{Q}$ and $\mathbf{K}$ are matrices of size $l\times d$, representing the queries and keys, respectively, where $l$ denotes the number of tokens and $d$ represents the feature dimensionality.
+$\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ are matrices of size $l\times d$, representing the queries, keys and values respectively, where $l$ denotes the number of tokens processed so far in the sequence and $d$ represents the feature dimensionality.
 Vectors $\mathbf{q}_{i}$, $\mathbf{k}_{i}$, and $\mathbf{v}_{i}$ correspond to the query, key, and value of the $i^{\text{th}}$ token.
-The dot product between $\mathbf{Q}$ and $\mathbf{K}$ measures the relevance between each pair of query and key.
-The softmax function normalizes these dot products to form a probability distribution, referred to as attention scores, which are used to compute a weighted sum of the value vectors and produce the output.
 To capture a broader range of interactions between tokens, Transformer models employ multi-head attention (MHA), where each head independently computes attention using distinct projections of queries, keys, and values.
-The outputs from different heads are concatenated and linearly transformed to create a more expressive representation.
-Grouped-query attention (GQA) [5] improves the efficiency of MHA by allowing a group of query heads to share the same projection of keys and values, referred to as a KV head.
-Autoregressive Generation and KV Cache.
-The autoregressive generation process of LLMs consists of two phases: the prompt phase, during which the model computes latent representations for all tokens in the prompt, the user-provided context, and generates the first new token, and the generation phase, where the model iteratively generates subsequent tokens by attending to both the prompt and previously generated tokens.
+Grouped-query attention (GQA) [6, 60] improves the efficiency of MHA by allowing a group of query heads to share the same projection of keys and values, referred to as a KV head.
+LLM execution involves two phases: the prompt phase, where the model computes latent representations for all tokens in the prompt and generates the first new token, and the generation phase, where subsequent tokens are generated iteratively.
 To avoid redundant computations across generation steps, KV cache is introduced to store the keys and values of all previous tokens.
@@
 KV Cache Quantization.
-Quantization [43, 16, 22, 27] reduce the KV cache size by approximating high-precision floating points with discrete low-bit values.
-To quantize a given tensor $\mathbf{X}$, we first calculate the scale $s$ and zero point $z$ as shown in Eq.
-(2), where $\mathbf{X}_{\text{max}}$ and $\mathbf{X}_{\text{min}}$ represent the maximum and minimum values in $\mathbf{X}$, and $b$ is the bit-width.
-Then, quantization is applied element-wise as: $\mathbf{Q}=\operatorname{round}\left(\frac{\mathbf{X}+z}{s}\right)$.
-During inference, the original values are approximately reconstructed via dequantization: $\hat{\mathbf{X}}=s\cdot\mathbf{Q}-z$.
+Quantization [50, 18, 25, 32] reduces the KV cache size by approximating high-precision floating points with discrete low-bit values.
+For a tensor $\mathbf{X}$, we first compute the scale $s$ and zero point $z$ based on $\mathbf{X}_{min}$ and $\mathbf{X}_{max}$, then apply quantization element-wise as: $\mathbf{Q}=\operatorname{round}\left(\frac{\mathbf{X}-z}{s}\right)$.
+During inference, the original tensor is approximately reconstructed via dequantization: $\hat{\mathbf{X}}=s\cdot\mathbf{Q}+z$.
 Unlike the quantized tensor $\mathbf{Q}$, the metadata $s$ and $z$ are kept in higher precision (e.g., FP16) to ensure more accurate dequantization.
-$s=\frac{\mathbf{X}_{\text{max}}-\mathbf{X}_{\text{min}}}{2^{b}-1},\quad z=-\mathbf{X}_{\text{min}}$ (2)
 State-of-the-art KV cache quantization methods such as Atom [77] and Qserve [44] apply this process to each key and value vector independently.
@@
 Complementary to KV cache pruning, another line of work [61, 40] addresses GPU memory constraints by offloading the entire KV cache to CPU memory and selectively loading important tokens to the GPU for attention computations.
-However, these methods do not inherently reduce the KV cache size and incur significant latency due to data transfers between the CPU and GPU.
+However, these methods do not inherently reduce the KV cache size and incur additional latency due to data transfers between the CPU and GPU.
 3 The Case for Differentiated KV
-We present the case for three levels of differentiation within the KV cache, motivating the design of LeanKV.
+We motivate the design of DiffKV by presenting the case for three levels of differentiation within the KV cache: (i) the differentiated impacts of keys and values (Section 3.1), (ii) differentiated token importance (Section 3.2), and (iii) per-head dynamic sparsity (Section 3.3).
+Building on these insights, Section 3.4 describes how DiffKV combines them together and highlights its contributions over prior systems.
 3.1 Differentiated Impacts of Keys and Values
@@
 To this end, we reformulate the attention mechanism as a weighted sum of unit vectors, as presented in Equation 2.
-This formulation decomposes each input token’s contribution into two components: the unit vector $\frac{v_{j}}{|v_{j}|}$, representing only the direction of the token in feature space, and a coefficient, defined as the product of the attention score and the norm of the value vector, which determines the relative significance of the token’s direction.
+This formulation decomposes each input token’s contribution into two components: the unit vector $\frac{v_{j}}{|v_{j}|}$, obtained by dividing the token’s value vector by its L2 norm to capture only its direction in feature space, and a coefficient, defined as the product of the attention score and the norm of the value vector, which determines the relative significance of the token’s direction.
 This allows us to assess the relative importance of attention scores and value vectors by examining their contributions to these coefficients.
@@
 We therefore conclude that key vectors exert a broader and more impactful influence than value vectors, motivating further exploration of processing key and value vectors at distinct precision levels.
+We validate this intuition in Section 7.2, showing that K8V4 and K4V2, which quantize keys to 8/4 bits and values to 4/2 bits respectively, outperform their mirror configurations K4V8 and K2V4.
 Figure 2: Distribution of attention score and value vector norm in Llama3-8B.
@@
 By exploiting these differences, we can apply finer-grained compression strategies that go beyond uniformly quantizing all tokens [77, 44] or exclusively pruning the least important ones [76, 11].
-Specifically, we store progressively less important tokens at decreasing precision levels, ultimately leading to the pruning of the least important tokens.
-To investigate the feasibility of this hierarchical approach, we experiment with a simple design: quantizing both the key and value vectors of significant tokens to 8 bits while pruning insignificant ones.
-Figure 2 compares the performance of the FP16 baseline, 8-bit quantization, pruning, and the combination of 8-bit quantization with pruning on the general knowledge question answering benchmark MMLU [26] using the Llama3-8B model.
-For both pruning and the combined approach, approximately 50% of insignificant tokens are pruned.
-The resulting accuracy is nearly indistinguishable among the four approaches.
-However, the combination of 8-bit quantization and pruning achieves the lowest memory usage.
-The improved efficiency of this combined approach motivates us to further explore the spectrum of KV cache compression by introducing varying levels of precision to accommodate tokens of differing significance.
-Figure 2: Accuracy and normalized KV cache memory usage of the Llama3-8B model on MMLU
-3.3 Dynamic Sparsity Patterns
-Existing approaches to KV cache pruning primarily rely on static memory allocation for each transformer layer in LLMs, either distributing memory uniformly [64, 69, 20, 35] or allocating it based on offline profiling [10].
-However, LLMs exhibit dynamic sparsity patterns that vary across both attention heads and requests, which static memory allocation cannot adequately handle.
-To evaluate the sparsity of attention computation, we analyze the minimum number of critical tokens required to retain the majority of information, specifically by preserving a target percentage (e.g., 95%) of the total attention score.
-We begin by investigating dynamic sparsity patterns across layers.
-Using the first thousand articles from the Wikitext dataset, we evaluate the Llama3-8B model.
+Figure 3 shows the per-token attention scores in the $8^{th}$ layer, selected as a representative layer of Llama3-8B [20], on a sequence randomly sampled from the Wikitext dataset [49].
+Similar sparsity patterns are observed across the other layers as well.
+This non-uniform distribution of token importance motivates a hierarchical compression strategy that allocates memory based on token significance: high-precision storage (e.g., K8V4) for the most important tokens, lower precision (e.g., K4V2) for moderately important tokens, and pruning for the least important.
+This hierarchical approach enables a smoother trade-off between memory savings and generation quality than uniform quantization methods (e.g., Atom [77], QServe [44]) and pruning-only approaches (e.g., H2O [76], SnapKV [42]).
+Figure 3: Per-token attention scores in the $8^{th}$ layer of Llama3-8B on a sequence randomly sampled from Wikitext.
+3.3 Per-head Dynamic Sparsity Patterns
+LLMs exhibit dynamic sparsity patterns that vary across both attention heads and requests: the number of critical tokens can differ not only between heads but also for the same head under different requests.
+To characterize the sparsity of attention, we analyze the minimum number of critical tokens required to retain the majority of information, specifically by preserving a target percentage (e.g., 95%) of the total attention score.
+We begin by investigating dynamic sparsity patterns across layers, and evaluate the Llama3-8B model with the first thousand articles from the Wikitext dataset.
 Figure 4 illustrates the average number of critical tokens per layer, aggregated across all KV heads, required to retain 95% of the total attention score.
@@
 Furthermore, within individual KV heads, the number of critical tokens can vary substantially across requests, as shown by the high standard deviation in some KV heads.
-These highly dynamic sparsity patterns, across layers, attention heads, and individual requests, underscore the necessity of adaptive memory management on a per-head and per-request basis.
 Figure 4: Number of critical tokens per layer in Llama3-8B to preserve 95% of the total attention score.
 Figure 5: Number of critical tokens per KV head in Llama3-8B to preserve 95% of the total attention score.
-3.4 Key Findings and Implications
+3.4 Main Insights and Implications
 Our findings reveal three critical levels of differentiation within the KV cache:
@@
   • Attention sparsity patterns vary across requests and attention heads, requiring dynamically managing memory resources on a per-request and per-head basis.
-Collectively, these insights guide the design of LeanKV, which integrates differentiated precision assignments for keys and values, hierarchical token importance handling, and per-request and per-head memory management.
+Collectively, these highly dynamic sparsity patterns, across layers, attention heads, and individual requests, underscore the necessity of adaptive memory management on a per-head and per-request basis.
+They motivate the design of DiffKV, which integrates differentiated precisions for keys and values, hierarchical compression based on importance, and per-request and per-head memory management.
+Contributions of DiffKV.
+DiffKV expands the design space for KV cache compression by jointly exploiting the three levels of differentiation in KV cache, enabling higher compression ratios with minimal quality degradation.
+Compared to prior work that explores importance-based KV cache pruning [76, 42, 5, 65] and mixed-precision quantization [77, 19, 30], DiffKV introduces the following key innovations:
+  • Request-aware differentiated KV quantization.
+DiffKV adaptively adjusts the mix of high- and low-precision tokens (e.g., K8V4, K4V2) on a per-request basis (Section 4), whereas prior work [77, 19, 5] typically applies a single static configuration across all requests.
+  • Sequence length-adaptive importance estimation.
+Diff-KV modulates the proportion of important tokens based on sequence length, retaining more tokens in shorter sequences to preserve quality, while enabling more aggressive compression for longer ones (Section 4).
+Prior approaches [76, 42, 5] typically use a fixed token retention ratio, regardless of sequence length.
+  • Per-head dynamic sparsity exploitation.
+DiffKV dynamically adjusts memory allocation per head, based on observed sparsity patterns (Section 3.3 & 4), whereas existing methods allocate memory uniformly across attention heads, missing optimization opportunities.
+  • Scalable on-GPU memory management.
+DiffKV introduces an efficient on-GPU memory manager that handles irregular per-head memory allocation patterns, effectively translating memory savings into performance gains (Section 5).
 4 KV Compression Policy
-We present the KV compression policy of LeanKV that addresses the three levels of differentiation within the KV cache.
-First, to reflect the greater impact of keys on attention computation than values, we propose quantizing keys at a higher precision than values.
+We present the KV compression policy of DiffKV, designed to address the three levels of differentiation within the KV cache.
+First, to reflect the greater impact of keys on attention computation compared to values, we propose quantizing keys at a higher precision than values.
 For example, keys and values can be quantized to 8 and 4 bits (K8V4), or 4 and 2 bits (K4V2).
-Second, to account for the varying importance of tokens, we introduce a hierarchical compression strategy that classifies tokens into three significance levels: the most important tokens are quantized at high precision (e.g., K8V4), moderately important tokens at lower precision (e.g., K4V2), and the least important tokens are pruned.
-Finally, to address the dynamic sparsity patterns across requests and attention heads, we propose an adaptive memory management approach.
+Second, to account for the varying importance of tokens, we introduce a hierarchical compression strategy that classifies tokens into three significance levels: the most important tokens are quantized at high precision (e.g., K8V4), moderately important tokens at lower precision (e.g., K4V2), and the least important tokens are pruned entirely.
+This strategy is both request-aware and sequence-length adaptive.
+It dynamically adjusts the mix of high- and low-precision tokens per request to reflect their varying sparsity patterns.
+Moreover, it retains a larger fraction of tokens at high precision for short sequences to preserve quality, while applying low precision quantization and pruning more aggressively for long sequences to maximize memory savings.
+Finally, to address the dynamic attention sparsity patterns across requests and attention heads, we propose an adaptive memory management approach.
 Rather than imposing a fixed memory budget, our approach allows each attention head to determine its memory requirements dynamically based on its specific sparsity pattern.
@@
 The compression policy then determines the appropriate precision for storing each token based on its significance.
-The significance of the $i^{th}$ token is calculated by averaging the $N-i$ attention scores it receives from subsequent tokens, where $N$ denotes the sequence length.
+The significance of the $i^{th}$ token is calculated by averaging the NiN-i attention scores it receives from subsequent tokens, where $N$ denotes the prompt sequence length.
 In the case of GQA and MHA, scores from all attention heads associated with the KV head are aggregated using the maximum operation.
 To mitigate premature compression, the most recent $W$ tokens are always quantized at high precision, where $W$ is typically set to 64.
-For the remaining tokens, the policy determines the precision level of the $i^{th}$ token by comparing its significance score to the theoretical average $\frac{1}{i}$, based on the intuition that tokens with scores below average are less critical.
-Specifically, the $i^{th}$ token is quantized at high precision if its significance score exceeds $\frac{\alpha_{h}}{i}$, at low precision if its score lies within the interval $[\frac{\alpha_{l}}{i},\frac{\alpha_{h}}{i}]$, and is pruned otherwise.
+For the remaining tokens, the policy determines the precision level of the $i^{th}$ token in a sequence-length adaptive manner, by comparing its significance score to the theoretical average $\frac{1}{i}$, based on the intuition that tokens with scores below average are less critical.
+Specifically, the $i^{th}$ token is quantized at high precision if its significance score exceeds $\frac{\alpha_{h}}{i}$, at low precision if its score lies within the interval $\frac{\alpha_{l}}{i},\frac{\alpha_{h}}{i}]$, and is pruned otherwise.
 The parameters $\alpha_{l}$ and $\alpha_{h}$, which define the thresholds for high- and low-precision quantization, are determined offline through profiling on a calibration dataset.
@@
 First, token $t_{c}$ is quantized at either a high or low precision and added to the corresponding section of the KV cache, or it may be completely pruned.
-Next, if $t_{c}$ is quantized, the least significant token $t_{v}$ in the corresponding precision section of the KV cache is either re-quantized to lower precision or pruned, depending on its significance.
+Next, if $t_{c}$ is quantized, the least significant token $t_{v}$ in the corresponding precision section of the KV cache is considered for further downgrading: it may be re-quantized to lower precision or pruned, depending on its significance.
 Essentially, this policy establishes a smooth downgrading path for the less important token: rather than being pruned directly, it is first re-quantized to low precision, with pruning occurring only if it remains insignificant.
@@
 Subsequently, the least significant token in $KV_{h}$ is identified as the victim token $t_{v}$ for more aggressive compression (line 7).
-If $t_{v}$’s significance exceeds $\frac{\alpha_{h}}{N}$, it remains in $KV_{h}$; if it falls within $[\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, $t_{v}$ is re-quantized to low precision and moved to $KV_{l}$, the low precision KV cache (line 9); otherwise, $t_{v}$ is pruned.
-Similarly, if $t_{c}$’s significance lies within $[\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, it is quantized to low precision and added to $KV_{l}$ (line 14).
+If $t_{v}$’s significance exceeds $\frac{\alpha_{h}}{N}$, it remains in $KV_{h}$; if it falls within $\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, $t_{v}$ is re-quantized to low precision and moved to $KV_{l}$, the low precision KV cache (line 9); otherwise, $t_{v}$ is pruned.
+Similarly, if $t_{c}$’s significance lies within $\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, it is quantized to low precision and added to $KV_{l}$ (line 14).
 The least important token in $KV_{l}$ is designated as the victim $t_{v}$ (line 15), and is further pruned if its significance falls below $\frac{\alpha_{l}}{N}$ (line 17).
 Discussion.
-The proposed KV compression policy is highly extensible, allowing for the potential use of more than three precision levels.
-We adopt three precision levels primarily to minimize metadata overhead and improve system efficiency, as will be detailed in the following section.
-Additionally, we use a shared set of thresholds for all attention heads, as our empirical studies demonstrate that this approach is sufficient to capture the varying sparsity patterns across different heads.
-Notably, LeanKV can efficiently support more flexible compression policies, such as tuning thresholds for each head individually, which could potentially further optimize the balance between model accuracy and memory efficiency.
+The proposed KV compression policy is highly extensible, allowing for the potential use of more than two quantization precision levels.
+We adopt two quantization precision levels primarily to minimize metadata overhead and improve system efficiency, as will be detailed in the following section.
+In Section 7.2, we empirically show that the two-level scheme (K8V4-K4V2) achieves near-lossless generation quality across multiple models and benchmarks.
+We use a shared set of thresholds for all attention heads, as our empirical studies demonstrate that this approach is sufficient to capture the varying sparsity patterns across different heads.
+Notably, DiffKV can efficiently support more flexible compression policies, such as tuning thresholds for each head individually, which could potentially further optimize the balance between model accuracy and memory efficiency.
 We leave the exploration of such flexible policies for future work.
-1:  Input: Parameters $\alpha_{h}$, $\alpha_{l}$; High & low precision $P_{h}$ $P_{l}$
-2:  Input: Candidate token $t_{c}$; Sequence length $N$
-3:  Input: High & low precision KV cache $KV_{h}$ $KV_{l}$
-4:  Function: Significance $Score$; Quantization $Quant$;
+1:Input: Parameters $\alpha_{h}$, $\alpha_{l}$; High & low precision $P_{h}$ & $P_{l}$
+2:Input: Candidate token $t_{c}$; Sequence length NN
+3:Input: High & low precision KV cache $KV_{h}$ & $KV_{l}$
+4:Function: Significance ScoreScore; Quantization QuantQuant;
 5:if Score($t_{c}$) $\geq$ $\frac{\alpha_{h}}{N}$ then
@@
 7:  $t_{v}$ = $argmin_{t\in KV_{h}}$(Score($t$))
-8:     if $T_{l}\leq$ Score($t_{v}$) $<T_{h}$ then
+8:  if $\frac{\alpha_{l}}{N}\leq$ Score($t_{v}$) < $\frac{\alpha_{h}}{N}$ then
 9:   $KV_{h}$.remove($t_{v}$),  $KV_{l}$.add(Quant($t_{v}$, $P_{l}$))
-10:     else if Score($t_{v}$) < $T_{l}$ then
+10:  else if Score($t_{v}$) < $\frac{\alpha_{l}}{N}$ then
 11:   $KV_{h}$.remove($t_{v}$)
@@
 15:  $t_{v}$ = $argmin_{t\in KV_{l}}$(Score($t$))
-16:     if Score ($t_{v}$) < $T_{l}$ then
+16:  if Score ($t_{v}$) < $\frac{\alpha_{l}}{N}$ then
 17:   $KV_{l}$.remove($t_{v}$)
@@
 5 Memory Management
-In this section, we describe LeanKV’s memory management mechanism which efficiently supports the three levels of differentiation in the KV cache.
+In this section, we describe DiffKV’s memory management mechanism which efficiently supports the three levels of differentiation in the KV cache.
 We first highlight the challenges posed by these differentiations in memory management, and then present parallel KV compaction, a novel memory management technique that effectively tackles the challenges.
 5.1 Challenges
-The Need for Flexible Paging.
+Flexible Paging.
 In PagedAttention [39], all tokens are stored at the same precision, allowing for a fixed page format.
-However, the differentiation of key and value vectors, along with varying token importance, introduces multiple precision levels both within individual tokens and across different tokens, making a fixed page format insufficient.
+However, the differentiation of key and value vectors, along with varying token importance, introduces multiple precision levels both within and across tokens, making a fixed page format insufficient.
 Specifically, a fixed page format requires conservatively allocating high-precision slots for all tokens regardless of their actual precisions, leading to considerable memory wastage.
@@
 Worse, such a fixed page format results in misaligned memory accesses, which hinder memory bandwidth utilization and reduce overall computational efficiency.
-Substantial KV Compaction Overhead.
-In PagedAttention, memory management complexity is $O(\#requests)$, as memory is partitioned uniformly across attention heads, with identical page IDs assigned to each head.
-In contrast, LeanKV must handle varying numbers of high- and low-precision tokens across different heads, mapping these irregular memory requirements to physical memory.
-We term this process KV compaction, whose complexity is $O(\#requests\times\#heads)$, significantly higher than that of PagedAttention.
-If KV compaction were performed sequentially on CPUs, its overhead could overshadow the performance gains from KV cache compression.
-Moreover, to track both high-precision and low-precision tokens, a straightforward approach would involve using separate data structures for each precision, resulting in increased metadata overhead.
+KV Compaction Scalability.
+At each LLM inference step, the memory management complexity is O(#requests)O(\#requests) for PagedAttention, as memory is partitioned uniformly among attention heads.
+In contrast, DiffKV must handle varying numbers of high- and low-precision tokens across different heads, mapping these heterogeneous memory requirements to physical memory.
+We term this process KV compaction, whose complexity is $O(\#requests\times\#heads)$ per inference step.
+This poses significant scalability challenges for memory management.
+Given that model execution time per step is on the order of tens of milliseconds, improper memory management could render the overhead of dynamic allocation prohibitive, negating the benefits of KV cache compression.
+Moreover, managing both high-precision and low-precision tokens would require separate data structures for each precision, resulting in increased metadata overhead.
 5.2 Parallel KV Compaction
-A detailed analysis of the KV compaction process reveals opportunities for parallelization and performance optimization.
+A detailed analysis of the KV compaction process reveals opportunities for parallelization.
 KV compaction can be divided into two phases: planning and coordination.
-In the planning phase, each head independently determines its memory allocation requirements.
+In the planning phase, each attention head independently determines its memory allocation requirements.
 The subsequent coordination phase synchronizes these per-head requirements and maps them to the GPU’s physical memory.
-For instance, consider two attention heads require 16 and 32 bytes, respectively, as determined in the planning phase, with all GPU physical memory available.
-Then in the coordination phase, physical memory bytes 0 to 15 are allocated to the first head, and bytes 16 to 48 to the second.
-The primary source of increased memory management complexity lies in the planning phase, which is perfectly parallel and well-suited to the GPU’s parallel compute capabilities.
+The primary source of increased memory management complexity lies in the planning phase, which is perfectly parallelizable and well-suited to the GPU’s parallel compute capabilities.
 Parallelizing the coordination phase is more challenging, as it requires synchronization across heads.
@@
 Both the available and unavailable regions in the list remain contiguous, enabling the coordination phase in memory management to be parallelized via parallel prefix sum.
-In parallel KV compaction, after each head determines the number of pages to be allocated or freed, a parallel prefix sum operation computes a unique offset for each head relative to the start or end pointer, ensuring non-conflicting operations within the list.
+In parallel KV compaction, after each head determines the number of pages to be allocated or freed, a parallel prefix sum operation computes a unique offset for each head relative to the start or end pointer.
+This step converts per-head memory demands into a globally consistent schedule, where each head is assigned a disjoint region of the free page list to read from (for allocation) or write to (for recycling), ensuring non-conflicting operations within the list.
 For memory allocation, each head concurrently retrieves its new page IDs from its designated region in the list, with the start pointer incremented by the cumulative number of pages required across all heads.
-Similarly, for memory recycling, each head concurrently writes freed page IDs to its designated region, with the end pointer incremented by the total number of pages released.
+Similarly, for memory recycling, each head writes freed page IDs to its designated region, with the end pointer incremented by the total number of pages released.
 Bidirectional Page Table.
-Similar to PagedAttention, LeanKV maintains a page table for each attention head, mapping each request to its corresponding list of page IDs.
-To avoid the doubled metadata overhead associated with maintaining separate page tables for high- and low-precision pages, LeanKV introduces a unified, GPU-resident data structure called the bidirectional page table.
-This structure efficiently supports the use of three precision levels for KV cache, as proposed in Section 4, minimizing the metadata overhead.
+Similar to PagedAttention, Diff-KV maintains a page table for each attention head, mapping each request to its corresponding list of page IDs.
+To avoid the doubled metadata overhead associated with maintaining separate page tables for high- and low-precision pages, DiffKV introduces a unified, GPU-resident data structure called the bidirectional page table.
+This structure efficiently supports the use of two quantization precision levels for KV cache, as proposed in Section 4, minimizing the metadata overhead.
 In each entry of the bidirectional page table, high-precision page IDs grow from the left side of the list, while low-precision page IDs grow from the right, dynamically adapting to the precision requirements of the workload.
@@
 This unified approach not only minimizes metadata overhead but also eliminates the need for separate lookups based on precision levels, thereby enhancing memory access efficiency during attention computations.
-The memory overhead of the bidirectional page table is minimal: for example, with a batch size of 128 on Llama-2-7B, which has 32 layers and 32 attention heads per layer, the total size of all bidirectional page tables is only 64 MB.
-By contrast, the KV cache for a single request occupies 2 GB.
+The memory overhead of the bidirectional page table is minimal: for example, with a batch size of 128 on Llama3-8B, which has 32 layers and 8 KV heads per layer, the total size of all bidirectional page tables is only 32 MB.
+By contrast, the KV cache for a single request occupies 1 GB.
 5.3 KV Compaction Workflow
@@
 Notably, KV compaction is executed once per inference step for all requests in the batch and all layers in the LLM.
-This design ensures sufficient parallelism for efficient GPU execution of parallel KV compaction, while amortizing the associated GPU kernel launch overheads.
+This design ensures sufficient parallelism for efficient GPU execution, while amortizing the associated GPU kernel launch overheads.
 Figure 6: Memory management flow in the prompt phase.
 Prompt Phase.
-Figure 5 illustrates the KV compaction workflow during the prompt phase using an example request with eight tokens.
-In this setting, one high-precision page can store two tokens, and one low-precision page can store four.
-Since the exact number of high- and low-precision pages required by each head is unknown prior to running the compression algorithm, we conservatively allocate four unified pages per head, assuming that all tokens will be stored at high precision.
-Each head then independently applies the KV cache compression algorithm to determine which tokens to store at high-precision, low-precision, or to prune.
-This step constitutes the planning phase, during which each head calculates its specific memory needs based on its unique sparsity pattern.
-In this example, the first head uses one high-precision page, while the second head uses two; both heads also require one low-precision page.
-High-precision pages are allocated from the left side of the bidirectional page table, while low-precision pages are allocated from the right side.
-Following the planning phase, any unused pages are marked for recycling, triggering the coordination phase.
-Here, we employ a parallel prefix-sum-based approach, as described in the previous section, to facilitate efficient and conflict-free page recycling across heads.
+Figure 6 illustrates the KV compaction workflow during the prompt phase for an example request with eight tokens.
+In this example, one high-precision page stores two tokens, and one low-precision page stores four.
+Since the exact numbers of high- and low-precision pages required by each head are unknown a priori, we conservatively allocate four unified pages per head, assuming all tokens are stored at high precision.
+Pages 5-8 and 9-12 are thus assigned to Head A and B, respectively, as shown in their page tables.
+The end pointer of circular free page list advances to Page 13, marking Pages 5-12 as allocated.
+Next comes the planning phase, during which each head independently applies the KV cache compression algorithm and calculates its specific memory requirements.
+In this example, Head A uses one high-precision page, while Head B uses two; both heads also require one low-precision page.
+We allocate high-precision pages left-to-right in the bidirectional page table and low-precision pages right-to-left: Head A uses Page 5 (high-precision) and Page 8 (low-precision), while Head B uses Pages 9–10 (high-precision) and Page 12 (low-precision).
+Following the planning phase, the coordination phase finally reclaims unused pages (Pages 6-7 from Head A and Page 11 from Head B) via a parallel prefix-sum.
+These recycled pages are appended to the circular free page list; since the end pointer already points to the list’s tail, it wraps around to the head to accommodate the recycled pages.
 Generation Phase.
@@
 Once a request is finished, all pages allocated for that request are recycled, freeing up memory for incoming requests.
+Supporting Additional Precision Levels.
+The proposed memory management system naturally extends to support more than two precision levels, by incorporating additional page tables.
+For example, accommodating three precision levels can be achieved by combining a unidirectional page table with a bidirectional one.
+Similarly, four precision levels can be supported using two bidirectional page tables.
+This composable design preserves memory efficiency while providing the flexibility needed to handle increasingly fine-grained precision differentiation.
 6 Implementation
-We implement LeanKV on top of vLLM, comprising 4.5K lines of CUDA/C++ code and 9K lines of Python.
-We first outline the architecture of LeanKV and then detail our custom GPU attention kernel, designed to efficiently support the differentiated KV cache compression.
-6.1 LeanKV Architecture
-The architecture of LeanKV is illustrated in Figure 6.
+We implement DiffKV on top of vLLM, comprising 4.5K lines of CUDA/C++ code and 9K lines of Python.
+We first outline the architecture of DiffKV and then detail our custom GPU attention kernel, designed to efficiently support the differentiated KV cache compression.
+6.1 DiffKV Architecture
+The architecture of DiffKV is illustrated in Figure 7.
 At each inference step, the scheduler batches as many requests as possible within the available GPU memory to maximize throughput, sending the selected requests to all workers.
-Each GPU hosts one worker, responsible for executing a segment of the model.
-LeanKV utilizes tensor parallelism [54], where each worker handles a partition of each matrix multiplication and a subset of the attention heads.
-Each worker includes a dedicated memory manager to oversee the KV cache for its assigned attention heads, using the method described in Section 5.
-Due to dynamic sparsity patterns, memory requirements vary per head for each request, leading to an irregular memory layout.
-Additionally, each worker incorporates an execution engine for model computation.
+Each GPU hosts one worker, responsible for executing a partition of the model [63, 33, 52, 53].
+To facilitate efficient execution, each worker includes a dedicated memory manager that oversees the KV cache for its assigned attention heads, as outlined in Section 5, and an execution engine for model computation.
 To support differentiated KV cache compression, the execution engine integrates a KV compressor and a custom GPU attention kernel.
@@
 The custom GPU attention kernel then efficiently computes the attention output using the compressed KV cache.
-Figure 6: LeanKV architecture.
+Figure 7: DiffKV architecture.
 6.2 Efficient Attention Kernel
-We develop a custom GPU attention kernel that efficiently supports differentiated KV cache compression, fully harnessing the reduction in memory access volume and translating it into accelerated performance.
-Overall, we assign each CUDA thread block to process a single attention head per sequence.
-However, mixed-precision quantization introduces potential load imbalance, as high-precision tokens demand more bandwidth, which is the bottleneck of attention, than low-precision ones.
-Thanks to unified pages that store tokens of the same precision together, we mitigate this issue by having thread warps iterate over high-precision pages first, followed by low-precision pages, ensuring that threads within a warp remain load-balanced.
+We develop a custom attention kernel that efficiently supports differentiated KV cache compression, leveraging reduced memory access volume to accelerate performance.
+Notably, the kernel design is portable across GPUs and other NPUs, as mainstream architectures support vectorized memory access and layout-aware tiling [37, 79].
+We detail the GPU implementation in the following paragraphs.
+Overall, we assign each thread block to process a single attention head per sequence.
+To mitigate the potential load imbalance caused by mixed-precision quantization, we let thread warps iterate over high-precision pages first, followed by low-precision pages.
 Each page is processed in two phases: dot product between the query and keys to derive attention scores, and weighted sum of values.
@@
 To ensure coalesced and vectorized memory accesses for both phases, we design tailored data layouts and parallelization strategies, elaborated as follows.
-For key processing, each warp handles one page at a time, with threads in the warp divided into groups responsible for distinct keys.
-Within a group, each thread fetches its assigned keys in a vectorized manner, performs dequantization on the fly, and computes a partial dot product between the keys and corresponding query.
-Using a straightforward layout such as [$F$, $N_{\text{tokens}}$] for keys in each page, where $F$ denotes the feature dimension length, would cause threads within a group to fetch non-contiguous elements strided by $N_{\text{tokens}}$ when parallelizing across the feature dimension.
-Alternatively, using [$N_{\text{tokens}}$, $F$] would cause thread groups to fetch non-contiguous elements strided by $F$ when parallelizing across tokens.
+For key processing, each warp handles one page at a time, with threads divided into groups responsible for distinct key vectors.
+Within a group, each thread fetches its assigned key elements in a vectorized manner, performs dequantization on the fly, and computes a partial dot product between the key elements and corresponding query elements.
+Using a straightforward layout such as [$F$, $N_{\text{tokens}}$] for keys, where $F$ denotes the feature dimension length, would cause threads within a group to fetch non-contiguous elements strided by $N_{\text{tokens}}$ when parallelizing across the feature dimension.
+Alternatively, using $N_{\text{tokens}}$, $F$] would cause thread groups to fetch non-contiguous elements strided by $F$ when parallelizing across tokens.
 Both layouts suffer inefficient strided memory accesses, leading to low bandwidth utilization.
@@
 As a result, the combined memory accesses of all threads in a warp are contiguous, enabling memory coalescing and maximizing bandwidth utilization.
-For value processing, tokens in a page are evenly distributed across thread groups in a warp; within each group, individual threads perform sum reductions over specific ranges of the feature dimension and save their accumulation results in registers.
+For value processing, tokens in a page are evenly distributed across thread groups in a warp; within each group, individual threads perform sum reductions over the feature dimension and save their accumulation results in registers.
 Once all groups finish, a tree reduction aggregates these partial results and produces the output.
 The number of registers required per thread is proportional to the feature dimension range assigned to it, while during key processing each thread only requires a single register to store the partial dot product.
-As a consequence, vectorization along the feature dimension, as used in key processing, is not suitable for value processing, because it would significantly increase register pressure by forcing each thread to handle a larger portion of the feature dimension.
-To better align with the computation pattern of reduction across tokens during value processing, we apply vectorization to the token dimension instead, which reduces register pressure and ensures more effective parallelization.
+Consequently, vectorization along the feature dimension, as used in key processing, is not suitable for value processing, because it would significantly increase register pressure by forcing each thread to handle a larger portion of the feature dimension.
+To better align with the computation pattern of reduction across tokens, we apply vectorization to the token dimension instead, which reduces register pressure and ensures more effective parallelization.
 We organize the layout of values as [ $\frac{F}{V_{\text{group}}}$, $\frac{N_{\text{tokens}}}{V_{\text{vec}}}$, $V_{\text{group}}$, $V_{\text{vec}}$] accordingly.
+Additionally, for ultra-long sequences, the attention kernel supports parallelization along the sequence dimension.
+It splits the sequence into multiple segments, processes them in parallel using the aforementioned method within separate thread blocks, and then merges the results with minimal computation overhead.
 7 Evaluation
-We first evaluate the effectiveness of LeanKV’s differentiated KV cache compression policy and analyze the sensitivity of different models to KV cache compression.
-Next, we evaluate the efficiency of the memory manager as well as the end-to-end throughput improvements achieved by LeanKV.
+We first evaluate the effectiveness of DiffKV’s differentiated KV cache compression policy.
+Next, we evaluate the efficiency of the memory manager as well as the end-to-end throughput improvements achieved by DiffKV.
 7.1 Experiment Setup
-We evaluate LeanKV on five models spanning three major families: Llama3-8B and 70B [60], Qwen2.5-7B and 32B [28], and the recently released thinking model QwQ-32B [58], which generates extended chains of thought and exhibits strong performance on complex reasoning tasks.
-Our evaluation covers a diverse set of mainstream benchmarks across multiple task domains, including general knowledge (MMLU [26] and MMLU-Pro [55]), mathematics (GSM8K [13] and MATH [34]), code generation (HumanEval+ [11, 39] and MBPP+ [7, 39]), and long-context understanding (LongBench [8]).
-In addition, we assess LeanKV’s performance on QwQ-32B using two particularly challenging benchmarks: AIME24 [3], a high-school level mathematics competition dataset, and GPQA [48], which evaluates graduate-level science reasoning.
-Throughput and latency are measured on NVIDIA L40 GPUs, each equipped with 48 GB of memory [45].
+We evaluate DiffKV on several models spanning three major families: Llama3-8B and 70B [68], Qwen2.5-7B and 32B [34], and the recent thinking models QwQ-32B [67], R1-Distill-Qwen-14B and R1-Distill-Llama-8B [26], which generate extended chains of thought and exhibit strong capabilities on complex reasoning tasks.
+We select benchmarks that align with those commonly used in recent LLM technical reports [20, 34, 27, 45, 67], covering a diverse set of task domains, including general knowledge (MMLU [31] and MMLU-Pro [64]), mathematics (GSM8K [15] and MATH [41]), code generation (HumanEval+ [12, 46] and MBPP+ [8, 46]), and long-context understanding (LongBench [9]).
+In addition, we assess DiffKV’s generation quality on thinking models using two particularly challenging benchmarks: AIME24 [3], a mathematics competition dataset, and GPQA [57], which evaluates graduate-level science reasoning.
+We measure the throughput and latency on NVIDIA L40 GPUs, each with 48 GB of memory [54].
+In all experiments, model weights are stored in FP16 precision.
 7.2 Differentiated KV Compression Policy
@@
 Evaluating Differentiated KV Quantization.
-Figure 7 shows the performance of differentiated KV quantization on GSM8K and HumanEval+, namely K8V4 and K4V2, compared to FP16 baselines, across Llama-3 8B & 70B and Qwen-2.5 7B.
-To validate our intuition that keys have a more significant impact than values (Section 3.1), we additionally evaluate the mirror configurations, namely K4V8 and K2V4, where values are stored with higher precision than keys.
+Figure 8 shows the model generation quality of differentiated KV quantization on GSM8K and HumanEval+, namely K8V4 and K4V2, compared to FP16 baselines, across Llama3-8B & 70B and Qwen2.5-7B.
+To validate our intuition that keys have a more significant impact than values (Section 3.1), we additionally evaluate the mirror configurations, namely K4V8 and K2V4, where values are stored with higher precision than keys, as well as more skewed variants, K8V2 and K4V1.
 Results show that K8V4 matches the accuracy of the FP16 baseline across all models and benchmarks.
-In contrast, its mirror configuration, K4V8, exhibits noticeable performance degradation, particularly for Qwen2.5-7B, where accuracy drops to nearly zero on both tasks.
-This pronounced sensitivity to 4-bit key quantization in Qwen2.5-7B is likely due to its GQA architecture, which applies aggressive KV compression with a queries-per-KV ratio of 7, substantially higher than the ratio of 4 used in Llama-3 8B.
-Similarly, K4V2 retains over 65% of FP16 accuracy across both benchmarks on Llama-3 8B and 70B, while the mirror configuration K2V4 results in near-zero accuracy.
-These results confirm that keys play a more critical role than values, and demonstrate the effectiveness of differentiated KV quantization in preserving accuracy while enabling aggressive compression.
-Figure 7: Performance of differentiated KV quantization.
-Figure 8: Performance of dynamic vs. static sparsity.
+In contrast, its mirror configuration, K4V8, exhibits noticeable accuracy degradation, particularly for Qwen2.5-7B, where accuracy drops to nearly zero on both tasks.
+This pronounced sensitivity to 4-bit key quantization in Qwen2.5-7B is likely due to its GQA architecture, which applies aggressive KV compression with a queries-per-KV ratio of 7, substantially higher than the ratio of 4 used in Llama3-8B.
+Similarly, K4V2 retains over 65% of FP16 accuracy across both benchmarks on Llama3-8B and 70B, while the mirror configuration K2V4 results in near-zero accuracy.
+The more skewed variant K8V2 retains at least 83% of FP16 accuracy across all models, even for Qwen2.5-7B, reinforcing the importance of assigning higher precision to keys.
+On the other hand, K4V1 yields almost zero accuracy, suggesting that 2 bits is the lower bound for effective value quantization.
+In summary, these results confirm that keys play a more critical role than values, and demonstrate the effectiveness of differentiated KV quantization in preserving accuracy while enabling aggressive compression.
+The accuracy trends across various differentiated KV quantization configurations motivate the design of adopting two precision levels: K8V4 and K4V2.
+K8V4 matches the generation quality of the FP16 baseline and is applied to important tokens to preserve model accuracy.
+K8V2 and K4V2, when applied uniformly, both lead to noticeable accuracy degradation.
+Among them, K4V2 offers better efficiency and is thus selected for compressing less significant tokens, enabling opportunistic KV cache savings with minimal quality loss.
+As we demonstrate later, the K8V4–K4V2 scheme achieves near-lossless generation quality across multiple models and benchmarks, obviating the need for an intermediate precision level like K8V2.
+Likewise, introducing an even lower precision level such as K4V1 is suboptimal: not only does K4V1 significantly degrade accuracy, but the additional memory savings from a three-level configuration K8V4–K4V2–K4V1 are marginal compared to the K8V4–K4V2 design.
+That said, future models employing more aggressive training-time KV compression (e.g., architectures with higher queries-per-KV ratios in GQA) may benefit from an additional high-precision level (e.g., FP16–K8V4 or FP16–K8V4–K4V2), which we leave to future investigation.
+Figure 8: Accuracy of differentiated KV quantization.
+Figure 9: Accuracy of dynamic vs. static sparsity.
 Evaluating Dynamic Sparsity.
-We evaluate the effectiveness of per-head dynamic sparsity (Section 3.3) in identifying critical tokens, comparing it to the static sparsity method used in SnapKV [35] and H2O [69], which allocates an equal memory budget to all attention heads.
+We evaluate the effectiveness of per-head dynamic sparsity (Section 3.3) in identifying critical tokens, comparing it to the static sparsity method used in SnapKV [42], which allocates an equal memory budget to all attention heads.
 Figure 9 reports results on Qwen2.5-7B and Llama3-8B across GSM8K and HumanEval+.
@@
 In summary, per-head dynamic sparsity is superior to static sparsity by leveraging head-specific significance.
-Figure 9: Calibrating the high- and low-precision threshold parameters $\alpha_{h}$ and $\alpha_{l}$ on the training split of the MATH dataset.
+Figure 10: Calibrating the high- and low-precision thresholds $\alpha_{h}$ and $\alpha_{l}$ on the training split of the MATH dataset.
 Parameter Calibration.
-As described in Section 4, LeanKV compares the significance score of a token with its theoretical average, $\frac{1}{N}$, where $N$ denotes the sequence length.
-A token is quantized to high precision if its score exceeds $\frac{\alpha_{h}}{N}$, to low precision if its score falls within the interval $[\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, and is pruned otherwise.
+As described in Section 4, DiffKV compares the significance score of a token with its theoretical average, $\frac{1}{N}$, where $N$ denotes the sequence length.
+A token is quantized to high precision if its score exceeds $\frac{\alpha_{h}}{N}$, to low precision if its score falls within the interval $\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, and is pruned otherwise.
 To calibrate the appropriate parameters $\alpha_{h}$ and $\alpha_{l}$, we utilize the training split of MATH [41].
 We select MATH as the calibration dataset as it is comprised of complex mathematical reasoning tasks with high text information density, where even the removal of a single symbol can invalidate an argument.
-This makes MATH a highly suitable choice for parameter fitting, as the resulting compression parameters are expected to generalize well across other tasks.
+This makes MATH highly suitable for parameter fitting, as the resulting compression parameters are expected to generalize well across other tasks.
 Additionally, MATH provides a dedicated training split, ensuring that parameter fitting does not involve testing data and mitigating potential overfitting.
@@
 The profiled values of $\alpha_{h}$ achieve FP16-equivalent accuracy across all models, except for Qwen2.5-7B, which exhibits heightened sensitivity to 4-bit key quantization.
-Based on our profiling, we select the values of $\alpha_{h}$ that yield the highest accuracy for each model: $\alpha_{h}=1$ for Llama-3 8B and 70B, $\alpha_{h}=3$ for Qwen2.5-32B and QwQ-32B, and $\alpha_{h}=0$ for Qwen2.5-7B.
+Accordingly, we select the values of $\alpha_{h}$ that yield the highest accuracy based on profiling results: $\alpha_{h}=1$ for Llama3-8B and 70B, and $\alpha_{h}=3$ for Qwen2.5-32B and QwQ-32B, and disable low-precision quantization for Qwen2.5-7B.
 Subsequently, we profile the low-precision threshold $\alpha_{l}$, while keeping $\alpha_{h}$ fixed.
@@
 As a result, we profile five equally spaced values within the range $[0,0.1]$ and select the one yielding the highest accuracy for each model.
-The final selected values of $\alpha_{l}$ are $0.02$ for Llama3-8B, $0.04$ for Qwen2.5-7B, and 00 for the remaining models.
-While the profiled values cover a limited portion of the entire parameter space, as will be demonstrated in the following sections, LeanKV significantly outperforms existing baselines with these fitted parameters, highlighting its potential effectiveness.
-Figure 10: KV cache memory normalized to vLLM vs. benchmark performance tradeoff of LeanKV.
+The final selected values of $\alpha_{l}$ are $0.02$ for Llama3-8B, $0.04$ for Qwen2.5-7B, and 0 for the remaining models.
+While the profiled values cover a limited portion of the entire parameter space, as will be demonstrated in the following sections, Diff-KV significantly outperforms existing baselines with these fitted parameters, highlighting its potential effectiveness.
+Figure 11: KV cache memory normalized to vLLM vs. benchmark accuracy tradeoff of DiffKV.
 Evaluating Differentiated Compression Policy.
-Finally, we evaluate the combined benefits of LeanKV’s KV cache compression policy.
-We compare LeanKV against several state-of-the-art baselines, including pruning-based methods such as H2O [69], SnapKV [35], and DuoAttention [63], quantization-based approaches such as 4-bit KV and KIVI [41], as well as KV partial loading techniques like Quest [56].
-Table 1 summarizes the accuracy and KV cache memory usage of LeanKV normalized to vLLM [32] baseline, with the parameters tuned as described, for the evaluated non-thinking models.
+Finally, we evaluate the combined benefits of DiffKV’s KV cache compression policy.
+We compare DiffKV against several state-of-the-art baselines, including pruning-based methods such as H2O [76], SnapKV [42], and DuoAttention [70], quantization-based approaches such as 4-bit KV, KIVI [48], and QAQ [19], as well as KV partial loading techniques like Quest [65].
+Table 1 compares the accuracy and KV cache memory usage of DiffKV and the best-performing baselines, both normalized to vLLM [39], for the evaluated non-thinking models.
+The parameters of DiffKV are tuned as described previously.
 For the pruning-based baselines and Quest, we tune their memory usage to 50% of the baseline.
-LeanKV achieves near-lossless accuracy relative to the FP16 baseline, with an average degradation of only $0.3\%$, while using $19.3\%$ to $36.7\%$ memory.
+DiffKV achieves near-lossless accuracy relative to the FP16 baseline, with an average degradation of only $0.3\%$, while using $19.3\%$ to $36.7\%$ memory.
 In contrast, all other baselines incur more significant accuracy degradation at comparable or higher memory usage.
-Furthermore, LeanKV dynamically adapts its memory usage based on the task’s information density.
-For example, LeanKV allocates less memory for the 5-shot MMLU benchmark compared to the 0-shot HumanEval+, where the prompt consists solely of the function name and comments.
-LeanKV also outperforms baseline methods in long-context scenarios, as shown in Table 2, where we present results from one benchmark in each category of LongBench.
-For Llama3-8B and Qwen2.5-7B, LeanKV achieves memory usage of $14.9\%$ and $27.0\%$, respectively, while baseline methods are tuned to $25\%$ memory usage.
-Notably, LeanKV consistently achieves superior accuracy across all categories, with an average degradation of only $0.4\%$, while utilizing less or comparable memory compared to the baseline methods.
-Furthermore, the performance gap between LeanKV and baseline methods becomes even more pronounced for the thinking model QwQ-32B on more challenging benchmarks, as demonstrated in Table 3.
-LeanKV matches the FP16 baseline performance using only $27.4\%$ of the memory on average, while all other methods experience significant performance degradation.
+Furthermore, DiffKV dynamically adapts its memory usage based on the task’s information density.
+For example, DiffKV allocates less memory for the 5-shot MMLU benchmark compared to the 0-shot HumanEval+, where the prompt consists solely of the function name and comments.
+DiffKV also outperforms baseline methods in long-context scenarios, as shown in Table 2, where we present results from one benchmark in each category of LongBench.
+For Llama3-8B and Qwen2.5-7B, DiffKV achieves memory usage of $14.9\%$ and $27.0\%$, respectively, while baseline methods are tuned to $25\%$ memory usage.
+Notably, DiffKV consistently achieves superior accuracy across all categories, with an average degradation of only $0.4\%$, while utilizing less or comparable memory compared to the baseline methods.
+Furthermore, the generation quality gap between DiffKV and baseline methods becomes even more pronounced for the thinking models QwQ-32B, R1-Distill-Qwen-14B and R1-Distill-Llama-8B on the more challenging benchmarks, as demonstrated in Table 3.
+DiffKV achieves FP16-comparable generation quality using $27.4\%$, $29.4\%$ and $23.5\%$ of the memory on average for QwQ-32B, R1-Distill-Qwen-14B and R1-Distill-Llama-8B, respectively, while all other methods experience significant generation quality degradation.
 On the math competition benchmark AIME24, the most accurate baseline, 4-bit KV quantization incurs a $16.0\%$ accuracy degradation.
-On the graduate-level science benchmark GPQA, the accuracy of pruning-based and KV partial loading methods drops to almost zero.
-Lastly, Figure 10 depicts the performance-memory trade-offs of LeanKV across representative benchmarks under the profiled parameter values, namely $\alpha_{h}\in[0,5]$ (step size of 1) for all models and $\alpha_{l}\in[0.02,0.1]$ (step size of 0.02) for Llama3-8B and Qwen2.5-7B.
+On the graduate-level science benchmark GPQA, the accuracy of pruning-based and KV partial loading methods drops to almost zero, while using $50\%$ of the full KV cache.
+The larger gap between DiffKV and prior work in long CoT generation highlights the increased challenges such tasks pose for KV cache compression.
+In long CoT generation, tokens are generated autoregressively based on previous tokens, and errors introduced by compression are accumulated and propagated across generation steps.
+In contrast, in long prompt scenarios, the majority of the text is provided as ground truth in the prompt, and the model-generated portion is significantly shorter, limiting the impact of errors introduced by KV cache compression.
+Lastly, Figure 11 depicts the accuracy-memory trade-offs of DiffKV under the profiled parameter values, namely $\alpha_{h}\in[0,5]$ (step size of 1) for all models and $\alpha_{l}\in[0.02,0.1]$ (step size of 0.02) for Llama3-8B and Qwen2.5-7B.
 We also include the accuracy of baseline methods within similar memory usage ranges for comparison.
-LeanKV consistently matches the FP16 baseline accuracy across different models and benchmarks within the profiled range of memory usage, demonstrating its robustness.
+DiffKV consistently matches the FP16 baseline accuracy across different models and benchmarks within the profiled range of memory usage, demonstrating its robustness.
 In contrast, baseline methods experience significant accuracy degradation, particularly pruning-based and KV partial loading methods, which suffer rapid accuracy declines as memory usage decreases.
-In summary, LeanKV provides a superior performance-memory tradeoff compared to state-of-the-art methods.
-Table 1: Accuracy and memory usage of LeanKV and baseline methods across models and benchmarks.
+In summary, DiffKV provides a superior accuracy-memory tradeoff compared to state-of-the-art methods.
+Table 1: Accuracy and memory usage of DiffKV and the best-performing baseline methods across models and benchmarks.
 Numbers in parentheses represent memory usage normalized to FP16.
-The memory usage is 50% for SnapKV, Quest, H2O and DuoAttn, 25% for INT4, and 12.5% for KIVI due to 2-bit quantization.
+The memory usage is 50.0% for SnapKV, Quest and DuoAttn, 25.0% for INT4, 18.8% for QAQ, and 12.5% for KIVI.
+DiffKV achieves accuracy close to FP16, consistently outperforming prior quantization and pruning approaches.
 […]
-Table 2: Evaluating LeanKV on LongBench. LeanKV uses 14.9% memory for Llama3.1-8B and 27.0% for Qwen2.5-7B. For Quest and SnapKV, the memory usage is 25%.
+Table 2: Evaluating DiffKV on LongBench. DiffKV uses 14.9% memory for Llama3.1-8B and 27.0% for Qwen2.5-7B. For Quest and SnapKV, the memory usage is 25%.
 […]
-Table 3: Evaluating LeanKV on QwQ-32B thinking model.
+Table 3: Evaluating DiffKV on thinking models.
 […]
+Qwen-14B
+[…]
+[…]
+Llama-8B
+[…]
+[…]
+KV Compression Breakdown.
+Figure 12 shows the fraction of tokens that are pruned, quantized to low precision, and quantized to high precision across three benchmarks (MMLU, HumanEval+, and MATH) for three different models.
+Two key observations emerge.
+First, sparsity levels vary substantially across workloads.
+The general knowledge benchmark MMLU exhibits the highest sparsity, likely due to the lower information density of natural language and the use of 5-shot prompts.
+In contrast, the reasoning-intensive benchmarks MATH and HumanEval+ display lower sparsity, with HumanEval+ being the least sparse, partly due to its 0-shot setting.
+Second, despite being calibrated on challenging reasoning-heavy workloads, DiffKV’s KV compression policy adapts effectively to the varying sparsity patterns across different tasks, highlighting its generalizability to diverse scenarios.
+Figure 12: Fraction of tokens pruned, quantized to low precision, and quantized to high precision across different benchmarks and models.
 7.3 System Performance
-In this section, we provide a comprehensive evaluation of LeanKV’s performance, comparing it against state-of-the-art systems including vLLM, Quest, and SnapKV across various models.
-All experiments are conducted on a machine equipped with eight NVIDIA L40 GPUs.
+In this section, we provide a comprehensive evaluation of DiffKV’s performance, comparing it against state-of-the-art systems including vLLM, Quest, and SnapKV across various models.
 For models that fit within the memory of a single GPU, such as Llama3-8B and Qwen2.5-7B, we use one GPU.
 For larger models, we parallelize the computation across multiple GPUs: four GPUs for Llama3-70B, and two GPUs for Qwen2.5-32B and QwQ-32B.
-Our evaluation focuses on latency breakdown, attention kernel speedup, and end-to-end throughput to demonstrate the efficiency and scalability of LeanKV.
+Our evaluation focuses on latency breakdown, attention kernel speedup, and end-to-end throughput to demonstrate the efficiency and scalability of DiffKV.
 Memory Management Overhead.
-Figure 11 compares LeanKV’s on-GPU parallel KV compaction with an alternative implementation that performs multi-threaded memory management on the CPU.
-Evaluated with a sequence length of 1024 tokens across different batch sizes, LeanKV reduces memory management latency by up to three orders of magnitude relative to the on-CPU approach.
-For one entire inference step, LeanKV significantly outperforms the on-CPU approach, particularly in the generation phase, where the memory management overhead on the CPU far exceeds the model execution time.
-Therefore, LeanKV’s on-GPU parallel memory management is essential for ensuring that the performance benefits of KV cache compression are fully realized without being overshadowed by memory management overhead.
-Figure 12 further illustrates the latency breakdown of LeanKV during one inference step.
-The memory management overhead is remarkably low thanks to LeanKV’s on-GPU parallel KV compaction, contributing less than 0.2% of the total latency in the prompt phase and under 0.9% in the generation phase.
+Figure 13 compares Diff-KV’s on-GPU parallel KV compaction with an alternative implementation that performs multi-threaded memory management on the CPU.
+Evaluated with a sequence length of 1024 tokens across different batch sizes, DiffKV reduces memory management latency by up to three orders of magnitude relative to the on-CPU approach.
+For one entire inference step, DiffKV significantly outperforms the on-CPU approach, particularly in the generation phase, where the memory management overhead on the CPU far exceeds the model execution time.
+Therefore, DiffKV’s on-GPU parallel memory management is essential for ensuring that the performance benefits of KV cache compression are fully realized without being overshadowed by memory management overhead.
+Figure 14 further illustrates the latency breakdown of Diff-KV during one inference step.
+The memory management overhead is remarkably low thanks to DiffKV’s on-GPU parallel KV compaction, contributing less than 0.2% of the total latency in the prompt phase and under 0.9% in the generation phase.
 Model execution dominates the latency in both phases, accounting for 96–97% of the latency in the prompt phase and 92–93% in the generation phase.
-Notably, as the batch size increases, the percentage of time spent on model execution rises slightly, reflecting the scalability of LeanKV.
+Notably, as the batch size increases, the percentage of time spent on model execution rises slightly, reflecting the scalability of DiffKV.
 (a) Memory management
@@
 Figure 13: Latency comparison between parallel KV compaction and on-CPU multi-threaded memory management.
-Figure 12: Latency breakdown of LeanKV.
-Attention Kernel Speedup.
-Figure 13 shows the speedup of the LeanKV’s custom attention kernel against vLLM under different quantization configurations.
-LeanKV achieves a near-linear speedup proportional to the reduction in KV cache size.
-For example, with K8V8, which halves the KV cache size relative to FP16 datatype, the theoretical speedup is 2$\times$, and LeanKV achieves 1.7$\times$.
-The slight gap is primarily due to the overhead of accessing quantization metadata and performing dequantization.
-Additionally, LeanKV achieves greater speedups on longer sequences, indicating that its bandwidth optimization techniques become increasingly effective as the sequence length grows.
-Figure 13: Attention kernel speedup.
-End-to-End Throughput.
-Figure 14 presents the end-to-end throughput and achieved batch sizes of LeanKV compared to vLLM, Quest, and SnapKV.
+Figure 14: Latency breakdown of DiffKV.
+[…]
+(a) Attention kernel
+(b) One batch end-to-end
+Figure 15: Latency speedup of DiffKV against vLLM.
+Throughput Speedup.
+Figure 17 presents the end-to-end throughput and achieved batch sizes of DiffKV compared to vLLM, pruning-based approaches including Quest and SnapKV, as well as quantization-based methods including Atom (4-bit) and KIVI (2-bit).
+Note that these pruning- and quantization-based baselines incur larger accuracy degradation than DiffKV, as shown in Table 1.
 The maximum generation length is set to 16K tokens for QwQ-32B, 8K for Qwen2.5-32B, and 4K for the other models.
-We evaluate on 1000 sequences sampled from the MATH dataset [34], which naturally elicits chain-of-thought reasoning and typically leads to long generations reaching the specified limit.
-Compression thresholds for LeanKV are adopted from Figure 9.
-Across all models, LeanKV consistently achieves higher throughput than prior systems.
-Notably, LeanKV achieves a remarkable 5.4×5.4$\times$5.4 × higher throughput over vLLM on the QwQ-32B thinking model, while Quest and SnapKV achieve only 1.6×1.6$\times$1.6 × and 1.8×1.8$\times$1.8 × speedups, respectively.
-This result highlights the superior efficiency of LeanKV in processing long sequences, making it particularly well-suited for tasks that require extended reasoning.
-The throughput improvements are directly correlated with the larger batch sizes supported by LeanKV, enabled by its KV cache compression techniques.
-For instance, with the QwQ-32B model, LeanKV sustains a batch size of 15.9, far exceeding vLLM’s batch size of 2.7.
-The increased batch size enables LeanKV to maximize the utilization of GPU compute capacity, leading to higher throughput.
+We evaluate on 1000 sequences sampled from the MATH dataset [41], which elicits chain-of-thought reasoning and typically leads to long generations reaching the specified limit.
+Compression thresholds for DiffKV are adopted from Figure 10.
+Across all models, DiffKV consistently achieves higher throughput than prior systems.
+Notably, DiffKV achieves a remarkable $5.4\times$ higher throughput over vLLM on the QwQ-32B thinking model, while Quest, SnapKV, Atom, and KIVI achieve $1.6\times$, $1.8\times$, $2.1\times$, and $3.4\times$ speedups, respectively.
+This result highlights the superior efficiency of DiffKV in processing long sequences, making it particularly well-suited for tasks that require extended reasoning.
+The throughput improvements are directly correlated with the larger batch sizes supported by DiffKV, enabled by its KV cache compression techniques.
+For instance, with the QwQ-32B model, DiffKV sustains a batch size of 15.9, far exceeding vLLM’s batch size of 2.7.
+The increased batch size enables DiffKV to maximize the utilization of GPU compute capacity, leading to higher throughput.
 In contrast, Quest supports the same batch size as vLLM because it retains the entire KV cache without reducing memory usage.
@@
 However, the process of estimating token importance incurs additional overhead, limiting the overall efficiency of Quest.
+Note that although Atom and KIVI substantially increase the batch size, approaching or even exceeding DiffKV, their throughput gains are not proportional.
+This is because both methods are built on the HuggingFace Transformers library, which incurs higher framework overhead than vLLM and DiffKV, and lacks high-performance GPU kernels that fuse dequantization with attention computation, as implemented in DiffKV.
+These results underscore that custom GPU kernels and an efficient runtime are essential for effectively translating KV cache compression into throughput gains.
+Serving Dynamic Workloads.
+We evaluate DiffKV under dynamic workloads and compare it to vLLM [39].
+Following the methodology of vLLM, we inject requests according to a Poisson arrival process and sweep the request rate.
+Figure 16 reports the average per-token latency (including both queuing and LLM processing time) for Llama3-8B and Qwen2.5-32B across a range of request rates.
+DiffKV consistently achieves lower latency and sustains higher loads before queuing delays grow sharply.
+These improvements stem from DiffKV’s reduced KV cache footprint, which enables larger batch sizes, and its parallel on-GPU KV compaction, which keeps memory management overhead low even under bursty traffic.
+Figure 16: Comparison of average latency between DiffKV and vLLM under dynamic workloads.
 Figure 17: Throughput and achieved batch size.
 8 Conclusion
-We propose LeanKV, a framework to enhance LLM serving efficiency by exploiting three levels of differentiation in KV cache: differentiated precision for keys and values, hierarchical compression based on varying token importance, and per-head dynamic sparsity.
-At the core of LeanKV is the parallel KV compaction technique that efficiently handles irregular memory requirements across requests and attention heads, effectively translating memory savings into performance gains.
-Our evaluation shows that LeanKV is able to compress the KV cache by 2.7×2.7$\times$2.7 × to 5.7×5.7$\times$5.7 × without compromising accuracy on complex workloads that require sophisticated reasoning and long-generation capabilities, and enhances throughput by 1.9×1.9$\times$1.9 × to 5.4×5.4$\times$5.4 ×.
+We proposed DiffKV, which enhances LLM serving efficiency by exploiting three levels of differentiation in KV cache: differentiated precision for keys and values, hierarchical compression based on token importance, and per-head dynamic sparsity.
+At the core of DiffKV is the parallel KV compaction technique that efficiently handles irregular memory requirements across requests and attention heads, effectively translating memory savings into performance gains.
+Our evaluation demonstrated that DiffKV is able to compress the KV cache by $2.7\times$ to $5.7\times$ with near-lossless accuracy, even on thinking models and complex workloads that require sophisticated reasoning and long-generation capabilities, and enhances throughput by $1.9\times$ to $5.4\times$, outperforming prior KV cache compression methods.
+9 Acknowledgment
+We sincerely thank our shepherd, Tim Harris, for his thoughtful feedback and careful guidance throughout the paper shepherding process.
+We are also thankful to Mingzhe Hao, Shuang Chen, other friends, and the anonymous reviewers for their insightful comments on earlier versions of this manuscript.